### v1

# URS: A Unified Neural Routing Solver for Cross-Problem Zero-Shot Generalization

Abstract
Multi-task neural routing solvers have emerged as a promising paradigm for their ability to solve multiple vehicle routing problems (VRPs) using a single model. However, existing neural solvers typically rely on predefined problem constraints or require per-problem fine-tuning, which substantially limits their zero-shot generalization ability to unseen VRP variants. To address this critical bottleneck, we propose URS, a unified neural routing solver capable of zero-shot generalization across a wide range of unseen VRPs using a single model without any fine-tuning. The key component of URS is the unified data representation (UDR), which replaces problem enumeration with data unification, thereby broadening the problem coverage and reducing reliance on domain expertise. In addition, we propose a Mixed Bias Module (MBM) to efficiently learn the geometric and relational biases inherent in various problems. On top of the proposed UDR, we further develop a parameter generator that adaptively adjusts the decoder and bias weights of MBM to enhance zero-shot generalization. Moreover, we propose an LLM-driven constraint satisfaction mechanism, which translates raw problem descriptions into executable stepwise masking functions to ensure solution feasibility. Extensive experiments demonstrate that URS can consistently produce high-quality solutions for more than 100 distinct VRP variants without any fine-tuning, which includes more than 90 unseen variants. To the best of our knowledge, URS is the first neural solver capable of handling over 100 VRP variants with a single model.
1 Introduction
The Vehicle Routing Problem (VRP) is an essential class of combinatorial optimization problems (COPs) with extensive applications in logistics and supply chain management (Tiwari & Sharma, 2023; Sar & Ghadimi, 2023). Solving VRPs efficiently is challenging due to their NP-hard nature. While exact solvers can produce optimal solutions, they often become computationally prohibitive for real-world scenarios. In recent decades, classical heuristic solvers (Helsgaun, 2017; Vidal, 2022) have achieved impressive performance within acceptable timeframes. Nevertheless, these methods require considerable domain expertise to design specialized rules for each routing problem. Given the growing diversity of VRP variants in real-world applications, manually crafting tailored rules for every case has become impractical.
In recent years, neural combinatorial optimization (NCO) methods have attracted substantial attention for their potential to reduce reliance on handcrafted rules while maintaining competitive solution quality (Bengio et al., 2021; Li et al., 2022a). This has led to the development of many high-performing neural routing solvers that can automatically learn implicit problem-specific patterns from data under different training paradigms such as supervised learning (SL) (Drakulic et al., 2023; Luo et al., 2023; 2025b), reinforcement learning (RL) (Bello et al., 2016; Kool et al., 2019; Zhou et al., 2024a), and emerging self-improved learning (SIL) (Luo et al., 2023; 2025a; Pirnay & Grimm, 2024). Although these solvers have demonstrated impressive performance on specific problems (Kwon et al., 2020; Kim et al., 2022; Zhou et al., 2025), they typically require architectural customization and per-problem retraining to accommodate the distinct constraints and features of different VRP variants. These limitations increase the overall training cost and hinder practical deployment on new problems.
Table 1: Comparison between our URS and existing neural solvers with multi-task learning.
Neural Solver Training Paradigm Routing Problems (#\#) Zero-shot Generalization Remarks
MTPOMO (Liu et al., 2024a) RL 16 $\checkmark$ Constraint Combination
MVMoE (Zhou et al., 2024b) RL 16 $\checkmark$ Constraint Combination
RouteFinder (Berto et al., 2024) RL 48 $\checkmark$ Constraint Combination
CaDA (Li et al., 2025a) RL 16 $\checkmark$ Constraint Combination
MTL-KD (Zheng et al., 2025a) RL+KD 16 $\checkmark$ Constraint Combination
TSP-FT (Lin et al., 2024) RL 3 $\times$ Adapter-based Fine-tuning
MTL-MAB (Wang et al., 2025a) RL 3 $\times$ Adapter-based Fine-tuning
GOAL (Drakulic et al., 2025) SL 9 $\times$ Adapter-based Fine-tuning
[…]
To address the challenge of cross-problem generalization, growing attention has been directed toward multi-task learning capable of handling diverse routing problems. As summarized in Table 1, existing efforts largely fall into two categories: 1) constraint combination-based multi-task learning and 2) adapter-based fine-tuning frameworks. In the first approach, VRP variants are treated as different combinations of constraint attributes, and a unified model is trained across these combinations to enable knowledge sharing for problems with seen constraints (Liu et al., 2024a; Zhou et al., 2024b; Li et al., 2025a; Berto et al., 2024; Zheng et al., 2025a). The second approach builds a shared model backbone for all problems and incorporates specific input/output adapters for each problem, thereby reducing the retraining costs (Lin et al., 2024; Drakulic et al., 2025; Wang et al., 2025a).
Despite the above advancements, existing methods still fall short in zero-shot generalization. The problem coverage of constraint combination-based approaches is inherently bounded by their manually specified constraint sets, while adapter-based approaches require fine-tuning and thus cannot perform zero-shot generalization. Moreover, both strategies fundamentally rely on the explicit enumeration of problems through a predefined set of problem tags. This practice is problematic because the constraint space of routing problems is open-ended and compositional, making any fixed tagging scheme fundamentally limited. More critically, creating and maintaining such a problem taxonomy requires considerable domain expertise, which is precisely what NCO always aims to avoid. A detailed review of related work on single-task and multi-task learning is provided in Appendix A.
In this paper, we propose a powerful Unified Neural Routing Solver (URS) to significantly improve the cross-problem zero-shot generalization ability for NCO methods. Unlike existing approaches, we introduce a unified data representation (UDR) that replaces problem enumeration with data unification, thereby broadening problem coverage and reducing reliance on domain expertise. In addition, we propose a Mixed Bias Module (MBM) to better learn the geometric and relational biases inherent in various problems. Based on UDR, we further develop a parameter generator that adaptively adjusts decoder and bias weights of MBM to improve zero-shot generalization. Furthermore, to handle complex and unseen routing problems, we introduce a constraint satisfaction mechanism that leverages a large language model (LLM) to translate raw problem constraint descriptions into executable node masking functions for each construction step, thereby ensuring our URS generates feasible solutions for various problems.
We evaluate URS on more than 100 VRP variants, including more than 90 unseen variants. The results demonstrate that URS not only obtains competitive performance against specialized neural solvers on seen variants but also exhibits strong zero-shot generalization capabilities on a wide range of unseen problems. To the best of our knowledge, URS is the first neural solver capable of efficiently solving over 100 distinct VRP variants using a single model without any retraining or fine-tuning.
2 Preliminaries
In this section, we first introduce the definition of VRP, then provide an overview of recent constructive neural solvers for solution generation (Kool et al., 2019; Kwon et al., 2020).
2.1 Vehicle Routing Problems
Consider a VRP instance with an optional depot indexed by 0 and $n$ customers indexed by $\{1,2,\ldots,n\}$, such as in the Capacitated VRP (CVRP). The instance can be represented as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where the node set $\mathcal{V}=\{v_{i}\}_{i=0}^{n}$ has a total size of $|\mathcal{V}|=1+n$ unless the variant has no depot (e.g., the Traveling Salesman Problem), and the edge set is $\mathcal{E}=\{e(v_{i},v_{j})\ |\ v_{i},v_{j}\in\mathcal{V},v_{i}\neq v_{j}\}$. The travel cost between nodes is given by a distance matrix $\bm{D}=\{d_{ij}|\ \forall\ i,j\in 0,\ldots,n\}$. In VRP, each node $v_{i}\in\mathcal{V}$ includes node coordinates $\{x_{i},y_{i}\}$ when available and problem-specific attributes (e.g., demands in CVRP). A feasible solution is a tour $\bm{\pi}=(\pi_{1},\pi_{2},\ldots,\pi_{m})$ (i.e., a finite sequence), which is a permutation of $m$ nodes. Let $\Omega$ denote the feasible sequence set that satisfies the constraints, given an objective function $f(\bm{\pi}|\mathcal{G})$, we aim to search for a sequence $\bm{\pi}^{*}$ with the maximum objective, i.e., $\bm{\pi}^{*}=\arg\max_{\bm{\pi}\in\Omega}f(\pi|\mathcal{G})$. In most routing problems, the objective can be represented as the negative value of the total distance of a feasible sequence $\bm{\pi}$.
In this paper, we consider over 100 VRP variants which may simultaneously have one or more constraints from the following categories: (1) Capacity (C); (2) Open Route (O); (3) Backhaul (B); (4) Backhaul and Priority (BP); (5) Duration Limit (L); (6) Time Windows (TW); (7) Multi-Depot (MD); (8) Prize Collecting (PC); (9) Asymmetric (A); (10) Pickup and Delivery (PD). Detailed definitions of these constraints are provided in Appendix B.
2.2 Constructive Neural Routing Solver
Most constructive NCO methods employ an encoder-decoder architecture (Luo et al., 2023; Kwon et al., 2020). Let learnable model parameters be denoted as $\bm{\theta}=\{\bm{\theta}_{enc},\bm{\theta}_{dec}\}$. Without loss of generality, we present the prevailing construction pipeline using AM (Kool et al., 2019). Given an instance $\mathcal{G}$, raw node features are first mapped by a linear projection to initial embeddings $H^{(0)}=\{\mathbf{h}_{i}^{(0)}\}_{i=0}^{n}$. They are then processed by an encoder $\bm{\theta}_{enc}$ with $L$ attention layers to produce refined node embeddings $H^{(L)}=\{\mathbf{h}_{i}^{(L)}\}_{i=0}^{n}$. At each decoding step $t$, the decoder $\bm{\theta}_{dec}$ sequentially selects a node to append to the current partial solution $\pi_{1:t-1}=(\pi_{1},\pi_{2},...,\pi_{t-1})$, where $\pi_{1},\pi_{t-1}\in\mathcal{V}$ are the first and last visited node, respectively. This sequential process continues until a complete solution is formed. To guarantee the feasibility of generated solutions, a masking function $\mathcal{M}_{t}$ sets the selection probabilities of the following nodes to -$\infty$ during the construction process: nodes that (1) are visited or (2) violate problem-specific constraints (e.g., capacity).
3 The Proposed Model: Unified Neural Routing Solver
In this section, we propose a powerful Unified Neural Routing Solver (URS), which can significantly improve cross-problem zero-shot generalization for NCO methods. The framework is illustrated in Figure 1, and its key components are elaborated below.
Figure 1: The overview of our URS with a 44‑node APDTSP instance as an example. (a) The solving pipeline of URS; (b) The proposed mixed bias module that incorporates multiple geometric priors, enabling it to efficiently learn the geometric and relational biases inherent in various problems.
3.1 Unified Data Representation
Existing multi-task neural solvers (Liu et al., 2024a; Zhou et al., 2024b; Drakulic et al., 2025; Wang et al., 2025a) fundamentally rely on the explicit problem enumeration through a predefined set of problem tags, which is challenging because the constraint space of routing problems is open-ended and compositional, making any fixed tagging scheme fundamentally limited. More critically, creating and maintaining such a problem taxonomy requires considerable domain expertise, which is precisely what NCO always aims to avoid.
From a data perspective, despite the substantial diversity in constraints among routing problems, their instances share a common instance representation, e.g., all input instances under the current constraint combinations (Liu et al., 2024a; Zhou et al., 2024b) can be formulated as instances of either CVRP or CVRPTW. Leveraging this commonality, we propose a Unified Data Representation (UDR) to replace problem enumeration with data unification, thereby broadening problem coverage and reducing reliance on domain expertise. For an instance $\mathcal{G}$, we define the unified representation $\bm{U}=\{\mathbf{u}_{i}\}_{i=0}^{n}$, each $\mathbf{u}_{i}=\{\bm{\rho}_{i},\ \bm{\omega}_{i},\ \bm{\xi}_{i}\}$, where $\bm{\rho}_{i},\ \bm{\omega}_{i},\ \bm{\xi}_{i}$ are a position identifier, unified attribute set, and node-type indicator, respectively.
Position Identifier ($\bm{\rho}_{i}$)
We define a unified positional identifier for each node as $\bm{\rho}_{i}=\{\eta_{i},x_{i},y_{i}\}\in[0,1]^{3}$. The $\eta_{i}\sim\mathrm{Uniform}(0,1)\in\mathbb{R}^{1}$ is an additional sampled scalar, which is used to address asymmetric problems, following  Drakulic et al. (2023). For symmetric cases, we simply set $\eta_{i}=0$.
Unified Attribute Set ($\bm{\omega}_{i}$)
We define the unified attribute set as $\bm{\omega}_{i}=\{\delta_{i},\epsilon_{i},\mu_{i},\ e_{i},\ l_{i},\ s_{i}\}$, where they correspond to demand, prize, penalty, earliest arrival time, latest arrival time, and service time, respectively. The unification empowers the model to learn the relative importance of individual attributes in a shared embedding space. Furthermore, it supports effortless extension or ablation of attributes through simple zero-filling, which eliminates the need for changes to the model architecture.
Node-type Indicator ($\bm{\xi}_{i}$)
We provide an additional 5-dimensional multi-hot type tensor $\bm{\xi}_{i}\in\{0,1\}^{5}$ for each node, encoding: depot, pickup node, delivery node, node in sub-routes, node in open route. The nodes in sub-routes and open routes mean the solution $\pi$ can have sub-routes or be an open route. We introduce explicit structural roles via $\bm{\xi}_{i}$ that help the model generalize over diverse routing problems. Further ablation analysis concerning $\bm{\xi}_{i}$ can be found in Appendix K.1.
Multi-hot Problem Representation ($\bm{\lambda}_{i}$)
Through the UDR, each problem instance activates only a specific subset of the unified feature space. Thus, for any instance $\mathcal{G}_{i}$ drawn from different problems, a corresponding multi-hot problem representation $\bm{\lambda}{i}$ is obtained by indicating active (non-zero) feature slots with a value of $1$. This representation captures only the presence of features, in contrast to their raw values, and it subsequently provides conditional guidance to the position bias weight (see Equation 6) and the adaptive decoder parameters $\bm{\theta}_{dec}(\bm{\lambda})$ (see Equation 8 and Equation 9), thereby helping the model in distinguishing problem variants and refining its node selection process.
Instead of relying on a predefined problem taxonomy with a discrete set of problem tags, our proposed UDR decouples instance representation from constraint definitions and operates in a unified space characterized by both continuous and discrete features. It allows a single neural solver to address a much larger, open-ended space of VRP variants, as new problems can be seamlessly incorporated into this unified representation. Problem-specific constraints are delegated to the model-agnostic masking function $\mathcal{M}_{t}$. For example, for all capacity-constrained problems, only the remaining load is explicitly provided as input, while diverse additional constraints (e.g., time windows, duration limit, and backhaul) are enforced implicitly by masking infeasible candidate nodes during solution construction, rather than being fully enumerated in the decoder’s input. This process is further automated by an LLM-driven constraint satisfaction mechanism that generates problem-specific mask generators, which convert constraint handling from manual engineering to automated programming (see Section 4 for more details). Together, our proposed UDR enables a single model to handle various explored and even unexplored VRP variants while significantly reducing reliance on domain expertise. Detailed descriptions of $\bm{U}$ and $\bm{\lambda}$ are provided in Appendix C and Appendix D, respectively.
3.2 Model Architecture
As shown in Figure 1, we adopt AM (Kool et al., 2019) as our basic model. While UDR provides cross-problem consistency, different problems have diverse geometric priors. To efficiently learn the geometric and relational biases inherent in various problems and obtain better generalization performance, we propose two key enhancements: (1) Mixed Bias Modules (MBM) that integrate diverse geometric priors, and (2) adaptive parameter generators conditioned on the UDR. Their implementations are detailed below.
Embedding Layer
Given an instance $\mathcal{G}$, for each $\mathbf{u}_{i}=\{\bm{\rho}_{i},\ \bm{\omega}_{i},\ \bm{\xi}_{i}\}$, we project each component into $d$-dimensional embeddings through three separate linear transformations:
$\mathbf{h}_{i}^{(0)}=\bm{\rho}_{i}W_{\bm{\rho}}+\bm{\omega}_{i}W_{\bm{\omega}}+\bm{\xi}_{i}W_{\bm{\xi}},$ (1)
where $W_{\bm{\rho}}\in\mathbb{R}^{3\times d},\ W_{\bm{\omega}}\in\mathbb{R}^{6\times d},\ W_{\bm{\xi}}\in\mathbb{R}^{5\times d}$ are learnable matrices. Then we obtain a set of initial embeddings $H^{(0)}=\{\mathbf{h}_{i}^{(0)}\}_{i=0}^{n}$ for all nodes in instance $\mathcal{G}$.
Encoder
The encoder $\bm{\theta}_{enc}$ with $L$ stacked attention layers transforms $H^{(0)}$ into advanced node embeddings $H^{(L)}=\{\mathbf{h}_{i}^{(L)}\}_{i=0}^{n}$. The detailed process is provided in Appendix E.1. Unlike prior work, we replace the plain attention (Vaswani et al., 2017) with the proposed mixed bias module (MBM) that incorporates three structural bias matrices via three separate attention calculations: (1) an outgoing distance matrix $\bm{D}$; (2) an incoming distance matrix $\bm{D^{\mathrm{T}}}$; and (3) an optional relation matrix $\bm{R}=\{r_{ij}|\ \forall\ i,j\in 0,\ldots,n\}$, where $r_{ij}=0$ if a predefined relation exists (e.g., pickup–delivery pairing in PDPs) and 1 otherwise, so that smaller values remain preferable. For the $\ell$-th layer, the MBM output $\hat{\mathbf{h}}_{i}^{(\ell)}$ can be expressed as:
$\bar{\mathbf{h}}_{i}^{(0)}=\mathrm{Attention}(\mathbf{h}_{i}^{(\ell-1)},H^{(\ell-1)},f\left(\alpha,N,\bm{D}_{i}\right)),$ (2)
$\bar{\mathbf{h}}_{i}^{(1)}=\mathrm{Attention}(\mathbf{h}_{i}^{(\ell-1)},H^{(\ell-1)},f\left(\alpha,N,\bm{D}_{i}^{\mathrm{T}}\right)),\\$ (3)
$\bar{\mathbf{h}}_{i}^{(2)}=\begin{cases}\mathrm{Attention}(\mathbf{h}_{i}^{(\ell-1)},H^{(\ell-1)},f\left(\alpha,\bm{R}_{i}\right))&\text{if}\ \bm{R}\neq\emptyset\\ \bm{0}&\text{otherwise},\end{cases},$ (4)
$\hat{\mathbf{h}}_{i}^{(\ell)}=\left[\hat{\mathbf{h}}_{i}^{(0)},\hat{\mathbf{h}}_{i}^{(1)},\hat{\mathbf{h}}_{i}^{(2)}\right]W^{O},$ (5)
where $\cdot$$\cdot$] denotes the horizontal concatenation operator, $W^{O}\in\mathbb{R}^{(3\times d)\times d}$ is a learnable matrices. For $\mathrm{Attention}(\cdot)$, we adopt the Adaptation Attention Free Module (AAFM) (Zhou et al., 2024a), which boosts the perception of diverse geometric patterns through an adaptation function $f(\alpha,N,d_{ij})$ with bias weight $\alpha$ (see Appendix F for implementation details). We generate $\alpha$ via a lightweight bias network BIAS($\bm{\lambda}$) whose input is the instance’s multi-hot representation $\bm{\lambda}$:
$\mathrm{BIAS}(\bm{\lambda})=\max\left(1,\ \left(\bm{\lambda}W_{1}+\mathbf{b}_{1}\right)W_{2}+\mathbf{b}_{2}\right),$ (6)
where $W_{1}\in\mathbb{R}^{|\bm{\lambda}|\times d},\ W_{2}\in\mathbb{R}^{d\times 1},\ \mathbf{b}_{1}\in\mathbb{R}^{d},\ \mathbf{b}_{2}\in\mathbb{R}^{1}$ are learnable parameters. We set the minimum value of the generated bias to $1$, resulting in faster model convergence. Due to the relation $r_{ij}$ is scale-independent, we remove scale $N$ in adaptation function for $\bm{R}$, and if $\bm{R}$ is absent, we substitute a zero vector for $\bar{\mathbf{h}}_{i}^{(2)}$. As shown in Figure 1(b), unlike MatNet (Kwon et al., 2021), which processes $\bm{D}$ and $\bm{D^{\mathrm{T}}}$ using two parallel attention layers, we mix them into one shared attention layer and let separate heads address different priors. This head‑level incorporation not only cuts architectural redundancy but also enhances model flexibility (e.g., integrating relation matrix), enabling URS can achieve impressive performance for various VRP variants.
Decoder
Figure 2: Comparison between our URS and existing adapter-based fine-tuning frameworks.
In URS, all parameters of $\bm{\theta}_{dec}(\bm{\lambda})$ are generated based on the multi-hot problem representation $\bm{\lambda}$. Following Kwon et al. (2020), given the first and last visited node embeddings $\mathbf{h}_{\pi_{1}}^{(L)}$ and $\mathbf{h}_{\pi_{t-1}}^{(L)}$, and an optional constraint state $C_{t}\in\mathbb{R}^{1}$ (e.g., remaining load), the context embedding $\mathbf{h}_{(C)}^{t}$ can be expressed as:
$\mathbf{h}_{(C)}^{t}=\left[\mathbf{h}_{\pi_{1}}^{(L)},\ \mathbf{h}_{\pi_{t-1}}^{(L)},\ C_{t}\right]W_{Q}(\bm{\lambda}),$ (7)
where $\cdot$$\cdot$] is the horizontal concatenation operator, $W_{Q}(\bm{\lambda})\in\mathbb{R}^{(2d+1)\times d}$ is a linear projection matrix conditioned on $\bm{\lambda}$. Missing $C_{t}$ is zero-padded (see Appendix E.3 for detailed settings). The new context embedding $\hat{\mathbf{h}}_{(C)}^{t}$ is obtained via $\mathrm{Attention}(\cdot)$ (Zhou et al., 2024a) on $\mathbf{h}_{(C)}^{t}$ and $H^{(L)}$:
$\hat{\mathbf{h}}_{(C)}^{t}=\mathrm{Attention}(\mathbf{h}_{(C)}^{t},H^{(L)}W_{K}(\bm{\lambda}),H^{(L)}W_{V}(\bm{\lambda}),\mathcal{M}_{t},f(\alpha,N,d_{t,i})),$ (8)
where $W_{K}(\bm{\lambda})\in\mathbb{R}^{d\times d}$, $W_{V}(\bm{\lambda})\in\mathbb{R}^{d\times d}$ are linear projection matrices conditioned on $\bm{\lambda}$. Finally, we compute the selection probability for each feasible node $p_{\bm{\theta}}(\pi_{t}=i\mid\mathcal{G},\pi_{1:t-1})$ based on the improved compatibility with adaptation bias (Zhou et al., 2024a):
$p_{\bm{\theta}}(\pi_{t}=i\mid\mathcal{G},\pi_{1:t-1})=\mathrm{softmax}(\zeta\cdot\text{tanh}(\frac{\hat{\mathbf{h}}_{(C)}^{t}(\mathbf{h}_{i}^{(L)})^{\mathrm{T}}}{\sqrt{d}}+f(\alpha,N,d_{t,i}))+\mathcal{M}_{t})$ (9)
where $\zeta$ is the clipping parameter. Inspired by Lin et al. (2022), we use a simple MLP hypernetwork $\mathrm{WEIGHT}(\bm{\lambda})$ to generate $\{W_{Q}(\bm{\lambda}),\ W_{K}(\bm{\lambda}),\ W_{V}(\bm{\lambda})\}$ (see Appendix E.2 for implementation details). For $\alpha$ in Equation 8 and Equation 9, we generate them via $\mathrm{BIAS}(\bm{\lambda})$ in Equation 6. As shown in Figure 2, compared with adapter-based frameworks (Lin et al., 2024; Drakulic et al., 2025; Wang et al., 2025a), we provide a unified decoding entrance, which $\mathrm{WEIGHT}(\bm{\lambda})$ and $\mathrm{BIAS}(\bm{\lambda})$ maps a problem representation vector $\bm{\lambda}$ directly to all parameters of the decoder. This conditional weight generation removes problem‑specific adapters, enabling zero-shot generalization to unseen problems while maintaining solution quality.
4 LLM-driven Constraint Satisfaction Mechanism
Figure 3: The overview of LLM-driven Constraint Satisfaction Mechanism. (a) The generating pipeline of LLM-driven Constraint Satisfaction Mechanism for a general routing problem; (b) The illustration of mask generator utilization, handling the complex constraint.
While the proposed UDR enables URS to handle a wide range of routing problems, ensuring solution feasibility across diverse and potentially unseen constraints remains a major challenge. Traditional NCO approaches rely on manually designed masks $\mathcal{M}_{t}$ for each specific VRP variant Kool et al. (2019); Kwon et al. (2020); Luo et al. (2023), which becomes impractical when solving over 100 distinct variants as targeted by URS. To address this bottleneck, we propose an LLM-driven constraint satisfaction mechanism that automatically produces problem-specific mask generators. This mechanism transforms constraint handling from a manual engineering task into an automated program design task, thereby effectively supporting URS’s excellent zero-shot generalization capabilities.
As illustrated in Figure 3 (a), our constraint satisfaction mechanism operates through a systematic generate-and-validate pipeline. The process begins by prompting an LLM with both a problem-specific natural language constraint description and a standardized Python code template. The LLM then designs a set of candidate mask generators that implement the constraint logic. Each candidate is validated on a set of problem instances using a problem-specific checker to verify its reliability in producing feasible solutions. Invalid candidates are discarded, and the generation process iterates until a reliable mask generator is obtained.
Specifically, during validation, the candidate mask generator guides the stepwise solution construction. As illustrated in Figure 3 (b), at each step $t$, the mask generator takes the current state of the partial solution $\pi_{1:t-1}$ as input, which contains various information such as visited nodes, remaining load, etc. It then produces a binary mask $\mathcal{M}_{t}$ that identifies all permissible nodes. This mask is directly integrated into the URS’s selection probability computation in Equation 9 to ensure feasibility.
To efficiently steer this generate-and-test process, we formulate it as an optimization problem that maximizes the solution validity rate. The well-developed LLM-based automated algorithm design frameworks, such as EoH (Liu et al., 2024b), can be effectively applied to tackle this optimization problem. The optimization process terminates immediately once a reliable mask generator is found, which significantly reduces computational overhead compared to exhaustive search approaches. Furthermore, generated mask generators can be cached and reused for different instances sharing the same problem variant, thereby amortizing the generation cost. For more details about our LLM-driven constraint satisfaction mechanism, please see Appendix G.
5 Experiments
In this section, we comprehensively evaluate URS against classical and learning-based solvers across more than one hundred distinct VRP variants. For URS, we focus on: (1) performance on seen problems and (2) zero-shot generalization under unseen ones. To ensure comparability with prior neural routing studies, we adopt the 100-node setting as our primary evaluation scenario for URS. All experiments are conducted on a single NVIDIA GeForce RTX 4090 GPU (24GB memory).
5.1 Experimental Setup
Problem Setting
We train URS on $11$ mixed VRP variants to ensure that each feature in UDR has been seen at least once, i.e., $\mathcal{P}=\{P_{i}\}_{i=1}^{11}$. Specifically, training problems and related data generations are as follows: (1) ATSP (Kwon et al., 2021);(2) TSP (Kool et al., 2019); (3) CVRP (Kool et al., 2019); (4) ACVRP (Kwon et al., 2021; Kool et al., 2019); (5) Orienteering Problem (OP) (Kool et al., 2019); (6) PCTSP (Kool et al., 2019); (7) PDTSP (Li et al., 2021); (8) CVRPTW (Zhou et al., 2024b); (9) OCVRP (Zhou et al., 2024b); (10) CVRPB (Zhou et al., 2024b), and (11) OCVRPTW (Zhou et al., 2024b). To evaluate the zero-shot generalization of URS, we construct 96 unseen variants to further highlight the broad applicability of URS. And each test dataset consists of 1000 instances. The oracle solver for each problem can be found in Appendix D.
Model & Training Setting
We use an embedding dimension $d$ of $128$ and a feed-forward layer dimension of $512$. We set the number of attention layers in the encoder to $12$. The clipping parameter $\zeta=50$ in Equation 9 for better training convergence (Jin et al., 2023). For training, URS is trained by the REINFORCE (Williams, 1992) algorithm, and the data is generated on the fly (see Appendix E.4 for more details). We train URS for $500$ epochs with $2\text{,}000$ batches per epoch, and the batch size is $128$. At each training step, we randomly select one problem $P_{i}$ from the problem set $\mathcal{P}$. We employ the AdamW optimizer (Loshchilov & Hutter, 2017) with an initial learning rate of $10^{-4}$, which is decayed by a factor of 0.1 at the 451451st epoch. Weight decay is set to $10^{-6}$. More details about the model training are provided in Appendix H.
Baseline
We compare URS with the following methods: (1)Classical Solvers: [^1] We run LKH3 (Helsgaun, 2017) with $10000$ trials and 1 run (Kool et al., 2019). For PyVRP and OR-Tools, we run them on a single CPU core with time limits of 20s (Li et al., 2025a). PyVRP (Wouda et al., 2024), LKH3 (Helsgaun, 2017), and OR-Tools (Perron & Furnon, 2023); (2)Neural Solvers: (i) specialized solvers that support multiple VRP variants: Sym-NCO (Kim et al., 2022), BQ (Drakulic et al., 2023), LEHD (Luo et al., 2023), ICAM (Zhou et al., 2024a) and POMO (Kwon et al., 2020); (ii) general neural solvers capable of handling multiple variants [^2] Given the differences in problem variants and visible task training among the baseline models, we only compare methods with the same training setup. Thus, RF Berto et al. (2024) and CaDA (Li et al., 2025a) are not included in our experiments to ensure a fair comparison.: MTPOMO (Liu et al., 2024a), MVMoE (Zhou et al., 2024b), ReLD-MTL (Huang et al., 2025), GOAL-MTL (Drakulic et al., 2025); (iii) Solvers for less-explored variants: MatNet (Kwon et al., 2021), Heter-AM (Li et al., 2021).
Metrics and Inference
We report the average objective value (Obj.), optimality gap (Gap), and total inference time (Time) for each method. The optimality gap quantifies the discrepancy between the solutions generated by the corresponding methods and the near-optimal solutions, which are obtained using classical solvers. For most NCO baseline methods, we execute the source code provided by the authors using default settings. We compare the performance of URS to the relevant baselines for each problem as well as its single-task version (i.e., URS-STL). For symmetric instances, we report the best result with $\times 8$ instance augmentation, following Kwon et al. (2020). For asymmetric instances, we report the best result with $\times 128$ instance augmentation, following Kwon et al. (2021).
Table 2: Performance on 1K test instances of $11$ seen VRPs.
Method ATSP TSP OP PCTSP PDTSP -
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 2m 0.00% 6m 0.00% 1.5m 0.00% 1.2h 0.00% 9.8m -
Sym-NCO - 0.14% 5s 0.68% 15s 0.14% 16s - -
BQ 1.27% 1m 0.35% 7s 0.63% 7s - - -
LEHD - 0.59% 2s - - - -
POMO - 0.13% 5s - - - -
ICAM 3.05% 1m 0.15% 4s - - - -
MatNet 0.95% 6.5m - - - - -
Heter-AM - - - - 6.61% 5m -
GOAL-MTL 1.77% 1m - 1.20% 38s - - -
[…]
URS 2.26% 1.1m 0.57% 6s 0.45% 4s 1.06% 5s 4.98% 4s -
Method ACVRP CVRP CVRPTW CVRPB OCVRP OCVRPTW
Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 9.1m 0.00% 19.6m 0.00% 20.8m 0.00% 5.3m 0.00% 20.8m
Sym-NCO - 1.46% 6s - - - -
BQ - 3.24% 8s - - - -
LEHD - 3.68% 2s - - - -
[…]
ICAM - 2.04% 4s - - - -
MVMoE - 1.65% 12s 4.90% 15s 1.28% 11s 3.14% 13s 3.85% 15s
MTPOMO - 1.85% 6s 5.31% 8s 1.67% 5s 3.46% 6s 4.41% 7s
[…]
GOAL-MTL - 4.22% 48s 4.66% 42s - - -
[…]
URS 3.06% 1.4m 1.81% 6s 6.13% 8s 1.46% 6s 3.24% 6s 5.07% 8s
Table 3: Zero-shot generalization on 1K test instances of $30$ unseen VRPs. Due to page limit, complete experimental results (including all 96 unseen variants) are deferred to Appendix I.
Method CVRPBP OCVRPBP CVRPBPL OCVRPBPTW CVRPBPLTW
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
MVMoE 13.95% 14s 17.36% 14s 13.38% 14s 2.29% 15s 10.46% 16s
MTPOMO 13.71% 7s 16.94% 8s 13.44% 8s 3.26% 9s 11.35% 10s
ReLD-MTL 13.57% 9s 14.83% 9s 12.96% 10s 2.27% 11s 10.80% 12s
[…]
Method MDCVRP MDCVRPL MDOCVRPB MDOCVRPBPTW MDOCVRPBPLTW
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
MVMoE 31.90% 31s 33.42% 44s 28.21% 26s 46.18% 40s 45.23% 40s
MTPOMO 23.75% 27s 24.32% 25s 25.06% 20s 34.19% 29s 33.37% 31s
ReLD-MTL 17.37% 36s 18.48% 28s 18.50% 21s 34.08% 33s 33.83% 36s
[…]
Method ACVRPB ACVRPL ACVRPTW AOPDCVRP ACVRPBL
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
[…]
Method AOCVRPBPTW ACVRPBPLTW ACVRPBPTW AOCVRPBPLTW ACVRPBTW
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
[…]
Method AMDCVRPL AMDCVRPBPLTW AMDOCVRPBPLTW PDCVRP OPDCVRP
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
[…]
Method AMDOCVRPBPTW AMDCVRPBPTW AMDCVRP APDTSP APDCVRP
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
[…]
5.2 Performance Evaluation
Performance on Seen VRPs
The experimental results on trained VRPs with scale $100$ are reported in Table 2. We observe that among all comparable neural solvers, URS-STL achieves the best performance on 8 out of the 11 routing problems, complemented by remarkably fast inference, highlighting the strength of the URS architecture when fully specialized. Meanwhile, mixing all problems into the training of URS incurs only a minor performance drop, yet still delivers competitive performance across all problems, demonstrating that potential connections among different problems are effectively captured by URS.
Generalization to Unseen VRPs
We evaluate our URS model on zero-shot generalization across 96 unseen VRP variants, including 44 well-studied VRP variants from the NCO community and 52 newly introduced variants that have not been explored before. The detailed results are presented in Table 3 and Appendix I. On widely-studied constraints (i.e., C, O, L, TW, MD, B, BP), URS continues to deliver high-quality solutions. Notably, URS merely selects the next node from a given feasible candidates without any additional domain knowledge, making its performance especially meaningful. Benefiting from our UDR, URS can directly address previously unexplored variants while maintaining solution quality, which compounds asymmetry and PD relations. Without any fine-tuning, URS still constructs high-quality solutions for more than 90 VRP variants that are not seen in training, highlighting its strong zero-shot generalization ability.
Results on Benchmark Dataset
We further evaluate the generalization ability of URS on benchmark instances from CVRPLIB Set-X (Uchoa et al., 2017) with scale $\in[500,1000]$, and we compare URS against several representative single-task and multi-task models. As demonstrated in Table 12 (Appendix J), URS maintains its position as the best overall performance across instances of varying scales. Notably, RL-based URS surpasses the SL-based specialized model LEHD (8.678% vs. 12.836%), underscoring its practical applicability in real-world scenarios.
6 Ablation Study
Table 4: Effects of different components of URS.
Method Avg.gap
w/o $\bm{\xi}$ 3.78%
w/o $C_{t}$ 2.97%
$\{\bm{D}\}$ 6.33%
$\{\bm{D},\bm{D^{\mathrm{T}}}\}$ 4.07%
$\{\bm{D},\bm{R}\}$ 7.34%
POMO-Dec. 3.39%
ReLD-Dec. 3.87%
[…]
In this section, we conduct an ablation study across 11 seen VRP variants to validate the effect of key components of URS, mainly including: (1) without node-type indicator $\bm{\xi}$ of UDR; (2) without problem state $C_{t}$; (3) effects of different priors in MBM ; (4) effects of alternative decoder architectures. As detailed in Table 4. URS maintains its competitive performance without $C_{t}$. And its effectiveness is further improved by the inclusion of a node-type indicator $\bm{\xi}$. For MBM, the results reveal that three priors $\{\bm{D},\bm{D^{\mathrm{T}}},\bm{R}\}$ are complementary, and their combined use allows MBM to effectively learn the geometric and relational biases inherent in various problems, resulting in superior overall performance. Furthermore, replacing the decoder with POMO or ReLD confirms the strong cross-problem robustness of URS, while the integration of the decoder in ICAM (Zhou et al., 2024a) leads to a further improvement. More detailed results and analyses can be found in Appendix K.
7 Conclusion, Limitation, and Future Work
In this work, we propose a novel RL-based Unified Routing Solver (URS) for cross-problem zero-shot generalization. The core of URS is a Unified Data Representation (UDR), which broadens problem coverage by substituting problem enumeration with data unification. Its generalization capability is further enhanced by a Mixed Bias Module (MBM) and an adaptive parameter generator. Moreover, an LLM-based constraint satisfaction mechanism guarantees solution feasibility and reduces dependency on domain expertise. Extensive experiments demonstrate that URS can produce high-quality solutions for more than 100 distinct VRP variants without any fine-tuning, more than 90 of which are unseen during training.
Limitation and Future Work
While URS achieves impressive results on a wide range of VRP variants, its current training scheme does not accommodate their inherent differences in complexity. In the future, we aim to explore more effective training schemes to further improve training efficiency.
Appendix A Related Work
A.1 Neural Routing Solvers with Single-Task Learning
The NCO methods have emerged as a promising paradigm for tackling various routing problems, achieving impressive empirical performance. Early seminal work based on Pointer Networks ignited interest in learning end-to-end construction policies for TSP and CVRP (Vinyals et al., 2015; Bello et al., 2016; Nazari et al., 2018). Subsequently, the Transformer-style heavy-encoder and light-decoder auto-regressive architecture has become the dominant paradigm (Kool et al., 2019), while POMO (Kwon et al., 2020) exploits permutation symmetry to enhance exploration diversity. Many advancements within this paradigm have pushed promising results on small instances (e.g., 100-node TSPs) (Xin et al., 2020; Kim et al., 2022; Xiao et al., 2024a; b; Sun et al., 2024). However, models trained on a fixed instance scale often degrade sharply when evaluated out of scale. To mitigate the poor generalization performance, recent studies introduce auxiliary local policies (Gao et al., 2024), adaptive perception across instance scales (Zhou et al., 2024a), or different search space reduction methods (Fang et al., 2024; Zhou et al., 2025; Chen et al., 2025). In parallel, ”heavy decoder” designs that dynamically recompute embeddings at each construction step (Drakulic et al., 2023; Luo et al., 2023) demonstrate strong generalization on large-scale instances. To improve the performance on large-scale routing instances, complementary directions are reducing solving difficulty via problem decomposition (Zheng et al., 2024b; Ye et al., 2024b; Li et al., 2025b) or adopting non-autoregressive generation augmented with additional searches (e.g., Monte Carlo tree search (MCTS) and 2-opt) (Fu et al., 2021; Sun & Yang, 2023; Qiu et al., 2022; Li et al., 2023; Ye et al., 2023). Moreover, heterogeneous vehicle routing (Li et al., 2022b), asymmetric distance settings (Kwon et al., 2021), multiple VRPs (Zheng et al., 2024a), and pickup–delivery problems (Li et al., 2021; Kong et al., 2024) have each motivated tailored neural designs, which have yielded competitive results. Despite these advancements, they still train a specialized model per problem type, limiting cross-problem transfer. We instead focus on a unified neural routing solver that aims to generalize across multiple routing problems without retraining from scratch for each specification.
A.2 Neural Routing Solvers with Multi-Task Learning
To address the challenge of cross-problem generalization, growing attention has shifted to multi-task learning capable of adapting to diverse routing problems. Existing efforts can broadly fall into two categories: (1) constraint combination-based multi-task learning and (2) adapter-based fine-tuning frameworks. The former line (e.g., MTPOMO (Liu et al., 2024a)) treats VRP variants as combinations of a predefined set of attributes and trains a shared backbone, achieving promising results on up to $16$ VRP variants. Follow-up work extends this paradigm by introducing mixture-of-experts (MoE) modules (Zhou et al., 2024b), heavy decoder designs plus knowledge distillation (Zheng et al., 2025a), and constraint-aware dual-attention (Li et al., 2025a) or updated Transformer structure (Berto et al., 2024) to improve model performance. While this enables knowledge sharing across problems, its coverage is inherently bounded by the hand-specified attribute set. The second family trains a shared backbone and performs adaptation to each variant via lightweight adapters. For instance, Lin et al. (2024) pretrain a TSP backbone and adapt to new variants through adapter-based fine-tuning. Drakulic et al. (2025) integrate edge features and adopt supervised multi-task pretraining for obtaining a powerful backbone before fine-tuning on novel tasks. Wang et al. (2025a) propose a multi-armed bandit strategy to realize more efficient multi-task training. Although these reduce retraining cost, they fail to achieve zero-shot generalization. In contrast, we focus on a constructive neural routing solver supported with zero-shot capability: once trained, it can directly generate high-quality solutions for previously unseen problems without any fine-tuning.
A.3 LLM-based Algorithm Generation in VRPs
Employing LLMs to solve VRPs is a new research direction. As an early attempt, Yang et al. (2023) utilizes the LLM as a direct optimizer, prompting an LLM to directly generate solutions for TSP instances. A more predominant stream leverages LLMs as automated algorithm designers to generate algorithms for solving VRPs. These methods often integrate LLMs with a search procedure. EoH (Liu et al., 2024b) combined an LLM with Evolutionary Computation to design heuristic components for a Guided Local Search (GLS) framework. ReEvo (Ye et al., 2024a) incorporated a reflection mechanism to refine heuristics for iterative solution construction. HSEvo (Dat et al., 2025) additionally introduced a diversity mechanism for designing heuristics solving TSP and OP. And AHD-MCTS (Zheng et al., 2025b) employed Monte Carlo Tree Search (MCTS) to guide the design of heuristics for solving TSP and CVRP. Despite their promise, these methods share common limitations rooted in their reliance on the LLM’s implicit, pre-trained knowledge as the core design logic. The process of discovering effective algorithms demands extensive search procedures and is computationally expensive. Furthermore, the heuristics generated by the LLM are susceptible to overfitting to the specific instances used during the design phase, limiting their generalization capabilities. In contrast, our URS overcomes these challenges by establishing a clear division of labor: the LLM is tasked with satisfying complex constraints, while our URS model is responsible for ensuring solution quality. This approach significantly reduces the reasoning burden on the LLM, mitigating the risk of overfitting and enabling URS to effectively solve a more extensive range of VRPs.
Appendix B Setups of VRP Variants
In this paper, we consider over 100 VRP variants which may simultaneously have one or more constraints from the following categories: (1) Capacity (C); (2) Open Route (O); (3) Backhaul (B); (4) Backhaul and Priority (BP); (5) Duration Limit (L); (6) Time Windows (TW); (7) Multi-Depot (MD); (8) Prize Collecting (PC); (9) Asymmetric (A); (10) Pickup and Delivery (PD). The node coordinates are randomly sampled from the unit square when available (in symmetric geometric cases, i.e., $d_{ij}=d_{ji}$), following AM (Kool et al., 2019). Below, we provide a comprehensive description of the data generation process for each constraint.
Capacity (C)
The selection of the vehicle capacity $\mathcal{C}$ is critical, as prior work (Luo et al., 2025c) demonstrates its significant impact on the solution structure. For PDVRP variants (e.g., PDCVRP, OPDCVRP, APDCVRP, and AOPDCVRP), an excessively large $\mathcal{C}$ causes the solution to approximate that of a PDTSP instance. Conversely, an overly restrictive $\mathcal{C}$ promotes a simplistic, repetitive Pickup-Delivery sequence. To ensure structurally sound solutions, we set $\mathcal{C}$=20 for PDCVRP variants. For all other VRP variants, we adopt the convention from AM (Kool et al., 2019), setting $\mathcal{C}$=50.
Open Route (O)
The open route setting implies that vehicles are not required to return to the depot after completing their service tours. This does not alter the data generation, and it specifically affects the final cost evaluation by excluding the travel distance from the last customer back to the depot.
Backhaul (B)
The standard CVRP involves vehicles delivering goods from a depot to a set of customers with positive demands. The introduction of backhauls extends this formulation to the CVRP with Backhauls (CVRPB), where vehicles must also load goods from backhaul customers and return them to the depot, in addition to serving the standard linehaul (delivery) customers. Following the experimental setup of MTPOMO (Liu et al., 2024a), we generate customer demands by sampling from a discrete uniform distribution over the integers $\{1,\ldots,9\}$. Subsequently, 20% of these nodes are randomly designated as backhaul customers, and their corresponding demand values are negative, while the remainder are classified as linehaul customers. Notably, no precedence constraints are imposed between the servicing of linehaul and backhaul customers.To ensure the feasibility of the solutions constructed by our neural solver, we adhere to the MVMoE (Zhou et al., 2024b) framework’s policy that the first customer visited must be a linehaul node.
Backhaul and Priority (BP)
This constraint imposes a strict precedence requirement on the standard CVRPB, dictating that all linehaul customers must be serviced prior to any backhaul customers. Consequently, once a vehicle serves a backhaul customer, it is prohibited from visiting any subsequent linehaul nodes within the same route. If the first customer served is a backhaul node, the vehicle’s load is initialized to zero. This ensures the neural solver can construct a feasible solution.
Duration Limit (L)
Following MTPOMO (Liu et al., 2024a), we impose a duration limit (i.e., the maximum length of each vehicle route) of 3 in symmetric cases. Since the node coordinates are randomly sampled from the unit square, the maximum possible distance between any two nodes is $\sqrt{2}$. The setting is sufficient to ensure a vehicle can deliver goods to at least one customer and return to the depot, thus guaranteeing the feasibility of the solution. For the setting of asymmetric cases, we provide details in the description of the following asymmetric instances.
Time Windows (TW)
Consistent with the MVMoE (Zhou et al., 2024b), we set time windows of the depot to $e_{0},l_{0}]$=[0,3] and its service time to 0. The service time $s_{i}$ for each customer node is set to 0.2. For the time windows for each customer node, we generate them following the methodology outlined in MVMoE (Zhou et al., 2024b).
Multi-Depot (MD)
The generation of coordinates for the multiple depots aligns with that of other nodes, with the number of depots set to 3 (Berto et al., 2024). A vehicle is required to return to the same depot it departed from. To guarantee that each starting node is considered in this multi-depot variant, we set the number of trajectories for an instance to be the product of the number of depots and the number of customer nodes, maximizing the exploration of high-quality solutions.
Prize Collecting (PC)
Under the Prize Collecting condition, each node (excluding the depot) is assigned both a prize and a penalty. A prize is collected for each visited node, while a penalty is incurred for each unvisited node. The objective is to minimize the total travel distance plus the sum of penalties for all unvisited nodes. A key constraint is that a minimum amount of total prize must be collected before the vehicle is permitted to return to the depot and terminate its route. Consistent with the AM (Kool et al., 2019) framework, the prizes and penalties are generated as follows: $prize\sim\text{Uniform}\left(0,\frac{4}{n}\right)$, and $penalty\sim\text{Uniform}\left(0,3\cdot\frac{k_{n}}{n}\right)$, where $n=100$ and $k_{n}=4$ in our experiments.
Asymmetric (A)
We adopt the data generation method from MatNet (Kwon et al., 2021) to create an asymmetric distance matrix. Due to the small pairwise distances in the generated distance matrix, the parameters for duration limit and time windows set for symmetric cases are not applicable to the asymmetric problems. To ensure the constraints remain valid, we adjust the duration limit and depot end time to 0.6 and 1.0, respectively.
Pickup and Delivery (PD)
For PDP variants, we adhere to the methodology in  Li et al. (2021). For a problem of size $N$, node 0 is set to the depot, nodes $1$ to $N/2$ are designated as pickup nodes, while nodes from $N/2+1$ to $N$ are designated as delivery nodes. For each pickup node $v_{i}$, its corresponding delivery node is $v_{i+N/2}$. A strict precedence constraint is imposed in PDP cases, requiring each vehicle to visit a pickup node before its corresponding delivery node. For PDCVRP variants, the demand for each delivery node is randomly sampled from a discrete uniform distribution over $\{1,2,\ldots,9\}$. The demand for each pickup node is then set to the negative of its corresponding delivery node’s demand. To ensure the feasibility of the constraints, the vehicle capacity is set to $20$.
Appendix C Unified Data Representation
In this section, we provide a detailed description of our UDR $\bm{U}=\{\mathbf{u}_{i}\}_{i=0}^{n}$ (see Table 5). Each node $\mathbf{u}_{i}$ consists of three components, i.e., $\mathbf{u}_{i}=\{\bm{\rho}_{i},\ \bm{\omega}_{i},\ \bm{\xi}_{i}\}$. where $\bm{\rho}_{i},\ \bm{\omega}_{i},\ \bm{\xi}_{i}$ are a position identifier, unified attribute set, and node-type indicator, respectively. To distinguish linehaul/delivery from backhaul/pickup nodes, we assign negative demand values to backhaul/pickup nodes (positive for linehaul/delivery), making demand the only signed component, following  (Liu et al., 2024a; Zhou et al., 2024b).
Table 5: Detailed description of each component in UDR.
Attribute Name Symbol Range Description Examples
Random Identifier $\eta$ $[0,1]$ Random scalar for asymmetric problems. ATSP
Node Coordinates $\{x,y\}$ $0,1]^{2}$ 2D coordinates of the node. TSP, CVRP
Demand $\delta$ [1,1][-1,1] Node demand (++ for delivery, - for pickup). CVRP, PDCVRP
Prize $\epsilon$ $[0,1]$ Prize collected for each node. OP, PCTSP
Penalty $\mu$ $[0,1]$ Penalty incurred for each node. PCTSP
Earliest Arrival Time $e$ $[0,1]$ Earliest permissible arrival time. CVRPTW
Latest Arrival Time $l$ $[0,1]$ Latest permissible arrival time (deadline). CVRPTW
Service Time $s$ $[0,1]$ Time required for service at the node. CVRPTW
Depot - $\{0,1\}$ Indicates if the node is a depot. CVRP
Pickup Node - $\{0,1\}$ Indicates if the node is a pickup location. PDTSP
Delivery Node - $\{0,1\}$ Indicates if the node is a delivery location. CVRP, PDTSP
Node in Sub-routes - $\{0,1\}$ The solution $\pi$ can have sub-routes. CVRP
Node in Open Route - $\{0,1\}$ Vehicles need not return to the depot. OCVRP
Appendix D Multi-hot Problem Representation
Leveraging the UDR, each problem instance activates only a subset of the unified feature space. Thus, for any instance $\mathcal{G}_{i}$ from different problems, we simply mark active (non-zero) feature slots with 1 to obtain a corresponding multi-hot problem representation $\bm{\lambda}_{i}$. Different from raw feature values, it encodes presence only and provides conditional signals to the following position bias weight (see Equation 6) and adaptive decoder (see Equation 8 and Equation 9), helping the model distinguish variants and refine node selection. Here, we provide detailed representations for each problem in our experiments, and each problem is a 13-dimensional multi-hot vector.
As shown in Table 6, we train URS on eleven classic and varied routing problems to ensure that each feature in the UDR has been seen at least once, i.e., $\mathcal{P}=\{P_{i}\}_{i=1}^{11}$. Then, we evaluate generalization on unseen problems on different combinations of ten constraint attributes (see Appendix B), and the detailed representations of each unseen problem can be found in Table 7 (documented VRP variants) and Table 8 (undocumented VRP variants).
Table 6: The representations for seen routing problems in our experiments. The total number of problems is $1$. In training, each feature in the UDR has been seen at least once. Here, RI represents the random identifier, and EAT, LAT, and ST indicate the earliest arrival time, the latest arrival time, and the service time, respectively.
Problem Oracle RI Coord. Demand Prize Penalty EAT LAT ST Depot Pickup Delivery Sub-routes Open Route
ATSP LKH-3 $\checkmark$
TSP LKH-3 $\checkmark$
OP Compass $\checkmark$ $\checkmark$ $\checkmark$
PCTSP ILS $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
PDTSP LKH-3 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ACVRP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CVRP PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CVRPTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CVRPB OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OCVRP LKH-3 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OCVRPTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Table 7: The problem representations of $44$ VRP variants that have been extensively explored in current literature (Liu et al., 2024a; Berto et al., 2024). Here, RI represents the random identifier, and EAT, LAT, and ST indicate the earliest arrival time, the latest arrival time, and the service time, respectively.
Problem Oracle RI Coord. Demand Prize Penalty EAT LAT ST Depot Pickup Delivery Sub-routes Open Route
CVRPL LKH-3 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OCVRPB OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CVRPBL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CVRPLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OCVRPBTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CVRPBLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OCVRPL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CVRPBTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OCVRPBL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OCVRPLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OCVRPBLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CVRPBP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OCVRPBP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CVRPBPL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OCVRPBPTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CVRPBLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CVRPBPTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OCVRPBPL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OCVRPBLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDCVRP PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDCVRPTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOCVRP PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDCVRPL PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDCVRPB PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOCVRPTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOCVRPB PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDCVRPBL PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDCVRPLTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOCVRPBTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDCVRPBLTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOCVRPL PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDCVRPBTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOCVRPBL PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOCVRPLTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOCVRPBLTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDCVRPBP PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOCVRPBP PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDCVRPBPL PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOCVRPBPTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDCVRPBLTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDCVRPBTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOCVRPBPL PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOCVRPBLTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
SPCTSP ILS $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Table 8: The problem representations of $52$ VRP variants that have not yet been explored in current literature. Here, RI represents the random identifier, and EAT, LAT, and ST indicate the earliest arrival time, the latest arrival time, and the service time, respectively.
Problem Oracle RI Coord. Demand Prize Penalty EAT LAT ST Depot Pickup Delivery Sub-routes Open Route
ACVRPTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AOCVRP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ACVRPL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ACVRPB OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AOCVRPTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AOCVRPB OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ACVRPBL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ACVRPLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AOCVRPBTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ACVRPBLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AOCVRPL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ACVRPBTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AOCVRPBL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AOCVRPLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AOCVRPBLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ACVRPBP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AOCVRPBP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ACVRPBPL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AOCVRPBPTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ACVRPBLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ACVRPBTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AOCVRPBPL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AOCVRPBLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDCVRP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDCVRPTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDOCVRP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDCVRPL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDCVRPB OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDOCVRPTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDOCVRPB OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDCVRPBL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDCVRPLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDOCVRPBTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDCVRPBLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDOCVRPL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDCVRPBTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDOCVRPBL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDOCVRPLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDOCVRPBLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDCVRPBP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDOCVRPBP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDCVRPBPL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDOCVRPBPTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDCVRPBLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDCVRPBTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDOCVRPBPL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDOCVRPBLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
APDTSP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
PDCVRP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OPDCVRP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
APDCVRP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AOPDCVRP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Appendix E Model Architecture
E.1 Encoder
Similar to previous work (Kool et al., 2019; Kwon et al., 2020), we also use the encoder composed of $L$ stacked attention layers, where each attention layer consists of two sub-layers: an attention sub-layer and a Feed-Forward (FF) sub-layer, both of which use Instance Normalization (Ulyanov et al., 2016) and skip-connection (He et al., 2016).
Let ${H}^{(\ell-1)}=\{\mathbf{h}_{i}^{(\ell-1)}\}_{i=0}^{n}$ denote the input to the $\ell$-th attention layer for $\ell=1,\ldots,L$. The outputs for the $i$-th node are computed as follows:
$\widetilde{\mathbf{h}}_{i}^{(\ell)}=\mathrm{IN}^{(\ell)}\left(\mathbf{h}_{i}^{(\ell-1)}+\mathrm{MBM}^{(\ell)}\left(\mathbf{h}_{i}^{(\ell-1)},H^{(\ell-1)},\bm{D},\bm{D}^{\mathrm{T}},\bm{R}\right)\right),$ (10)
$\mathbf{h}_{i}^{(\ell)}=\mathrm{IN}^{(\ell)}\left(\mathbf{\widetilde{h}}_{i}^{(\ell)}+\mathrm{FF}^{(\ell)}\left(\mathbf{\widetilde{h}}_{i}^{(\ell)}\right)\right),$ (11)
where $\mathrm{IN}(\cdot)$ denotes instance normalization (Ulyanov et al., 2016), $\mathrm{MBM}$ represents the adopted mixed bias module (see Equation 5 for details), and $\mathrm{FF}(\cdot)$ in Equation 11 corresponds to a fully connected neural network with ReLU activation.
After $L$ attention layers, the final node embeddings $H^{(L)}=\{\mathbf{h}_{i}^{(L)}\}_{i=0}^{n}$ encapsulate the advanced feature representations of each node.
Notably, to ensure dimensional consistency across bias channels, we normalize $\bm{D}$ and $\bm{D^{\mathrm{T}}}$ by dividing their maximum value before passing the MBM.
E.2 Decoder
In URS, all parameters $\bm{\theta}_{dec}(\bm{\lambda})$ are adaptively generated conditioned on the multi-hot problem representation $\bm{\lambda}$ (see Appendix D for detailed problem representations). $\bm{\theta}_{dec}(\bm{\lambda})$ includes two MLP networks, which are $\mathrm{WEIGHT}(\bm{\lambda})$ and $\mathrm{BIAS}(\bm{\lambda})$. $\mathrm{WEIGHT}(\bm{\lambda})$ is used to generate linear projection matrices $W_{Q}(\bm{\lambda}),\ W_{K}(\bm{\lambda}),\ W_{V}(\bm{\lambda})]$. $\mathrm{BIAS}(\bm{\lambda})$ in Equation 6 is used to generate bias weights for $\mathrm{Attention}(\cdot)$(see Equation 8) and compatibility (see Equation 9).
For $\mathrm{WEIGHT}(\bm{\lambda})$, we use a multi-layer MLP hypernetwork conditioned on a 13-dimensional problem representation $\bm{\lambda}$ to generate problem-specific adaptive parameters. Note that we factor $W_{Q}(\bm{\lambda})\in\mathbb{R}^{(2d+1)\times d}$ into three vertically concatenated blocks $W_{\text{first}}(\bm{\lambda})\in\mathbb{R}^{d\times d},W_{\text{last}}(\bm{\lambda})\in\mathbb{R}^{d\times d}$ and $W_{C}(\bm{\lambda})\in\mathbb{R}^{1\times d}$, applied respectively to the first visited node $\mathbf{h}_{\pi_{1}}^{(L)}$, the last visited node $\mathbf{h}_{\pi_{t-1}}^{(L)}$, and a scalar state feature $C_{t}\in\mathbb{R}^{1}$ (see Appendix E.3 for more details). Thus $\mathrm{WEIGHT}(\bm{\lambda})$ outputs the linear projection matrices $\{W_{\text{first}}(\bm{\lambda}),W_{\text{last}}(\bm{\lambda}),W_{C}(\bm{\lambda}),W_{K}(\bm{\lambda}),W_{V}(\bm{\lambda})\}$.
Furthermore, we adopt the parameter compression technique of Ha et al. (2016) to control model size. Specifically, we first generate an advanced embedding matrix $\bm{H}_{\text{dec}}(\bm{\lambda})=[\bm{h}_{\text{dec}}^{i}(\bm{\lambda})|i=1,2,\ldots,5]\in\mathbb{R}^{5\times|\bm{\lambda}|}$ whose five row embeddings each parameterize one target projection matrix. It can be computed via three consecutive linear projections:
$\bm{H}_{\text{dec}}(\bm{\lambda})=((\bm{\lambda}W_{1}+\mathbf{b}_{1})W_{2}+\mathbf{b}_{2})W_{3}+b_{3},$ (12)
where $W_{1}\in\mathbb{R}^{|\lambda|\times d_{h}},\ W_{2}\in\mathbb{R}^{d_{h}\times d_{h}},\ W_{3}\in\mathbb{R}^{d_{h}\times(5\times|\bm{\lambda}|)},\ \mathbf{b}_{1}\in\mathbb{R}^{d_{h}},\ \mathbf{b}_{2}\in\mathbb{R}^{d_{h}},\mathbf{b}_{2}\in\mathbb{R}^{(5\times|\bm{\lambda}|)}$ are learnable parameters, $d_{h}$ is $256$ in this work, following Lin et al. (2022). Next, we use five different linear projections to map the new representation $\bm{h}_{\text{dec}}^{i}(\bm{\lambda})$ of each row to the corresponding decoder parameters. The computation can be expressed as:
$W_{\text{first}}(\bm{\lambda})=\bm{h}_{\text{dec}}^{1}(\bm{\lambda})W^{1}_{\textbf{hyper}},\ W_{\text{last}}(\bm{\lambda})=\bm{h}_{\text{dec}}^{2}(\bm{\lambda})W^{2}_{\textbf{hyper}},\ W_{C}(\bm{\lambda})=\bm{h}_{\text{dec}}^{3}(\bm{\lambda})W^{3}_{\textbf{hyper}},$ (13)
$W_{K}(\bm{\lambda})=\bm{h}_{\text{dec}}^{4}(\bm{\lambda})W^{4}_{\textbf{hyper}},\ W_{V}(\bm{\lambda})=\bm{h}_{\text{dec}}^{5}(\bm{\lambda})W^{5}_{\textbf{hyper}},$ (14)
where $W^{1}_{\textbf{hyper}}\in\mathbb{R}^{|\lambda|\times(d\times d)}$, $W^{2}_{\textbf{hyper}}\in\mathbb{R}^{|\lambda|\times(d\times d)}$, $W^{3}_{\textbf{hyper}}\in\mathbb{R}^{|\lambda|\times(1\times d)}$, $W^{4}_{\textbf{hyper}}\in\mathbb{R}^{|\lambda|\times(d\times d)}$, $W^{5}_{\textbf{hyper}}\in\mathbb{R}^{|\lambda|\times(d\times d)}$ are learnable parameters. In this way, we use a simple MLP hypernetwork $\mathrm{WEIGHT}(\bm{\lambda})$ to generate used linear projection matrices $\{W_{Q}(\bm{\lambda}),\ W_{K}(\bm{\lambda}),\ W_{V}(\bm{\lambda})\}$ conditioned on the multi-hot problem representation $\bm{\lambda}$, enabling the adaptive construction of high-quality solutions for different problems.
E.3 Problem State $C_{t}$
In this section, we detail the problems that require additional state signals. If a problem is not mentioned in this section, it means we will not apply any extra state to it, i.e., $C_{t}=0$.
For all problems with the capacity constraint, we explicitly impose only remaining load, while diverse additional constraints (e.g., time windows, duration limit, and backhaul) are enforced implicitly as $\mathcal{M}_{t}$ masks infeasible candidate nodes to guarantee valid solutions, instead of being fully enumerated in the input of the decoder.
For OP, we keep track of the remaining maximum length at time $t$, and $C_{t}$ represents the remaining routing length that can be moved, and we normalize it to $[0,1]$ by dividing by the maximum length (i.e., 4.0 in OP100). The setting is the same as AM (Kool et al., 2019).
For PCTSP and SPCTSP, $C_{t}$ is the remaining prize to collect, and we do not provide any information about the penalties, as this is irrelevant for the remaining decisions, following Kool et al. (2019).
For the problem state, we provide an ablation study in Appendix K.2.
E.4 Training
Following Kwon et al. (2020), we use $n$ trajectories with distinct starting nodes in training. URS is trained by the REINFORCE (Williams, 1992) algorithm with a shared baseline (Kwon et al., 2020):
$\nabla_{\theta}\mathcal{L}(\bm{\theta})=\mathbb{E}_{p(\bm{\pi}|\mathcal{G},{\bm{\theta}})}[(f(\bm{\pi}|\mathcal{G})-\frac{1}{n}\sum_{i=1}^{n}f(\bm{\pi}^{i}|\mathcal{G}))\nabla_{\bm{\theta}}\log p_{\bm{\theta}}\left(\bm{\pi}|\mathcal{G}\right)],$ (15)
$p_{\bm{\theta}}(\bm{\pi}\mid\mathcal{G})=\prod_{t=2}^{n}p_{\bm{\bm{\theta}}}(\pi_{t}\mid\mathcal{G},\pi_{1:t-1}),$ (16)
where the objective function $f(\bm{\pi}|\mathcal{G})$ represents the total reward (e.g., the negative value of tour length) of instance $\mathcal{G}$ given a specific solution $\bm{\pi}$.
Appendix F Adaptation Attention Free Module
Following Zhou et al. (2024a), we implement $\mathrm{Attention}(\cdot)$ using an adaptation attention free module (AAFM) to enhance geometric pattern recognition for routing problems. Given the input $X$, AAFM first transforms it into $Q$, $K$, and $V$ through corresponding linear projection operations:
$Q=XW^{Q},\quad K=XW^{K},\quad V=XW^{V},$ (17)
where $W^{Q}$, $W^{K}$, and $W^{V}$ are learnable matrices. The AAFM computation is then expressed as:
$\mathrm{Attention}(Q,K,V,A)=\sigma(Q)\odot\frac{\exp(A)(\exp(K)\odot V)}{\exp(A)\exp(K)},$ (18)
where $\sigma$ denotes the sigmoid function, $\odot$ represents the element-wise product, and $A=\{a_{ij}\}$ denotes the pair-wise adaptation bias. For distance matrices, the corresponding adaptation bias $f\left(\alpha,N,d_{ij}\right)=-\alpha\cdot\log_{2}{N}\cdot d_{i,j}$. Here, $N$ denotes the total number of nodes (i.e., problem size), and $\alpha$ is generated via a lightweight network $\mathrm{BIAS}(\bm{\lambda})$ in our paper (see Equation 6). Notably, for the relation matrix, we remove the scale $N$ in the calculation of adaptation bias because the relation is scale-independent.
Compared to multi-head attention (MHA) (Vaswani et al., 2017), AAFM enables the model to explicitly capture instance-specific knowledge by updating pair-wise adaptation biases while exhibiting lower computational overhead. Further details are provided in the related work section mentioned above.
Appendix G LLM-driven Constraint Satisfaction Mechanism
To operationalize the UDR framework for a vast spectrum of routing problems, we introduce an automated constraint satisfaction mechanism based on LLMs. This mechanism obviates the need for manual, problem-specific masks by automatically designing executable mask generators. This approach follows the generate-and-validate pipeline, transforming constraint specifications into reliable code.
G.1 Mask Generator Design
This process begins by prompting an LLM with a prompt that comprises two key elements: 1) a natural language task description outlining the constraint logic, and 2) a standardized Python code template that specifies the function signature, inputs, and outputs. Here is an example for generating a mask for the TSP that enforces the fundamental visit-once constraint:
Task Description: Given the current node, visited nodes, and the number of nodes, design a function to mask visited nodes.
Code Template:
1import torch
2
3def mask_visited_nodes(current_node: torch.Tensor, visited_nodes: torch.Tensor, num_nodes: int) -> torch.Tensor:
4 """
5 Design a novel algorithm to mask the visited nodes in each step.
6 The operation processes batches in parallel, so no loops over the batch dimension are allowed-inputs must use batch-first shapes.
7
8 Args:
9 current_node: ID of the current node. Shape: [batch_size, 1]
10 visited_nodes: Array of IDs of visited nodes. Shape: [batch_size, num_visited_nodes]
11 num_nodes: The number of nodes.
12
13 Return:
14 mask: Boolean mask indicating valid nodes (True = selectable, False = masked). Shape: [batch_size, num_nodes]
15 """
16 return mask
The LLM completes the function body with the prompt guidance. The candidate mask generator is an executable Python snippet implementing the described logic:
1import torch
2
3def mask_visited_nodes(current_node: torch.Tensor, visited_nodes: torch.Tensor, num_nodes: int) -> torch.Tensor:
4 """
5 Design a novel algorithm to mask the visited nodes in each step.
6 The operation processes batches in parallel, so no loops over the batch dimension are allowed-inputs must use batch-first shapes.
7
8 Args:
9 current_node: ID of the current node. Shape: [batch_size, 1]
10 visited_nodes: Array of IDs of visited nodes. Shape: [batch_size, num_visited_nodes]
11 num_nodes: The number of nodes.
12
13 Return:
14 mask: Boolean mask indicating valid nodes (True = selectable, False = masked). Shape: [batch_size, num_nodes]
15 """
16 mask = torch.ones(current_node.size(0), num_nodes, dtype=torch.bool)
17 mask.scatter_(1, visited_nodes, 0)
18 mask.scatter_(1, current_node, 0)
19 return mask
G.2 Automated Validation
Each candidate generator is then validated. We execute it within the URS solver or a simple simulated solver on a set of problem instances and use a problem-specific checker function to verify if the resulting solutions are constraint-compliant. For the TSP example, the checker ensures that every node is visited exactly once:
1import torch
2
3def checker(solutions: torch.Tensor, problem_size: int, num_instances: int) -> float:
4 """
5 Check the validity of TSP solutions that use the generated masks.
6
7 Args:
8 solutions: A batch of generated solutions to be validated. Each row represents a single solution sequence. Shape: [batch_size, problem_size].
9 problem_size: The problem size.
10
11 Returns:
12 valid_rate: The validity rate, representing the fraction of valid solutions in the batch.
13 """
14 num_instances = solutions.shape[0]
15 sorted, _ = torch.sort(solutions, dim=1)
16 valid = (sorted == torch.arange(problem_size)[None, :].expand(num_instances, problem_size)).all(1)
17 valid_rate = torch.mean(valid.float()).item()
18 return valid_rate
G.3 AHD Framework Integration
The mask generator design process can be formulated as an optimization problem aimed at maximizing the solution validity rate across test instances. This allows us to leverage sophisticated LLM-based automated algorithm design frameworks, such as EoH (Liu et al., 2024b), to improve the efficiency, especially for the complex VRP variants. The search terminates as soon as a fully reliable generator is found, making the process significantly more efficient than general optimization tasks. This LLM-driven synthesis, combined with UDR, empowers URS with true zero-shot generalization capabilities while rigorously upholding solution feasibility across an unprecedented range of routing problems.
Appendix H Training and Model Settings
The detailed information about the hyperparameter settings can be found in Table 9. The training of URS is conducted on a single NVIDIA GeForce RTX 4090 GPU (24GB memory). The training procedure requires only about $95$ hours ($\approx 4$ days) in total (500 epochs at approximately $12$ minutes each) to obtain a unified parameter set capable of addressing a broad class of routing problems.
Table 9: The hyperparameter settings of URS.
Hyperparameter Value
Optimizer AdamW
Clipping parameter $\zeta$ 50
Initial learning rate $10^{-4}$
Epochs of learning rate decay [451]
Factor of learning rate decay 0.1
Weight decay $10^{-6}$
The number of encoder layers $L$ $12$
Embedding dimension $d$ $128$
Feed forward dimension $512$
Dimension $d_{h}$ in Equation 12 $256$
Training scale $100$
Batches of each epoch $2,000$
Batch size $128$
Training Epochs $500$
Appendix I Detailed Experimental Results
To comprehensively evaluate zero-shot generalization of URS, we test its performance on a diverse set of 96 unseen problems. This set is designed to cover both established and novel challenges, including 44 well-studied VRP variants from the NCO literature and 52 newly introduced variants that have not been explored before. The detailed results are presented in Table 10 and Table 11. On widely-studied constraints (i.e., C, O, L, TW, MD, B, BP), URS continues to deliver high-quality solutions. Notably, URS merely selects the next node from a given feasible candidates without any additional domain knowledge, making its performance especially meaningful. Benefiting from our UDR, URS can directly address previously unexplored variants, which compound asymmetry and PD relations. Without any fine-tuning, URS still constructs high-quality solutions for them, strengthening the broad problem generality and deployment efficiency.
Table 10: Zero-shot generalization performance across $44$ VRP variants that have been explored in current literature (Liu et al., 2024a; Berto et al., 2024) ($N=100$).
Method CVRPBP OCVRPBP CVRPBPL OCVRPBPTW CVRPBPLTW
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
MVMoE 13.95% 14s 17.36% 14s 13.38% 14s 2.29% 15s 10.46% 16s
MTPOMO 13.71% 7s 16.94% 8s 13.44% 8s 3.26% 9s 11.35% 10s
ReLD-MTL 13.57% 9s 14.83% 9s 12.96% 10s 2.27% 11s 10.80% 12s
[…]
Method CVRPBPTW OCVRPBPL OCVRPBPLTW CVRPL CVRPBL
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 16m 0.00% 3.5h
MVMoE 10.85% 15s 16.88% 14s 1.84% 16s 0.26% 12s 1.35% 12s
MTPOMO 11.53% 9s 16.39% 8s 2.77% 10s 0.48% 6s 1.79% 6s
[…]
Method OCVRPB CVRPLTW OCVRPBTW CVRPBLTW OCVRPL
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 3.5h 0.00% 3.5h 0.00% 3.5h 0.00% 3.5h 0.00% 3.5h
MVMoE 7.09% 12s 1.47% 16s 9.95% 15s 7.33% 16s 3.15% 13s
MTPOMO 7.34% 5s 1.92% 9s 10.45% 6s 7.75% 8s 3.44% 6s
[…]
URS 9.35% 6s 2.67% 9s 13.77% 7s 9.18% 8s 3.22% 7s
Method CVRPBTW OCVRPBL OCVRPLTW OCVRPBLTW MDCVRPBP
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 3.5h 0.00% 3.5h 0.00% 3.5h 0.00% 3.5h 0.00% 6.6m
MVMoE 7.08% 15s 7.12% 12s 3.90% 15s 10.01% 15s 64.91% 33s
MTPOMO 7.41% 7s 7.34% 7s 4.37% 8s 10.50% 7s 45.66% 24s
[…]
Method MDOCVRPBP MDCVRPBPL MDOCVRPBPTW MDCVRPBPLTW MDCVRPBPTW
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
MVMoE 52.16% 33s 64.79% 36s 46.18% 38s 54.23% 42s 54.76% 39s
MTPOMO 40.15% 24s 45.84% 26s 34.19% 29s 40.26% 32s 40.32% 30s
ReLD-MTL 34.53% 27s 39.06% 29s 34.08% 34s 36.59% 37s 37.16% 35s
[…]
Method MDOCVRPBPL MDOCVRPBPLTW MDCVRP MDCVRPTW MDOCVRP
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
MVMoE 51.72% 34s 45.23% 40s 31.90% 31s 53.64% 41s 29.08% 31s
MTPOMO 39.34% 27s 33.37% 32s 23.75% 23s 36.10% 31s 27.46% 23s
[…]
Method MDCVRPL MDCVRPB MDOCVRPTW MDOCVRPB MDCVRPBL
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
MVMoE 33.42% 34s 21.70% 26s 51.57% 40s 28.21% 26s 20.35% 29s
MTPOMO 24.32% 25s 12.89% 19s 35.71% 29s 25.06% 20s 11.10% 20s
[…]
Method MDCVRPLTW MDOCVRPBTW MDCVRPBLTW MDOCVRPL MDCVRPBTW
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
MVMoE 55.55% 44s 62.47% 35s 63.18% 37s 29.33% 33s 65.90% 35s
MTPOMO 37.34% 32s 45.55% 25s 45.07% 27s 27.21% 25s 45.94% 25s
[…]
Method MDOCVRPBL MDOCVRPLTW MDOCVRPBLTW SPCTSP -
Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 1.6h -
MVMoE 28.70% 28s 51.09% 43s 62.10% 37s - -
MTPOMO 25.01% 21s 35.39% 31s 45.48% 27s - -
ReLD-MTL 18.50% 22s 33.26% 35s 43.27% 30s - -
AM - - - 2.69% 5s -
[…]
Table 11: Zero-shot generalization performance across $52$ VRP variants that have not yet been explored in the current literature ($N=100$).
Method ACVRPL ACVRPBL ACVRPLTW ACVRPBLTW AOCVRPL
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
[…]
Method AOCVRPBL AOCVRPLTW AOCVRPBLTW ACVRPBPL ACVRPBPLTW
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
[…]
Method AOCVRPBPL AOCVRPBPLTW ACVRPB ACVRPTW AOCVRP
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
[…]
Method AOCVRPTW AOCVRPB AOCVRPBTW ACVRPBTW ACVRPBP
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
[…]
Method AOCVRPBP AOCVRPBPTW ACVRPBPTW APDTSP AMDCVRPL
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
[…]
Method AMDCVRPBL AMDCVRPLTW AMDCVRPBLTW AMDOCVRPL AMDOCVRPBL
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
[…]
Method AMDOCVRPLTW AMDOCVRPBLTW AMDCVRPBPL AMDCVRPBPLTW AMDOCVRPBPL
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
[…]
Method AMDOCVRPBPLTW AMDCVRP AMDCVRPTW AMDOCVRP AMDCVRPB
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
[…]
Method AMDOCVRPTW AMDOCVRPB AMDOCVRPBTW AMDCVRPBTW AMDCVRPBP
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
[…]
Method AMDOCVRPBP AMDOCVRPBPTW AMDCVRPBPTW PDCVRP OPDCVRP
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
[…]
Method APDCVRP AOPDCVRP - - -
Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m - - -
[…]
Appendix J Results on Benchmark Dataset
Our test results on Benchmark datasets are presented in Table 12. All models are trained on instances of size $N=100$. Some results for comparative models are sourced from MVMoE (Zhou et al., 2024b) and RouteFinder (Berto et al., 2024). URS maintains its position as the best overall performance across instances of varying scales. Notably, RL-based URS surpasses the SL-based representative specialized model LEHD (8.678% vs. 12.836%), underscoring its practical applicability in real-world scenarios.
Table 12: Results on large-scale CVRPLIB instances (Set-X) (Uchoa et al., 2017) ($500\leq N\leq 1000$). All models are trained on instances of size $N=100$.
Set-X POMO LEHD MTPOMO MVMoE/4E MVMoE/4E-L RF-MVMoE RF-TE URS
Instance Opt. Obj. Gap Obj. Gap Obj. Gap Obj. Gap Obj. Gap Obj. Gap Obj. Gap Obj. Gap
[…]
Appendix K Ablation Study
In this section, we conduct a detailed ablation study and analysis to demonstrate the effectiveness and robustness of URS. Please note that, unless stated otherwise, the results presented in the ablation study reflect the best result from multiple trajectories, and we employ instance augmentation to improve performance in the ablation study. We adopt the widely used 100-node setting as our primary evaluation scenario for URS.
K.1 Effects of Node-type Indicator
Table 13: The ablation study of the node-type indicator on trained VRPs
Method ATSP TSP OP PCTSP PDTSP ACVRP CVRP CVRPTW CVRPB OCVRP OCVRPTW Avg.gap
[…]
To evaluate the effectiveness of node-type indicator $\bm{\xi}$, we conduct an ablation study about including and excluding $\bm{\xi}$. Table 13 shows that adding the node-type indicator reduces the average optimality gap from 3.78% to 2.74%, with large gains on the routing problem with relatively rich instance node roles, such as PDTSP, where the coexistence of a depot and paired pickup–delivery nodes (with inherent precedence coupling) allows the explicit node-role encoding to further boost performance. Meanwhile, URS maintains a competitive performance in addressing the problem of single-node roles.
K.2 Effects of Problem State
Table 14: Performance impact of removing the problem state feature $C_{t}$
Method ATSP TSP OP PCTSP PDTSP ACVRP CVRP CVRPTW CVRPB OCVRP OCVRPTW Avg.gap
[…]
As described in Appendix E.3, the majority of VRP variants are already solved without supplying an additional problem state $C_{t}$. To test whether the remaining use of $C_{t}$ matters, we conduct an ablation in which we eliminate it entirely, enforcing $C_{t}\equiv 0$ for all problems. The results in Table 14 show that URS maintains its competitive performance even in the absence of $C_{t}$. This indicates that explicit constraint-state features are not essential for the effectiveness of URS. Interestingly, on (O)CVRPTW instances, the model without any explicit constraint features outperforms the version that supplies only capacity information. We guess that capacity and time windows information are equally important, so enforcing only capacity creates a bias toward load considerations.
K.3 Effects of Mixed Bias Module
Table 15: Ablation of prior components in the Mixed Bias Module (MBM).
$\bm{D}$ $\bm{D^{\mathrm{T}}}$ $\bm{R}$ ATSP TSP OP PCTSP PDTSP ACVRP CVRP CVRPTW CVRPB OCVRP OCVRPTW Avg.gap
[…]
To further assess the effectiveness of MBM, we conduct an ablation study that incorporates different combinations of problem-specific priors into the MBM. Owing to its flexibility, the MBM permits imposing multiple priors on the same shared attention layer. The results in Appendix K.3 underscore the complementary roles of the three prior components: (1) The transposed distance matrix $\bm{D^{\mathrm{T}}}$ captures directional asymmetry absent from the raw distance matrix, and its inclusion markedly stabilizes optimization on asymmetric variants; (2) Removing the pickup–delivery relation matrix $\bm{R}$ degrades performance on PDTSPs, confirming that explicit relational coupling is indispensable for these instances; (3) When all priors are fused, MBM can simultaneously learn the geometric and relational biases inherent in various problems, yielding the best overall performance.
K.4 Effects of Different Decoder Architectures
Table 16: Effect of different decoder architectures on URS cross‑problem performance
Decoder ATSP TSP OP PCTSP PDTSP ACVRP CVRP CVRPTW CVRPB OCVRP OCVRPTW Avg.gap
URS-POMO 4.83% 0.74% 0.72% 1.29% 5.45% 5.09% 1.99% 6.42% 1.83% 3.62% 5.31% 3.39%
URS-ReLD 6.48% 0.90% 0.86% 1.51% 5.81% 6.51% 2.33% 6.56% 2.33% 3.62% 5.64% 3.87%
[…]
We adopt an AM (Kool et al., 2019) as our basic encoder-decoder model. Multiple decoder refinements have been proposed around AM (Kwon et al., 2020; Huang et al., 2025; Zhou et al., 2024a). To assess how decoder structure impacts multi-task performance, we additionally train two URS variants whose decoders follow POMO and ReLD, denoted URS-POMO and URS-ReLD. For stronger and fairer baselines, we apply Distance-aware Attention Reshaping (DAR) (Wang et al., 2025b) to both, and all decoder parameters are generated by a hypernetwork conditioned on the problem representation $\bm{\lambda}$ (see Appendix E.2). As shown in Table 16, URS-POMO and URS-ReLD still exhibit strong cross-problem robustness, while replacing the decoder with ICAM (Zhou et al., 2024a) yields a further overall improvement.
Appendix L Licenses for Used Resources
Table 17: List of licenses for the codes and datasets we used in this work.
Resource Type Link License
HGS-PyVRP (Wouda et al., 2024) Code https://github.com/PyVRP/PyVRP MIT License
LKH3 (Helsgaun, 2017) Code http://webhotel4.ruc.dk/~keld/research/LKH-3/ Available for academic research use
OR-Tools (Perron & Furnon, 2023) Code https://github.com/google/or-tools Apache-2.0 License
POMO (Kwon et al., 2020) Code https://github.com/yd-kwon/POMO/tree/master/NEW_py_ver MIT License
Sym-NCO (Kim et al., 2022) Code https://github.com/alstn12088/Sym-NCO Available for any non-commercial use
LEHD (Luo et al., 2023) Code https://github.com/CIAM-Group/NCO_code/tree/main/single_objective/LEHD Available for any non-commercial use
BQ (Drakulic et al., 2023) Code https://github.com/naver/bq-nco CC BY-NC-SA 4.0 license
ICAM (Zhou et al., 2024a) Code https://github.com/CIAM-Group/ICAM MIT License
MatNet (Kwon et al., 2021) Code https://github.com/yd-kwon/MatNet/tree/main MIT License
Heter-AM (Li et al., 2021) Code https://github.com/jingwenli0312/Heterogeneous-Attentions-PDP-DRL MIT License
MTPOMO (Liu et al., 2024a) Code https://github.com/FeiLiu36/MTNCO MIT License
MVMoE (Zhou et al., 2024b) Code https://github.com/RoyalSkye/Routing-MVMoE MIT License
RouteFinder (Berto et al., 2024) Code https://github.com/ai4co/routefinder MIT License
CaDA (Li et al., 2025a) Code https://github.com/CIAM-Group/CaDA MIT License
ReLD (Huang et al., 2025) Code https://github.com/ziweileonhuang/reld-nco MIT License
GOAL (Drakulic et al., 2025) Code https://github.com/naver/goal-co Available for any non-commercial use
CVRPLIB Set-X (Uchoa et al., 2017) Dataset http://vrp.galgos.inf.puc-rio.br/index.php/en/ Available for academic research use
We list the used existing codes and datasets in Table 17, and all of them are open-sourced resources for academic usage.
Appendix M LLM usage
During the preparation of this manuscript, we utilized a Large Language Model (LLM) as a writing assistance tool. The primary purpose of its use was to improve the linguistic quality of the paper. Specifically, the LLM is used to polish the writing for enhanced clarity and fluency and to identify and correct grammatical, spelling, and punctuation errors.
We confirm that all core research ideas, study design, data analysis, and conclusions are independently developed by the authors. The LLM does not contribute to the research ideation, methodology, or substantive academic content of the paper. Its function was strictly confined to language editing and proofreading to ensure the manuscript’s professionalism and readability.

### v2

@@
 However, existing neural solvers typically rely on predefined problem constraints or require per-problem fine-tuning, which substantially limits their zero-shot generalization ability to unseen VRP variants.
-To address this critical bottleneck, we propose URS, a unified neural routing solver capable of zero-shot generalization across a wide range of unseen VRPs using a single model without any fine-tuning.
-The key component of URS is the unified data representation (UDR), which replaces problem enumeration with data unification, thereby broadening the problem coverage and reducing reliance on domain expertise.
-In addition, we propose a Mixed Bias Module (MBM) to efficiently learn the geometric and relational biases inherent in various problems.
-On top of the proposed UDR, we further develop a parameter generator that adaptively adjusts the decoder and bias weights of MBM to enhance zero-shot generalization.
-Moreover, we propose an LLM-driven constraint satisfaction mechanism, which translates raw problem descriptions into executable stepwise masking functions to ensure solution feasibility.
-Extensive experiments demonstrate that URS can consistently produce high-quality solutions for more than 100 distinct VRP variants without any fine-tuning, which includes more than 90 unseen variants.
-To the best of our knowledge, URS is the first neural solver capable of handling over 100 VRP variants with a single model.
+To address this critical bottleneck, we propose URS, a unified neural routing solver that achieves zero-shot generalization across a wide range of unseen VRPs with a single model.
+We propose a unified data representation (UDR) that replaces problem enumeration with data unification, thereby broadening the problem coverage and reducing reliance on domain expertise.
+In addition, we introduce a Mixed Bias Module (MBM) during encoding to improve node embeddings, which efficiently captures multiple priors inherent to various problems.
+On top of the UDR, we develop a problem-conditioned parameter generator to further improve zero-shot generalization.
+Extensive experiments show that URS consistently produces high-quality solutions for 110 VRP variants (including 99 unseen variants) while demonstrating impressive scalability to large-scale instances with up to 7000 nodes.
+To the best of our knowledge, URS is the first neural solver to handle over 100 VRP variants with a single model.
+Our code is available at https://github.com/CIAM-Group/URS.
+Machine Learning, ICML
 1 Introduction
-The Vehicle Routing Problem (VRP) is an essential class of combinatorial optimization problems (COPs) with extensive applications in logistics and supply chain management (Tiwari & Sharma, 2023; Sar & Ghadimi, 2023).
+The Vehicle Routing Problem (VRP) is an essential class of combinatorial optimization problems (COPs) with extensive applications in logistics and supply chain management (Tiwari and Sharma, 2023; Sar and Ghadimi, 2023).
 Solving VRPs efficiently is challenging due to their NP-hard nature.
@@
 In recent decades, classical heuristic solvers (Helsgaun, 2017; Vidal, 2022) have achieved impressive performance within acceptable timeframes.
-Nevertheless, these methods require considerable domain expertise to design specialized rules for each routing problem.
+Nevertheless, these methods require considerable domain expertise to design specialist rules for each routing problem.
 Given the growing diversity of VRP variants in real-world applications, manually crafting tailored rules for every case has become impractical.
-In recent years, neural combinatorial optimization (NCO) methods have attracted substantial attention for their potential to reduce reliance on handcrafted rules while maintaining competitive solution quality (Bengio et al., 2021; Li et al., 2022a).
-This has led to the development of many high-performing neural routing solvers that can automatically learn implicit problem-specific patterns from data under different training paradigms such as supervised learning (SL) (Drakulic et al., 2023; Luo et al., 2023; 2025b), reinforcement learning (RL) (Bello et al., 2016; Kool et al., 2019; Zhou et al., 2024a), and emerging self-improved learning (SIL) (Luo et al., 2023; 2025a; Pirnay & Grimm, 2024).
-Although these solvers have demonstrated impressive performance on specific problems (Kwon et al., 2020; Kim et al., 2022; Zhou et al., 2025), they typically require architectural customization and per-problem retraining to accommodate the distinct constraints and features of different VRP variants.
-These limitations increase the overall training cost and hinder practical deployment on new problems.
+In recent years, neural combinatorial optimization (NCO) methods have attracted substantial attention for their potential to reduce reliance on handcrafted rules while maintaining competitive solution quality (Bengio et al., 2021; Ba et al., 2026).
+This has led to the development of many high-performing neural routing solvers that automatically learn implicit problem-specific rules from data under different training paradigms such as supervised learning (SL) (Drakulic et al., 2023; Luo et al., 2023, 2025b), reinforcement learning (RL) (Kool et al., 2019; Zhou et al., 2026), and self-improved learning (SIL) (Luo et al., 2025a; Pirnay and Grimm, 2024).
+Although these solvers have demonstrated impressive performance on specific problems (Huang et al., 2025; Zhou et al., 2025), they typically require architectural customization and per-problem retraining to accommodate the distinct constraints and features of different VRP variants, thereby increasing overall training costs and hindering practical deployment to new problems.
 Table 1: Comparison between our URS and existing neural solvers with multi-task learning.
-Neural Solver Training Paradigm Routing Problems (#\#) Zero-shot Generalization Remarks
-MTPOMO (Liu et al., 2024a) RL 16 $\checkmark$ Constraint Combination
-MVMoE (Zhou et al., 2024b) RL 16 $\checkmark$ Constraint Combination
-RouteFinder (Berto et al., 2024) RL 48 $\checkmark$ Constraint Combination
-CaDA (Li et al., 2025a) RL 16 $\checkmark$ Constraint Combination
-MTL-KD (Zheng et al., 2025a) RL+KD 16 $\checkmark$ Constraint Combination
-TSP-FT (Lin et al., 2024) RL 3 $\times$ Adapter-based Fine-tuning
-MTL-MAB (Wang et al., 2025a) RL 3 $\times$ Adapter-based Fine-tuning
-GOAL (Drakulic et al., 2025) SL 9 $\times$ Adapter-based Fine-tuning
-[…]
+Note that ”#\#VRP Variants” and ”Generalizable Scale” refer to the total number of tested problems and the maximum scale reported in the original papers, respectively.
+Multi-task Neural Routing Solver Training #\#VRP Variants Symmetric Asymmetric Pickup and Delivery Zero-shot Training Generalizable Remarks
+Paradigm VRPs VRPs Problems Generalization Scale Scale
+MTPOMO (Liu et al., 2024) RL 16 $\checkmark$ $\times$ $\times$ $\checkmark$ $100$ $1000$ Constraint Combination
+MVMoE (Zhou et al., 2024) RL 16 $\checkmark$ $\times$ $\times$ $\checkmark$ $100$ $1000$ Constraint Combination
+RouteFinder (Berto et al., 2025) RL 48 $\checkmark$ $\times$ $\times$ $\checkmark$ $100$ $1000$ Constraint Combination
+CaDA (Li et al., 2025a) RL 16 $\checkmark$ $\times$ $\times$ $\checkmark$ $100$ $200$ Constraint Combination
+MTL-KD (Zheng et al., 2025) RL+KD 16 $\checkmark$ $\times$ $\times$ $\checkmark$ $100$ $1000$ Constraint Combination
+TSP-FT (Lin et al., 2024) RL 4 $\checkmark$ $\times$ $\times$ $\times$ $100$ $1000$ Adapter-based Fine-tuning
+MTL-MAB (Wang et al., 2025a) RL 3 $\checkmark$ $\times$ $\times$ $\times$ $100$ $1000$ Adapter-based Fine-tuning
+GOAL (Drakulic et al., 2025) SL 10 $\checkmark$ $\checkmark$ $\times$ $\times$ $100$ $1000$ Adapter-based Fine-tuning
+URS (Ours) RL 110 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 100 7000 Unified Data Representation
 To address the challenge of cross-problem generalization, growing attention has been directed toward multi-task learning capable of handling diverse routing problems.
-As summarized in Table 1, existing efforts largely fall into two categories: 1) constraint combination-based multi-task learning and 2) adapter-based fine-tuning frameworks.
-In the first approach, VRP variants are treated as different combinations of constraint attributes, and a unified model is trained across these combinations to enable knowledge sharing for problems with seen constraints (Liu et al., 2024a; Zhou et al., 2024b; Li et al., 2025a; Berto et al., 2024; Zheng et al., 2025a).
-The second approach builds a shared model backbone for all problems and incorporates specific input/output adapters for each problem, thereby reducing the retraining costs (Lin et al., 2024; Drakulic et al., 2025; Wang et al., 2025a).
-Despite the above advancements, existing methods still fall short in zero-shot generalization.
-The problem coverage of constraint combination-based approaches is inherently bounded by their manually specified constraint sets, while adapter-based approaches require fine-tuning and thus cannot perform zero-shot generalization.
-Moreover, both strategies fundamentally rely on the explicit enumeration of problems through a predefined set of problem tags.
-This practice is problematic because the constraint space of routing problems is open-ended and compositional, making any fixed tagging scheme fundamentally limited.
-More critically, creating and maintaining such a problem taxonomy requires considerable domain expertise, which is precisely what NCO always aims to avoid.
+As summarized in Table 1, existing efforts largely fall into two categories: 1) constraint combination-based multi-task learning and 2) adapter-based fine-tuning.
+In the first approach, VRP variants are treated as different combinations of constraint attributes, and a unified model is trained across these combinations to enable knowledge sharing for problems with seen constraints (Liu et al., 2024; Zhou et al., 2024; Li et al., 2025a; Berto et al., 2025; Zheng et al., 2025).
+The second approach builds a shared model backbone for all problems and incorporates problem-specific input/output adapters, thereby reducing retraining costs (Lin et al., 2024; Drakulic et al., 2025; Wang et al., 2025a).
+Despite the above advancements, existing methods still fall short in cross-problem zero-shot generalization.
+The problem coverage of constraint combination-based approaches is inherently bounded by their manually specified constraint sets, whereas adapter-based approaches require additional fine-tuning and thus cannot perform zero-shot generalization.
+Furthermore, both strategies fundamentally rely on the explicit problem enumeration through a predefined set of problem tags.
+This practice is challenging because the constraint space of VRP variants is open-ended and compositional.
+More critically, creating and maintaining such a problem taxonomy requires considerable domain expertise, which NCO always aims to avoid.
 A detailed review of related work on single-task and multi-task learning is provided in Appendix A.
-In this paper, we propose a powerful Unified Neural Routing Solver (URS) to significantly improve the cross-problem zero-shot generalization ability for NCO methods.
-Unlike existing approaches, we introduce a unified data representation (UDR) that replaces problem enumeration with data unification, thereby broadening problem coverage and reducing reliance on domain expertise.
-In addition, we propose a Mixed Bias Module (MBM) to better learn the geometric and relational biases inherent in various problems.
-Based on UDR, we further develop a parameter generator that adaptively adjusts decoder and bias weights of MBM to improve zero-shot generalization.
-Furthermore, to handle complex and unseen routing problems, we introduce a constraint satisfaction mechanism that leverages a large language model (LLM) to translate raw problem constraint descriptions into executable node masking functions for each construction step, thereby ensuring our URS generates feasible solutions for various problems.
-We evaluate URS on more than 100 VRP variants, including more than 90 unseen variants.
-The results demonstrate that URS not only obtains competitive performance against specialized neural solvers on seen variants but also exhibits strong zero-shot generalization capabilities on a wide range of unseen problems.
-To the best of our knowledge, URS is the first neural solver capable of efficiently solving over 100 distinct VRP variants using a single model without any retraining or fine-tuning.
+In this paper, we propose a powerful Unified Neural Routing Solver (URS) to improve the cross-problem zero-shot generalization ability for NCO methods.
+Our contributions can be summarized as follows: (1) We propose a unified data representation (UDR) that replaces existing problem enumeration with data unification, significantly broadening problem coverage while mitigating reliance on domain-specific expertise; (2) We introduce a Mixed Bias Module (MBM) during encoding to improve node embeddings, which efficiently captures multiple priors inherent to various problems while reducing architectural redundancy; (3) On top of the UDR, we develop a problem-conditioned parameter generator to further improve zero-shot generalization; (4) Extensive experiments on 110 VRP variants show that URS not only achieves competitive performance against specialist neural solvers on seen variants but also exhibits strong zero-shot generalization across 99 unseen variants.
+Notably, URS demonstrates impressive scalability to large-scale instances with up to 7000 nodes.
+To the best of our knowledge, URS is the first neural solver to efficiently solve over 100 VRP variants with a single model, without retraining or fine-tuning.
 2 Preliminaries
 In this section, we first introduce the definition of VRP, then provide an overview of recent constructive neural solvers for solution generation (Kool et al., 2019; Kwon et al., 2020).
-2.1 Vehicle Routing Problems
+Vehicle Routing Problems
 Consider a VRP instance with an optional depot indexed by 0 and $n$ customers indexed by $\{1,2,\ldots,n\}$, such as in the Capacitated VRP (CVRP).
@@
 In VRP, each node $v_{i}\in\mathcal{V}$ includes node coordinates $\{x_{i},y_{i}\}$ when available and problem-specific attributes (e.g., demands in CVRP).
-A feasible solution is a tour $\bm{\pi}=(\pi_{1},\pi_{2},\ldots,\pi_{m})$ (i.e., a finite sequence), which is a permutation of $m$ nodes.
-Let $\Omega$ denote the feasible sequence set that satisfies the constraints, given an objective function $f(\bm{\pi}|\mathcal{G})$, we aim to search for a sequence $\bm{\pi}^{*}$ with the maximum objective, i.e., $\bm{\pi}^{*}=\arg\max_{\bm{\pi}\in\Omega}f(\pi|\mathcal{G})$.
-In most routing problems, the objective can be represented as the negative value of the total distance of a feasible sequence $\bm{\pi}$.
-In this paper, we consider over 100 VRP variants which may simultaneously have one or more constraints from the following categories: (1) Capacity (C); (2) Open Route (O); (3) Backhaul (B); (4) Backhaul and Priority (BP); (5) Duration Limit (L); (6) Time Windows (TW); (7) Multi-Depot (MD); (8) Prize Collecting (PC); (9) Asymmetric (A); (10) Pickup and Delivery (PD).
+Formally, a feasible solution is denoted by a node permutation $\bm{\pi}=(\pi_{1},\pi_{2},\ldots,\pi_{m})$ (Toth and Vigo, 2002) (i.e., a finite sequence) that represents a set of valid routes.
+Each route is traversed by a single vehicle originating and terminating at its own depot or starting node, while strictly satisfying all constraints.
+Let $\Omega$ denote the feasible sequence set, given an objective function $f(\bm{\pi}|\mathcal{G})$, we aim to search for a sequence $\bm{\pi}^{*}$ with the maximum objective, i.e., $\bm{\pi}^{*}=\arg\max_{\bm{\pi}\in\Omega}f(\pi|\mathcal{G})$.
+In most VRPs, $f(\bm{\pi}|\mathcal{G})$ can be defined as the negative value of the total distance of $\bm{\pi}$.
+In this paper, we consider 110 VRP variants which may simultaneously have one or more constraints from the following categories: (1) Capacity (C); (2) Open Route (O); (3) Backhaul (B); (4) Backhaul and Priority (BP); (5) Duration Limit (L); (6) Time Windows (TW); (7) Multi-Depot (MD); (8) Prize Collecting (PC); (9) Asymmetric (A); and (10) Pickup and Delivery (PD).
 Detailed definitions of these constraints are provided in Appendix B.
-2.2 Constructive Neural Routing Solver
-Most constructive NCO methods employ an encoder-decoder architecture (Luo et al., 2023; Kwon et al., 2020).
+Figure 1: The pipeline of our URS for solving 110 VRP variants using a single model without any fine-tuning, illustrated on APDTSP.
+Constructive Neural Routing Solver
+Most constructive NCO methods employ an encoder-decoder architecture for solution construction (Luo et al., 2023; Kwon et al., 2020).
 Let learnable model parameters be denoted as $\bm{\theta}=\{\bm{\theta}_{enc},\bm{\theta}_{dec}\}$.
-Without loss of generality, we present the prevailing construction pipeline using AM (Kool et al., 2019).
+Without loss of generality, we present the prevailing autoregressive construction pipeline using AM (Kool et al., 2019).
 Given an instance $\mathcal{G}$, raw node features are first mapped by a linear projection to initial embeddings $H^{(0)}=\{\mathbf{h}_{i}^{(0)}\}_{i=0}^{n}$.
 They are then processed by an encoder $\bm{\theta}_{enc}$ with $L$ attention layers to produce refined node embeddings $H^{(L)}=\{\mathbf{h}_{i}^{(L)}\}_{i=0}^{n}$.
-At each decoding step $t$, the decoder $\bm{\theta}_{dec}$ sequentially selects a node to append to the current partial solution $\pi_{1:t-1}=(\pi_{1},\pi_{2},...,\pi_{t-1})$, where $\pi_{1},\pi_{t-1}\in\mathcal{V}$ are the first and last visited node, respectively.
+At each decoding step $t$, the decoder $\bm{\theta}_{dec}$ sequentially selects a node to append to the current partial solution $\pi_{1:t-1}=(\pi_{1},\pi_{2},\ldots,\pi_{t-1})$, where $\pi_{1},\pi_{t-1}\in\mathcal{V}$ are the first and last visited node, respectively.
 This sequential process continues until a complete solution is formed.
 To guarantee the feasibility of generated solutions, a masking function $\mathcal{M}_{t}$ sets the selection probabilities of the following nodes to -$\infty$ during the construction process: nodes that (1) are visited or (2) violate problem-specific constraints (e.g., capacity).
-3 The Proposed Model: Unified Neural Routing Solver
-In this section, we propose a powerful Unified Neural Routing Solver (URS), which can significantly improve cross-problem zero-shot generalization for NCO methods.
-The framework is illustrated in Figure 1, and its key components are elaborated below.
-Figure 1: The overview of our URS with a 44‑node APDTSP instance as an example.
-(a) The solving pipeline of URS; (b) The proposed mixed bias module that incorporates multiple geometric priors, enabling it to efficiently learn the geometric and relational biases inherent in various problems.
+3 Methodology
+In this section, as illustrated in Figure 1, we propose a powerful Unified Neural Routing Solver (URS), which significantly improves cross-problem zero-shot generalization for NCO methods.
+Its key components are elaborated below.
 3.1 Unified Data Representation
-Existing multi-task neural solvers (Liu et al., 2024a; Zhou et al., 2024b; Drakulic et al., 2025; Wang et al., 2025a) fundamentally rely on the explicit problem enumeration through a predefined set of problem tags, which is challenging because the constraint space of routing problems is open-ended and compositional, making any fixed tagging scheme fundamentally limited.
-More critically, creating and maintaining such a problem taxonomy requires considerable domain expertise, which is precisely what NCO always aims to avoid.
-From a data perspective, despite the substantial diversity in constraints among routing problems, their instances share a common instance representation, e.g., all input instances under the current constraint combinations (Liu et al., 2024a; Zhou et al., 2024b) can be formulated as instances of either CVRP or CVRPTW.
-Leveraging this commonality, we propose a Unified Data Representation (UDR) to replace problem enumeration with data unification, thereby broadening problem coverage and reducing reliance on domain expertise.
-For an instance $\mathcal{G}$, we define the unified representation $\bm{U}=\{\mathbf{u}_{i}\}_{i=0}^{n}$, each $\mathbf{u}_{i}=\{\bm{\rho}_{i},\ \bm{\omega}_{i},\ \bm{\xi}_{i}\}$, where $\bm{\rho}_{i},\ \bm{\omega}_{i},\ \bm{\xi}_{i}$ are a position identifier, unified attribute set, and node-type indicator, respectively.
-Position Identifier ($\bm{\rho}_{i}$)
+Existing multi-task neural solvers (Zhou et al., 2024; Drakulic et al., 2025) fundamentally rely on explicit problem enumeration via a predefined set of problem tags, which is challenging because the constraint space of VRPs is open-ended and compositional.
+More critically, creating and maintaining such a problem taxonomy requires considerable domain expertise, which NCO always aims to avoid.
+From a data perspective, despite the substantial diversity in constraints, their instances share a common structural representation.
+For existing constraint combination-based methods (Liu et al., 2024; Zhou et al., 2024), their evaluated problems can be formulated as CVRPTW instances.
+Leveraging this structural commonality, we propose a Unified Data Representation (UDR) $\bm{U}=\{\mathbf{u}_{i}\}_{i=0}^{n}$.
+By replacing explicit problem enumeration with data unification, our UDR broadens problem coverage and reduces reliance on domain expertise.
+Notably, the inputs to existing multi-task neural solvers based on either problem-specific adapters (Drakulic et al., 2025) or constraint combinations (Zhou et al., 2024) can be viewed as special subsets of our proposed UDR.
+Instead of relying on a predefined problem taxonomy with a discrete set of problem tags, UDR decouples instance representation from constraint definitions and operates in a unified space characterized by both continuous and discrete features.
+It allows a single neural solver to address a much larger, open-ended space of VRP variants, as new problems can be seamlessly incorporated into this unified representation.
+Problem-specific constraints are delegated to the model-agnostic masking function $\mathcal{M}_{t}$.
+For example, for all CVRP variants, only the remaining load is explicitly provided as input, while diverse additional constraints (e.g., time windows) are enforced implicitly by masking infeasible candidate nodes during solution construction, rather than being fully enumerated in the decoder’s input.
+Thus, UDR enables a single model to handle various VRP variants while significantly reducing reliance on domain expertise.
+Specifically, for an instance $\mathcal{G}$, each $\mathbf{u}_{i}=\{\bm{\rho}_{i},\ \bm{\omega}_{i},\ \bm{\xi}_{i}\}$, where $\bm{\rho}_{i},\ \bm{\omega}_{i},\ \bm{\xi}_{i}$ are a positional identifier, unified attribute set, and node-type indicator, respectively.
+Positional Identifier ($\bm{\rho}_{i}$)
 We define a unified positional identifier for each node as $\bm{\rho}_{i}=\{\eta_{i},x_{i},y_{i}\}\in[0,1]^{3}$.
-The $\eta_{i}\sim\mathrm{Uniform}(0,1)\in\mathbb{R}^{1}$ is an additional sampled scalar, which is used to address asymmetric problems, following  Drakulic et al.
+The $\eta_{i}\sim\mathrm{Uniform}(0,1)\in\mathbb{R}^{1}$ is a sampled scalar, which is used to address asymmetric problems, following  Drakulic et al.
 (2023).
@@
 The unification empowers the model to learn the relative importance of individual attributes in a shared embedding space.
-Furthermore, it supports effortless extension or ablation of attributes through simple zero-filling, which eliminates the need for changes to the model architecture.
+It also supports effortless attribute extension or ablation via simple zero-filling, eliminating the need to change the architecture.
 Node-type Indicator ($\bm{\xi}_{i}$)
-We provide an additional 5-dimensional multi-hot type tensor $\bm{\xi}_{i}\in\{0,1\}^{5}$ for each node, encoding: depot, pickup node, delivery node, node in sub-routes, node in open route.
+We provide a 5-dimensional binary vector $\bm{\xi}_{i}\in\{0,1\}^{5}$ for each node, encoding: depot, pickup node, delivery node, node in sub-routes, node in open route.
 The nodes in sub-routes and open routes mean the solution $\pi$ can have sub-routes or be an open route.
-We introduce explicit structural roles via $\bm{\xi}_{i}$ that help the model generalize over diverse routing problems.
-Further ablation analysis concerning $\bm{\xi}_{i}$ can be found in Appendix K.1.
+We introduce explicit structural roles to enable the model to generalize over diverse variants.
+Further ablation analysis concerning $\bm{\xi}_{i}$ is provided in Appendix K.1.
 Multi-hot Problem Representation ($\bm{\lambda}_{i}$)
 Through the UDR, each problem instance activates only a specific subset of the unified feature space.
-Thus, for any instance $\mathcal{G}_{i}$ drawn from different problems, a corresponding multi-hot problem representation $\bm{\lambda}{i}$ is obtained by indicating active (non-zero) feature slots with a value of $1$.
+For any instance $\mathcal{G}_{i}$ drawn from different problems, a corresponding multi-hot problem representation $\bm{\lambda}_{i}$ is obtained by indicating active (non-zero) feature slots with a value of $1$.
 This representation captures only the presence of features, in contrast to their raw values, and it subsequently provides conditional guidance to the position bias weight (see Equation 6) and the adaptive decoder parameters $\bm{\theta}_{dec}(\bm{\lambda})$ (see Equation 9 and Equation 10), thereby helping the model in distinguishing problem variants and refining its node selection process.
-Instead of relying on a predefined problem taxonomy with a discrete set of problem tags, our proposed UDR decouples instance representation from constraint definitions and operates in a unified space characterized by both continuous and discrete features.
-It allows a single neural solver to address a much larger, open-ended space of VRP variants, as new problems can be seamlessly incorporated into this unified representation.
-Problem-specific constraints are delegated to the model-agnostic masking function $\mathcal{M}_{t}$.
-For example, for all capacity-constrained problems, only the remaining load is explicitly provided as input, while diverse additional constraints (e.g., time windows, duration limit, and backhaul) are enforced implicitly by masking infeasible candidate nodes during solution construction, rather than being fully enumerated in the decoder’s input.
-This process is further automated by an LLM-driven constraint satisfaction mechanism that generates problem-specific mask generators, which convert constraint handling from manual engineering to automated programming (see Section 4 for more details).
-Together, our proposed UDR enables a single model to handle various explored and even unexplored VRP variants while significantly reducing reliance on domain expertise.
-Detailed descriptions of $\bm{U}$ and $\bm{\lambda}$ are provided in Appendix C and Appendix D, respectively.
+A more detailed description of $\bm{U}$ and $\bm{\lambda}$ is provided in Appendix C and Appendix D, respectively.
 3.2 Model Architecture
@@
 While UDR provides cross-problem consistency, different problems have diverse geometric priors.
-To efficiently learn the geometric and relational biases inherent in various problems and obtain better generalization performance, we propose two key enhancements: (1) Mixed Bias Modules (MBM) that integrate diverse geometric priors, and (2) adaptive parameter generators conditioned on the UDR.
+To efficiently learn the multiple priors inherent in various problems and obtain better cross-problem zero-shot generalization, we propose two key enhancements: (1) a Mixed Bias Module (MBM) that captures multiple priors inherent to various problems, and (2) problem-conditioned parameter generators conditioned on the UDR.
 Their implementations are detailed below.
 Embedding Layer
-Given an instance $\mathcal{G}$, for each $\mathbf{u}_{i}=\{\bm{\rho}_{i},\ \bm{\omega}_{i},\ \bm{\xi}_{i}\}$, we project each component into $d$-dimensional embeddings through three separate linear transformations:
+Given an instance $\mathcal{G}$, for each $\mathbf{u}_{i}=\{\bm{\rho}_{i},\ \bm{\omega}_{i},\ \bm{\xi}_{i}\}$, we project it into $d$-dimensional embeddings through three separate linear transformations:
 $\mathbf{h}_{i}^{(0)}=\bm{\rho}_{i}W_{\bm{\rho}}+\bm{\omega}_{i}W_{\bm{\omega}}+\bm{\xi}_{i}W_{\bm{\xi}},$ (1)
@@
 Then we obtain a set of initial embeddings $H^{(0)}=\{\mathbf{h}_{i}^{(0)}\}_{i=0}^{n}$ for all nodes in instance $\mathcal{G}$.
-Encoder
-The encoder $\bm{\theta}_{enc}$ with $L$ stacked attention layers transforms $H^{(0)}$ into advanced node embeddings $H^{(L)}=\{\mathbf{h}_{i}^{(L)}\}_{i=0}^{n}$.
-The detailed process is provided in Appendix E.1.
-Unlike prior work, we replace the plain attention (Vaswani et al., 2017) with the proposed mixed bias module (MBM) that incorporates three structural bias matrices via three separate attention calculations: (1) an outgoing distance matrix $\bm{D}$; (2) an incoming distance matrix $\bm{D^{\mathrm{T}}}$; and (3) an optional relation matrix $\bm{R}=\{r_{ij}|\ \forall\ i,j\in 0,\ldots,n\}$, where $r_{ij}=0$ if a predefined relation exists (e.g., pickup–delivery pairing in PDPs) and 1 otherwise, so that smaller values remain preferable.
+This embedding is the initial input of encoder $\bm{\theta}_{enc}$.
+After passing through $L$ stacked attention layers, $H^{(0)}$ is transformed into advanced node embeddings $H^{(L)}=\{\mathbf{h}_{i}^{(L)}\}_{i=0}^{n}$.
+The detailed encoding process is provided in Appendix E.1.
+Figure 2: The structure of the proposed Mixed Bias Module, explained on a 44-node APDTSP instance.
+We replace plain attention with MBM during encoding to improve node embeddings, which effectively captures multiple priors inherent to various problems while reducing architectural redundancy.
+Mixed Bias Module
+Given the distinct priors across VRP variants, efficiently capturing the intrinsic geometric bias is pivotal for enhancing the cross-problem generalization of RL-based multi-task neural solvers.
+Current leading RL-based models generally adopt a heavy encoder and light decoder architecture, where the quality of the encoder-generated node embeddings plays a key role in overall performance.
+While prior studies like MatNet (Kwon et al., 2021) have attempted to handle asymmetry by integrating directional distance biases via parallel layers, such designs incur high model complexity and lack universality.
+Consequently, effectively encoding multiple inherent biases, such as symmetric and asymmetric distances, and relational dependencies, remains challenging, hindering the acquisition of high-quality solutions across a wide range of problems.
+To address the limitation, we propose the mixed bias module (MBM) to replace the vanilla attention mechanism (Vaswani et al., 2017), which enhances cross-problem generalization while reducing architectural redundancy.
+As shown in Figure 2, MBM can efficiently capture three problem-specific prior matrices via three separate attention calculations: (1) an outgoing distance matrix $\bm{D}$; (2) an incoming distance matrix $\bm{D}^{\mathrm{T}}$; and (3) an optional relation matrix $\bm{R}=\{r_{ij}|\ \forall\ i,j\in 0,\ldots,n\}$, where $r_{ij}=0$ if a predefined relation exists (e.g., pickup-delivery pairing) and 1 otherwise.
+This ensures that smaller values are preferred, aligning with the distance metric, where smaller values correspond to a larger selection bias.
 For the $\ell$-th layer, the MBM output $\hat{\mathbf{h}}_{i}^{(\ell)}$ can be expressed as:
@@
 $\bar{\mathbf{h}}_{i}^{(1)}=\mathrm{Attention}(\mathbf{h}_{i}^{(\ell-1)},H^{(\ell-1)},f\left(\alpha,N,\bm{D}_{i}^{\mathrm{T}}\right)),\\$ (3)
-$\bar{\mathbf{h}}_{i}^{(2)}=\begin{cases}\mathrm{Attention}(\mathbf{h}_{i}^{(\ell-1)},H^{(\ell-1)},f\left(\alpha,\bm{R}_{i}\right))&\text{if}\ \bm{R}\neq\emptyset\\ \bm{0}&\text{otherwise},\end{cases},$ (4)
-$\hat{\mathbf{h}}_{i}^{(\ell)}=\left[\hat{\mathbf{h}}_{i}^{(0)},\hat{\mathbf{h}}_{i}^{(1)},\hat{\mathbf{h}}_{i}^{(2)}\right]W^{O},$ (5)
-where $\cdot$$\cdot$] denotes the horizontal concatenation operator, $W^{O}\in\mathbb{R}^{(3\times d)\times d}$ is a learnable matrices.
-For $\mathrm{Attention}(\cdot)$, we adopt the Adaptation Attention Free Module (AAFM) (Zhou et al., 2024a), which boosts the perception of diverse geometric patterns through an adaptation function $f(\alpha,N,d_{ij})$ with bias weight $\alpha$ (see Appendix F for implementation details).
-We generate $\alpha$ via a lightweight bias network BIAS($\bm{\lambda}$) whose input is the instance’s multi-hot representation $\bm{\lambda}$:
+$\bar{\mathbf{h}}_{i}^{(2)}=\begin{cases}\mathrm{Attention}(\mathbf{h}_{i}^{(\ell-1)},H^{(\ell-1)},f\left(\alpha,\bm{R}_{i}\right))&\text{if}\ \bm{R}\neq\emptyset\\ \bm{0}&\text{otherwise},\end{cases}$ (4)
+$\hat{\mathbf{h}}_{i}^{(\ell)}=\left[\bar{\mathbf{h}}_{i}^{(0)},\bar{\mathbf{h}}_{i}^{(1)},\bar{\mathbf{h}}_{i}^{(2)}\right]W^{O},$ (5)
+where $\cdot$$\cdot$] denotes the horizontal concatenation operator, $W^{O}\in\mathbb{R}^{(3\times d)\times d}$ is a learnable matrix.
+For $\mathrm{Attention}(\cdot)$, we adopt the attention mechanism introduced by Zhou et al.
+(2026), which boosts the perception of diverse geometric patterns through an adaptation function $f(\alpha,N,d_{ij})$ with bias weight $\alpha$ (see Appendix F for implementation details).
+Note that since the relation $r_{ij}$ is scale-independent, we remove the scale $N$ in the adaptation function for $\bm{R}$.
+If $\bm{R}$ is absent, we substitute a zero vector for $\bar{\mathbf{h}}_{i}^{(2)}$.
+For a comparison between our MBM and existing attentions (Vaswani et al., 2017; Kwon et al., 2021), please see Appendix G.
+Problem-Conditioned Parameter Generation
+Using a single static set of parameters to solve 110 VRP variants is a significant challenge.
+To mitigate this, we introduce a problem-conditioned mechanism based on problem representation $\bm{\lambda}$, which consists of two components: (1) $\mathrm{BIAS}(\bm{\lambda})$ adjusts the degree to which different priors are integrated into the attention; and (2) $\mathrm{WEIGHT}(\bm{\lambda})$ generates decoder parameters for each problem.
+As shown in Figure 3, unlike existing adapter-based fine-tuning methods (Lin et al., 2024; Drakulic et al., 2025), we provide a unified decoding entrance and enable URS to adaptively generate decoder parameters for each problem.
+This conditional weight generation removes problem-specific adapters, enabling zero-shot generalization to unseen problems while maintaining solution quality.
+For MBM in Equation 5, we generate $\alpha$ via a lightweight bias network $\mathrm{BIAS}(\bm{\lambda})$:
 $\mathrm{BIAS}(\bm{\lambda})=\max\left(1,\ \left(\bm{\lambda}W_{1}+\mathbf{b}_{1}\right)W_{2}+\mathbf{b}_{2}\right),$ (6)
 where $W_{1}\in\mathbb{R}^{|\bm{\lambda}|\times d},\ W_{2}\in\mathbb{R}^{d\times 1},\ \mathbf{b}_{1}\in\mathbb{R}^{d},\ \mathbf{b}_{2}\in\mathbb{R}^{1}$ are learnable parameters.
-We set the minimum value of the generated bias to $1$, resulting in faster model convergence.
-Due to the relation $r_{ij}$ is scale-independent, we remove scale $N$ in adaptation function for $\bm{R}$, and if $\bm{R}$ is absent, we substitute a zero vector for $\bar{\mathbf{h}}_{i}^{(2)}$.
-As shown in Figure 1(b), unlike MatNet (Kwon et al., 2021), which processes $\bm{D}$ and $\bm{D^{\mathrm{T}}}$ using two parallel attention layers, we mix them into one shared attention layer and let separate heads address different priors.
-This head‑level incorporation not only cuts architectural redundancy but also enhances model flexibility (e.g., integrating relation matrix), enabling URS can achieve impressive performance for various VRP variants.
-Decoder
-Figure 2: Comparison between our URS and existing adapter-based fine-tuning frameworks.
-In URS, all parameters of $\bm{\theta}_{dec}(\bm{\lambda})$ are generated based on the multi-hot problem representation $\bm{\lambda}$.
+We set the minimum bias to $1$, resulting in faster model convergence.
+For decoder $\bm{\theta}_{dec}(\bm{\lambda})$, all parameters are generated based on $\bm{\lambda}$.
 Following Kwon et al.
@@
 where $\cdot$$\cdot$] is the horizontal concatenation operator, $W_{Q}(\bm{\lambda})\in\mathbb{R}^{(2d+1)\times d}$ is a linear projection matrix conditioned on $\bm{\lambda}$.
-Missing $C_{t}$ is zero-padded (see Appendix E.3 for detailed settings).
-The new context embedding $\hat{\mathbf{h}}_{(C)}^{t}$ is obtained via $\mathrm{Attention}(\cdot)$ (Zhou et al., 2024a) on $\mathbf{h}_{(C)}^{t}$ and $H^{(L)}$:
-$\hat{\mathbf{h}}_{(C)}^{t}=\mathrm{Attention}(\mathbf{h}_{(C)}^{t},H^{(L)}W_{K}(\bm{\lambda}),H^{(L)}W_{V}(\bm{\lambda}),\mathcal{M}_{t},f(\alpha,N,d_{t,i})),$ (8)
-where $W_{K}(\bm{\lambda})\in\mathbb{R}^{d\times d}$, $W_{V}(\bm{\lambda})\in\mathbb{R}^{d\times d}$ are linear projection matrices conditioned on $\bm{\lambda}$.
-Finally, we compute the selection probability for each feasible node $p_{\bm{\theta}}(\pi_{t}=i\mid\mathcal{G},\pi_{1:t-1})$ based on the improved compatibility with adaptation bias (Zhou et al., 2024a):
-$p_{\bm{\theta}}(\pi_{t}=i\mid\mathcal{G},\pi_{1:t-1})=\mathrm{softmax}(\zeta\cdot\text{tanh}(\frac{\hat{\mathbf{h}}_{(C)}^{t}(\mathbf{h}_{i}^{(L)})^{\mathrm{T}}}{\sqrt{d}}+f(\alpha,N,d_{t,i}))+\mathcal{M}_{t})$ (9)
+Missing $C_{t}$ is zero-padded (see Appendix E.3 for details).
+The new context embedding $\hat{\mathbf{h}}_{(C)}^{t}$ is obtained via $\mathrm{Attention}(\cdot)$ (Zhou et al., 2026) on $\mathbf{h}_{(C)}^{t}$ and $H^{(L)}$:
+$K=H^{(L)}W_{K}(\bm{\lambda}),\qquad V=H^{(L)}W_{V}(\bm{\lambda}),$ (8)
+$\hat{\mathbf{h}}_{(C)}^{t}=\mathrm{Attention}(\mathbf{h}_{(C)}^{t},K,V,\mathcal{M}_{t},f(\alpha,N,d_{t,i})),$ (9)
+where $W_{K}(\bm{\lambda}),W_{V}(\bm{\lambda})\in\mathbb{R}^{d\times d}$ are linear projection matrices conditioned on $\bm{\lambda}$.
+Finally, we compute probabilities for feasible nodes $p_{\bm{\theta}}(\pi_{t}=i\mid\mathcal{G},\pi_{1:t-1})=\mathrm{softmax}(\mathbf{u})$ based on the improved compatibility (Zhou et al., 2026):
+$u^{t}_{i}=\begin{cases}\zeta\cdot\text{tanh}(\frac{\hat{\mathbf{h}}_{(C)}^{t}(\mathbf{h}_{i}^{(L)})^{\mathrm{T}}}{\sqrt{d}}+f(\alpha,N,d_{t,i}))&\text{if}\ i\not\in\{\pi_{1:t-1}\}\\ -\infty&\text{otherwise,}\end{cases}$ (10)
+Figure 3: Comparison between our URS and existing adapter-based fine-tuning frameworks.
 where $\zeta$ is the clipping parameter.
@@
 For $\alpha$ in Equation 9 and Equation 10, we generate them via $\mathrm{BIAS}(\bm{\lambda})$ in Equation 6.
-As shown in Figure 2, compared with adapter-based frameworks (Lin et al., 2024; Drakulic et al., 2025; Wang et al., 2025a), we provide a unified decoding entrance, which $\mathrm{WEIGHT}(\bm{\lambda})$ and $\mathrm{BIAS}(\bm{\lambda})$ maps a problem representation vector $\bm{\lambda}$ directly to all parameters of the decoder.
-This conditional weight generation removes problem‑specific adapters, enabling zero-shot generalization to unseen problems while maintaining solution quality.
-4 LLM-driven Constraint Satisfaction Mechanism
-Figure 3: The overview of LLM-driven Constraint Satisfaction Mechanism.
-(a) The generating pipeline of LLM-driven Constraint Satisfaction Mechanism for a general routing problem; (b) The illustration of mask generator utilization, handling the complex constraint.
-While the proposed UDR enables URS to handle a wide range of routing problems, ensuring solution feasibility across diverse and potentially unseen constraints remains a major challenge.
-Traditional NCO approaches rely on manually designed masks $\mathcal{M}_{t}$ for each specific VRP variant Kool et al.
-(2019); Kwon et al.
-(2020); Luo et al.
-(2023), which becomes impractical when solving over 100 distinct variants as targeted by URS.
-To address this bottleneck, we propose an LLM-driven constraint satisfaction mechanism that automatically produces problem-specific mask generators.
-This mechanism transforms constraint handling from a manual engineering task into an automated program design task, thereby effectively supporting URS’s excellent zero-shot generalization capabilities.
-As illustrated in Figure 3 (a), our constraint satisfaction mechanism operates through a systematic generate-and-validate pipeline.
-The process begins by prompting an LLM with both a problem-specific natural language constraint description and a standardized Python code template.
-The LLM then designs a set of candidate mask generators that implement the constraint logic.
-Each candidate is validated on a set of problem instances using a problem-specific checker to verify its reliability in producing feasible solutions.
-Invalid candidates are discarded, and the generation process iterates until a reliable mask generator is obtained.
-Specifically, during validation, the candidate mask generator guides the stepwise solution construction.
-As illustrated in Figure 3 (b), at each step $t$, the mask generator takes the current state of the partial solution $\pi_{1:t-1}$ as input, which contains various information such as visited nodes, remaining load, etc.
-It then produces a binary mask $\mathcal{M}_{t}$ that identifies all permissible nodes.
-This mask is directly integrated into the URS’s selection probability computation in Equation 9 to ensure feasibility.
-To efficiently steer this generate-and-test process, we formulate it as an optimization problem that maximizes the solution validity rate.
-The well-developed LLM-based automated algorithm design frameworks, such as EoH (Liu et al., 2024b), can be effectively applied to tackle this optimization problem.
-The optimization process terminates immediately once a reliable mask generator is found, which significantly reduces computational overhead compared to exhaustive search approaches.
-Furthermore, generated mask generators can be cached and reused for different instances sharing the same problem variant, thereby amortizing the generation cost.
-For more details about our LLM-driven constraint satisfaction mechanism, please see Appendix G.
+Table 2: Performance comparison on $8$ seen less-explored variants.
+Method OP100 PCTSP100 PDTSP100 ACVRP100 CVRPTW100 CVRPB100 OCVRP100 OCVRPTW100
+Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time
+Oracle 0.00% 1.5m 0.00% 1.2h 0.00% 9.8m 0.00% 6.6m 0.00% 19.6m 0.00% 20.8m 0.00% 5.3m 0.00% 20.8m
+Sym-NCO 0.68% 15s 0.14% 16s - - - - - -
+BQ 0.63% 7s - - - - - - -
+Heter-AM - - 6.61% 5m - - - - -
+MVMoE - - - - 4.90% 15s 1.28% 11s 3.14% 13s 3.85% 15s
+MTPOMO - - - - 5.31% 8s 1.67% 5s 3.46% 6s 4.41% 7s
+ReLD-MTL - - - - 4.56% 10s 0.90% 6s 2.32% 7s 3.10% 9s
+GOAL-MTL 1.20% 38s - - - 4.66% 42s - - -
+URS-STL 0.11% 4s -0.09% 5s 3.22% 4s 2.01% 1.4m 4.50% 8s 1.59% 6s 2.25% 6s 3.22% 8s
+URS 0.45% 4s 1.06% 5s 4.98% 4s 3.06% 1.4m 6.13% 8s 1.46% 6s 3.24% 6s 5.07% 8s
+Table 3: Experimental results on seen ATSP, TSP, and CVRP.
+Method ATSP100 TSP100 CVRP100
+Gap Time Gap Time Gap Time
+Oracle 0.00% 2m 0.00% 6m 0.00% 9.1m
+Sym-NCO - 0.14% 5s 1.46% 6s
+BQ 1.27% 1m 0.35% 7s 3.24% 8s
+LEHD - 0.59% 2s 3.68% 2s
+POMO - 0.13% 5s 1.25% 6s
+ICAM 3.05% 1m 0.15% 4s 2.04% 4s
+MatNet 0.95% 6.5m - -
+MVMoE - - 1.65% 12s
+MTPOMO - - 1.85% 6s
+ReLD-MTL - - 1.42% 8s
+GOAL-MTL 1.77% 1m - 4.22% 48s
+URS-STL 0.62% 1.1m 0.08% 6s 1.63% 6s
+URS 2.26% 1.1m 0.57% 6s 1.81% 6s
 4 Experiments
-In this section, we comprehensively evaluate URS against classical and learning-based solvers across more than one hundred distinct VRP variants.
-For URS, we focus on: (1) performance on seen problems and (2) zero-shot generalization under unseen ones.
-To ensure comparability with prior neural routing studies, we adopt the 100-node setting as our primary evaluation scenario for URS.
-All experiments are conducted on a single NVIDIA GeForce RTX 4090 GPU (24GB memory).
+In this section, we present a comprehensive evaluation of URS against both classical and neural solvers.
+For URS, we focus on four key aspects: (1) performance on 11 seen problems; (2) zero-shot generalization on 99 unseen problems; (3) scalability to large-scale VRP variants; and (4) results on benchmark datasets.
+All experiments are conducted using a single NVIDIA GeForce RTX 4090 GPU (24GB of memory) for neural solvers and an Intel Xeon Gold 6240R CPU (2.40 GHz) for classical solvers.
+Table 4: Zero-shot generalization on 1K test instances of $32$ unseen VRPs ($N=100$). Due to page limit, complete experimental results (including all 99 unseen variants) are deferred to Appendix I.
+Method CVRPBP CVRPBPL MDOCVRPBP MDOCVRPBPTW MDOCVRPBPL MDOCVRPBPLTW MDCVRP MDOCVRP
+Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time
+Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
+MVMoE 13.95% 14s 13.38% 14s 56.60% 33s 63.77% 38s 56.26% 34s 62.47% 40s 33.06% 31s 29.08% 31s
+MTPOMO 13.71% 7s 13.44% 8s 42.94% 24s 47.09% 29s 41.98% 27s 46.26% 32s 23.38% 23s 27.46% 23s
+ReLD-MTL 13.57% 9s 12.96% 10s 37.74% 27s 45.16% 34s 37.45% 29s 44.74% 36s 18.39% 25s 22.32% 25s
+URS 12.95% 7s 12.65% 8s 24.44% 22s 26.31% 26s 24.22% 24s 26.20% 28s 15.05% 23s 22.01% 23s
+Method MDCVRPL MDOCVRPTW MDOCVRPB MDOCVRPBL MDOCVRPLTW SPCTSP PDCVRP OPDCVRP
+Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time
+Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 1.6h 0.00% 1.6h 0.00% 1.6h
+MVMoE 33.42% 34s 51.57% 40s 28.21% 26s 28.70% 28s 51.09% 43s - - -
+MTPOMO 24.32% 25s 35.71% 29s 25.06% 20s 25.01% 21s 35.39% 31s - - -
+ReLD-MTL 18.48% 28s 33.55% 33s 18.50% 21s 18.50% 22s 33.26% 35s - - -
+URS 15.32% 26s 26.05% 25s 15.17% 18s 14.90% 20s 25.59% 27s -2.37% 5s -1.47% 5s 4.93% 5s
+Method ACVRPL ACVRPBL ACVRPB ACVRPBTW APDTSP AMDCVRPL AMDCVRP APDCVRP
+Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time
+Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
+URS -2.24% 1.9m 6.18% 2.4m 7.10% 2.1m 9.95% 2.2m 6.21% 1m 7.15% 6.1m 8.11% 5.3m 7.03% 1.2m
+Method ACVRPLTW ACVRPBLTW ACVRPBPLTW AOCVRPBLTW ACVRPTW ACVRPBPTW AOPDCVRP AOP
+Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time
+Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
+URS 12.22% 2.7m 10.18% 2.4m 13.82% 2.8m 14.32% 2.5m 11.28% 2.5m 14.52% 2.6m 11.13% 1.2m -5.47% 1.2m
+Table 5: Generalization on different large-scale VRP variants.
+All models are trained on instances of size $N=100$ and employ instance augmentation $\times$ 8.
+PyVRP and OR-Tools can process all instances of a given size concurrently to ensure a fair comparison.
+Problem Method N = 1000 N = 2000 N = 3000 N = 4000 N = 5000
+Obj.(Gap) Time Obj.(Gap) Time Obj.(Gap) Time Obj.(Gap) Time Obj.(Gap) Time
+CVRPTW PyVRP 159.35(0.00%) 20m 337.15(0.00%) 40m 516.97(0.00%) 1h 618.58(0.00%) 2.7h 787.25(0.00%) 3.3h
+OR-Tools 178.13(11.79%) 20m 357.18(5.94%) 40m 533.97(3.29%) 1h 647.92(4.74%) 2.7h 807.128(2.52%) 3.3h
+MTPOMO 219.22(37.57%) 1.7m 488.40(44.86%) 14.6m 773.00(49.53%) 49.5m 975.61(57.72%) 1.9h OOM
+MVMoE 233.53(46.55%) 1.7m 624.73(85.30%) 14.6m 1094.25(111.67%) 51.9m 1486.34(140.28%) 2h OOM
+ReLD-MTL 183.31(15.04%) 1.6m 396.28(17.54%) 13.4m 613.58(18.69%) 45.7m 748.65(21.03%) 1.8h OOM
+URS 172.58(8.30%) 1.3m 365.27(8.34%) 11.1m 559.03(8.14%) 30.1m 673.63(8.90%) 1.2h 853.79(8.45%) 2.3h
+CVRPB PyVRP 36.52(0.00%) 20m 51.80(0.00%) 40m 75.12(0.00%) 1h 93.52(0.00%) 2.7h 112.21(0.00%) 3.3h
+OR-Tools 44.23(21.11%) 20m 63.33(22.26%) 40m 88.6(17.94%) 1h 108.82(16.36%) 2.7h 128.11(14.17%) 3.3h
+MTPOMO 45.92(25.77%) 1.1m 83.50(61.20%) 8.8m 136.52(81.73%) 31.4m 197.99(111.70%) 1.2h OOM
+MVMoE 98.60(170.03%) 1m 250.47(383.57%) 8.1m 332.93(343.19%) 28.9m 423.19(352.49%) 1.1h OOM
+ReLD-MTL 38.52(5.48%) 1m 59.21(14.30%) 8m 88.36(17.62%) 28.3m 114.88(22.83%) 1.1h OOM
+URS 34.86(-4.54%) 40s 50.39(-2.72%) 5.1m 72.01(-4.15%) 17.3m 91.74(-1.91%) 41.2m 111.93(-0.24%) 1.3h
+OCVRPTW PyVRP 90.91(0.00%) 20m 164.00(0.00%) 40m 224.16(0.00%) 1h 299.45(0.00%) 2.7h 367.86(0.00%) 3.3h
+OR-Tools 105.98(16.58%) 20m 207.041(26.24%) 40m 280.88(25.30%) 1h 370.85(23.84%) 2.7h 447.02(21.52%) 3.3h
+MTPOMO 147.14(61.85%) 1.6m 297.53(81.43%) 12.8m 437.73(95.28%) 44.3m 601.63(100.91%) 1.7h OOM
+MVMoE 147.35(62.09%) 1.6m 389.60(137.57%) 12.9m 634.11(182.88%) 44.7m 899.18(200.28%) 1.7h OOM
+ReLD-MTL 110.00(21.01%) 1.5m 208.60(27.20%) 12.3m 292.70(30.57%) 42m 394.42(31.72%) 1.6h OOM
+URS 98.76(8.63%) 1m 178.07(8.58%) 7.6m 243.88(8.80%) 25.4m 325.64(8.74%) 1h 398.72(8.39%) 1.9h
 4.1 Experimental Setup
 Problem Setting
-We train URS on $11$ mixed VRP variants to ensure that each feature in UDR has been seen at least once, i.e., $\mathcal{P}=\{P_{i}\}_{i=1}^{11}$.
-Specifically, training problems and related data generations are as follows: (1) ATSP (Kwon et al., 2021);(2) TSP (Kool et al., 2019); (3) CVRP (Kool et al., 2019); (4) ACVRP (Kwon et al., 2021; Kool et al., 2019); (5) Orienteering Problem (OP) (Kool et al., 2019); (6) PCTSP (Kool et al., 2019); (7) PDTSP (Li et al., 2021); (8) CVRPTW (Zhou et al., 2024b); (9) OCVRP (Zhou et al., 2024b); (10) CVRPB (Zhou et al., 2024b), and (11) OCVRPTW (Zhou et al., 2024b).
-To evaluate the zero-shot generalization of URS, we construct 96 unseen variants to further highlight the broad applicability of URS.
-And each test dataset consists of 1000 instances.
-The oracle solver for each problem can be found in Appendix D.
+We train URS on $11$ mixed variants to ensure that most features in UDR have been seen at least twice, which prevents URS from overfitting a specific feature to a single problem.
+Training problems and related data generations are as follows: (1) ATSP (Kwon et al., 2021);(2) TSP (Kool et al., 2019); (3) CVRP (Kool et al., 2019); (4) ACVRP (Kwon et al., 2021; Kool et al., 2019); (5) Orienteering Problem (OP) (Kool et al., 2019); (6) PCTSP (Kool et al., 2019); (7) PDTSP (Li et al., 2021); (8) CVRPTW (Zhou et al., 2024); (9) OCVRP (Zhou et al., 2024); (10) CVRPB (Zhou et al., 2024), and (11) OCVRPTW (Zhou et al., 2024).
+We provide further discussion of problem selection in Appendix J.1.
+Each of 110 variants comprises 1,000 instances.
+The related oracle solvers and problem representations are detailed in Appendix D.
 Model & Training Setting
@@
 We set the number of attention layers in the encoder to $12$.
-The clipping parameter $\zeta=50$ in Equation 9 for better training convergence (Jin et al., 2023).
-For training, URS is trained by the REINFORCE (Williams, 1992) algorithm, and the data is generated on the fly (see Appendix E.4 for more details).
-We train URS for $500$ epochs with $2\text{,}000$ batches per epoch, and the batch size is $128$.
-At each training step, we randomly select one problem $P_{i}$ from the problem set $\mathcal{P}$.
-We employ the AdamW optimizer (Loshchilov & Hutter, 2017) with an initial learning rate of $10^{-4}$, which is decayed by a factor of 0.1 at the 451451st epoch.
+The clipping parameter $\zeta$ is set to $50$ in Equation 10.
+URS is trained by the REINFORCE (Williams, 1992) algorithm, and the data is generated on the fly (see Appendix E.4 for more details).
+We train URS for $500$ epochs with $2\text{,}000$ batches per epoch, and each batch has a size of $128$.
+At each training step, we randomly select one problem from $11$ problems.
+We employ the AdamW optimizer (Loshchilov and Hutter, 2017) with an initial learning rate of $10^{-4}$, which is decayed by a factor of 0.1 at the 451451st epoch.
 Weight decay is set to $10^{-6}$.
-More details about the model training are provided in Appendix H.
+More details are provided in Appendix H.
 Baseline
-We compare URS with the following methods: (1)Classical Solvers: [^1] We run LKH3 (Helsgaun, 2017) with $10000$ trials and 1 run (Kool et al., 2019).
-For PyVRP and OR-Tools, we run them on a single CPU core with time limits of 20s (Li et al., 2025a).
-PyVRP (Wouda et al., 2024), LKH3 (Helsgaun, 2017), and OR-Tools (Perron & Furnon, 2023); (2)Neural Solvers: (i) specialized solvers that support multiple VRP variants: Sym-NCO (Kim et al., 2022), BQ (Drakulic et al., 2023), LEHD (Luo et al., 2023), ICAM (Zhou et al., 2024a) and POMO (Kwon et al., 2020); (ii) general neural solvers capable of handling multiple variants [^2] Given the differences in problem variants and visible task training among the baseline models, we only compare methods with the same training setup.
-Thus, RF Berto et al.
-(2024) and CaDA (Li et al., 2025a) are not included in our experiments to ensure a fair comparison.: MTPOMO (Liu et al., 2024a), MVMoE (Zhou et al., 2024b), ReLD-MTL (Huang et al., 2025), GOAL-MTL (Drakulic et al., 2025); (iii) Solvers for less-explored variants: MatNet (Kwon et al., 2021), Heter-AM (Li et al., 2021).
+We compare URS with the following methods: (1)Classical Solvers: PyVRP (Wouda et al., 2024), LKH3 (Helsgaun, 2017), and OR-Tools (Perron and Furnon, 2023).
+We run LKH3 (Helsgaun, 2017) with $10000$ trials and 1 run (Kool et al., 2019).
+For PyVRP and OR-Tools, we run them on a single CPU core with time limits of 20s (Li et al., 2025a); (2)Neural Solvers: (i) specialist solvers: Sym-NCO (Kim et al., 2022), BQ (Drakulic et al., 2023), LEHD (Luo et al., 2023), ICAM (Zhou et al., 2026), POMO (Kwon et al., 2020), ELG (Gao et al., 2024), MatNet (Kwon et al., 2021), and Heter-AM (Li et al., 2021); (ii) generalist neural solvers: MTPOMO (Liu et al., 2024), MVMoE (Zhou et al., 2024), ReLD-MTL (Huang et al., 2025), and GOAL-MTL (Drakulic et al., 2025).
+Given the disparities in problem formulations and training settings, detailed comparisons against RouteFinder (Berto et al., 2025) and CaDA (Li et al., 2025a) are deferred to Appendix J.2.
 Metrics and Inference
-We report the average objective value (Obj.), optimality gap (Gap), and total inference time (Time) for each method.
-The optimality gap quantifies the discrepancy between the solutions generated by the corresponding methods and the near-optimal solutions, which are obtained using classical solvers.
-For most NCO baseline methods, we execute the source code provided by the authors using default settings.
-We compare the performance of URS to the relevant baselines for each problem as well as its single-task version (i.e., URS-STL).
-For symmetric instances, we report the best result with $\times 8$ instance augmentation, following Kwon et al.
-(2020).
-For asymmetric instances, we report the best result with $\times 128$ instance augmentation, following Kwon et al.
-(2021).
-Table 2: Performance on 1K test instances of $11$ seen VRPs.
-Method ATSP TSP OP PCTSP PDTSP -
-Gap Time Gap Time Gap Time Gap Time Gap Time
-Oracle Solver 0.00% 2m 0.00% 6m 0.00% 1.5m 0.00% 1.2h 0.00% 9.8m -
-Sym-NCO - 0.14% 5s 0.68% 15s 0.14% 16s - -
-BQ 1.27% 1m 0.35% 7s 0.63% 7s - - -
-LEHD - 0.59% 2s - - - -
-POMO - 0.13% 5s - - - -
-ICAM 3.05% 1m 0.15% 4s - - - -
-MatNet 0.95% 6.5m - - - - -
-Heter-AM - - - - 6.61% 5m -
-GOAL-MTL 1.77% 1m - 1.20% 38s - - -
-[…]
-URS 2.26% 1.1m 0.57% 6s 0.45% 4s 1.06% 5s 4.98% 4s -
-Method ACVRP CVRP CVRPTW CVRPB OCVRP OCVRPTW
-Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time
-Oracle Solver 0.00% 6.6m 0.00% 9.1m 0.00% 19.6m 0.00% 20.8m 0.00% 5.3m 0.00% 20.8m
-Sym-NCO - 1.46% 6s - - - -
-BQ - 3.24% 8s - - - -
-LEHD - 3.68% 2s - - - -
-[…]
-ICAM - 2.04% 4s - - - -
-MVMoE - 1.65% 12s 4.90% 15s 1.28% 11s 3.14% 13s 3.85% 15s
-MTPOMO - 1.85% 6s 5.31% 8s 1.67% 5s 3.46% 6s 4.41% 7s
-[…]
-GOAL-MTL - 4.22% 48s 4.66% 42s - - -
-[…]
-URS 3.06% 1.4m 1.81% 6s 6.13% 8s 1.46% 6s 3.24% 6s 5.07% 8s
-Table 3: Zero-shot generalization on 1K test instances of $30$ unseen VRPs. Due to page limit, complete experimental results (including all 96 unseen variants) are deferred to Appendix I.
-Method CVRPBP OCVRPBP CVRPBPL OCVRPBPTW CVRPBPLTW
-Gap Time Gap Time Gap Time Gap Time Gap Time
-Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
-MVMoE 13.95% 14s 17.36% 14s 13.38% 14s 2.29% 15s 10.46% 16s
-MTPOMO 13.71% 7s 16.94% 8s 13.44% 8s 3.26% 9s 11.35% 10s
-ReLD-MTL 13.57% 9s 14.83% 9s 12.96% 10s 2.27% 11s 10.80% 12s
-[…]
-Method MDCVRP MDCVRPL MDOCVRPB MDOCVRPBPTW MDOCVRPBPLTW
-Gap Time Gap Time Gap Time Gap Time Gap Time
-Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
-MVMoE 31.90% 31s 33.42% 44s 28.21% 26s 46.18% 40s 45.23% 40s
-MTPOMO 23.75% 27s 24.32% 25s 25.06% 20s 34.19% 29s 33.37% 31s
-ReLD-MTL 17.37% 36s 18.48% 28s 18.50% 21s 34.08% 33s 33.83% 36s
-[…]
-Method ACVRPB ACVRPL ACVRPTW AOPDCVRP ACVRPBL
-Gap Time Gap Time Gap Time Gap Time Gap Time
-Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
-[…]
-Method AOCVRPBPTW ACVRPBPLTW ACVRPBPTW AOCVRPBPLTW ACVRPBTW
-Gap Time Gap Time Gap Time Gap Time Gap Time
-Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
-[…]
-Method AMDCVRPL AMDCVRPBPLTW AMDOCVRPBPLTW PDCVRP OPDCVRP
-Gap Time Gap Time Gap Time Gap Time Gap Time
-Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
-[…]
-Method AMDOCVRPBPTW AMDCVRPBPTW AMDCVRP APDTSP APDCVRP
-Gap Time Gap Time Gap Time Gap Time Gap Time
-Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
-[…]
+We report the optimality gap (Gap) and total inference time (Time) for each method.
+The optimality gap quantifies the discrepancy between the corresponding solvers and oracle solvers.
+For neural solvers, we execute the source code provided by the authors using default settings.
+We compare the performance of URS with relevant baselines for each problem, as well as its single-task version (i.e., URS-STL).
+The URS-STL variant uses exactly the same architecture and hyperparameters (see Appendix H.2 for more details).
+For URS, we report the best result with $\times 8$ (Kwon et al., 2020) and $\times 128$ (Kwon et al., 2021) instance augmentation for symmetric and asymmetric instances, respectively.
+The mark (-) indicates that the method does not support the corresponding setting.
 4.2 Performance Evaluation
 Performance on Seen VRPs
-The experimental results on trained VRPs with scale $100$ are reported in Table 2.
-We observe that among all comparable neural solvers, URS-STL achieves the best performance on 8 out of the 11 routing problems, complemented by remarkably fast inference, highlighting the strength of the URS architecture when fully specialized.
-Meanwhile, mixing all problems into the training of URS incurs only a minor performance drop, yet still delivers competitive performance across all problems, demonstrating that potential connections among different problems are effectively captured by URS.
+As presented in Table 2 and Table 3, among comparable neural solvers, URS-STL achieves the lowest gap on 8 of the 11 VRPs while offering fast inference, highlighting the strength of the URS architecture when fully specialized.
+Although URS-MTL exhibits a performance degradation relative to STL variants, it remains competitive across all 11 problems.
+Using ATSP as an example, we attribute this drop to geometric structural inequality, which likely biases the model’s learning capacity toward symmetric VRPs.
+While the lack of explicit constraint states yields weaker performance compared to ReLD-MTL and MVMoE, URS shows clear advantages when generalizing to unseen VRP variants (Table 4) and scaling to large-size instances (Table 5).
 Generalization to Unseen VRPs
-We evaluate our URS model on zero-shot generalization across 96 unseen VRP variants, including 44 well-studied VRP variants from the NCO community and 52 newly introduced variants that have not been explored before.
-The detailed results are presented in Table 3 and Appendix I.
-On widely-studied constraints (i.e., C, O, L, TW, MD, B, BP), URS continues to deliver high-quality solutions.
-Notably, URS merely selects the next node from a given feasible candidates without any additional domain knowledge, making its performance especially meaningful.
-Benefiting from our UDR, URS can directly address previously unexplored variants while maintaining solution quality, which compounds asymmetry and PD relations.
-Without any fine-tuning, URS still constructs high-quality solutions for more than 90 VRP variants that are not seen in training, highlighting its strong zero-shot generalization ability.
+We evaluate our URS on zero-shot generalization across 99 unseen VRP variants.
+As presented in Table 4 and Appendix I, on 44 widely-studied VRPs with diverse constraint combinations (i.e., C, O, L, TW, MD, B, BP), URS continues to deliver high-quality solutions.
+Notably, URS merely selects the next node from a given set of feasible candidates without any additional domain knowledge, making its performance especially meaningful.
+Benefiting from our UDR, URS can directly address 55 less-explored VRP variants that compound asymmetry and PD relations.
+These results highlight URS’s strong capacity for cross-problem zero-shot generalization without any fine-tuning.
+Scalability to Larger-scale VRP Variants
+[…]
 Results on Benchmark Dataset
-We further evaluate the generalization ability of URS on benchmark instances from CVRPLIB Set-X (Uchoa et al., 2017) with scale $\in[500,1000]$, and we compare URS against several representative single-task and multi-task models.
-As demonstrated in Table 12 (Appendix J), URS maintains its position as the best overall performance across instances of varying scales.
-Notably, RL-based URS surpasses the SL-based specialized model LEHD (8.678% vs.
-12.836%), underscoring its practical applicability in real-world scenarios.
+We further evaluate URS on benchmark instances from CVRPLib Set-X (Uchoa et al., 2017) and CVRPLib Set-AGS (also known as Set-XXL) (Arnold et al., 2019).
+We compare URS against several representative specialist and generalist solvers.
+As presented in Table 6, URS can significantly outperform existing generalist neural routing solvers in large-scale benchmark instances.
+Notably, as a generalist solver, URS also greatly surpasses many representative specialist models (e.g., LEHD, BQ, and ELG).
+In addition, the detailed results (Appendix J.3) on the Set-X dataset show that URS still maintains its position as the best overall performance.
+These results further underscore the practical applicability of URS in real-world large-scale scenarios.
+Table 6: Comparison on Set-AGS with $N\in[3000,7000]$. All models are trained on instances of size $N=100$.
+Type Method Leuven1 Leuven2 Antwerp1 Antwerp2 Avg.gap
+(N=3000) (N=4000) (N=6000) (N=7000)
+Specialist LEHD greedy 16.60% 34.86% 14.66% 22.77% 22.22%
+BQ greedy 18.53% 30.70% 16.48% 27.67% 23.34%
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+GOAL greedy OOM OOM OOM OOM -
+[…]
 5 Ablation Study
-Table 4: Effects of different components of URS.
-Method Avg.gap
-w/o $\bm{\xi}$ 3.78%
-w/o $C_{t}$ 2.97%
-$\{\bm{D}\}$ 6.33%
-$\{\bm{D},\bm{D^{\mathrm{T}}}\}$ 4.07%
-$\{\bm{D},\bm{R}\}$ 7.34%
-POMO-Dec. 3.39%
-ReLD-Dec. 3.87%
-[…]
-In this section, we conduct an ablation study across 11 seen VRP variants to validate the effect of key components of URS, mainly including: (1) without node-type indicator $\bm{\xi}$ of UDR; (2) without problem state $C_{t}$; (3) effects of different priors in MBM ; (4) effects of alternative decoder architectures.
-As detailed in Table 4.
-URS maintains its competitive performance without $C_{t}$.
-And its effectiveness is further improved by the inclusion of a node-type indicator $\bm{\xi}$.
+In this section, we conduct a comprehensive ablation study and analysis to demonstrate the effectiveness and robustness of URS.
+The detailed results and discussion are presented below.
+Note that, unless otherwise stated, the instance augmentation technique is used for the ablation studies.
+Effects of Key Components
+We conduct an ablation study across 11 seen VRP variants to validate the effect of key components of URS, mainly including: (1) without node-type indicator $\bm{\xi}$ of UDR; (2) without problem state $C_{t}$; (3) effects of different priors in MBM ; (4) effects of alternative decoder architectures.
+As detailed in Figure 4, URS maintains its competitive performance without $C_{t}$; its effectiveness is further improved by the inclusion of a node-type indicator $\bm{\xi}$.
 For MBM, the results reveal that three priors $\{\bm{D},\bm{D}^{\mathrm{T}},\bm{R}\}$ are complementary, and their combined use allows MBM to effectively learn the geometric and relational biases inherent in various problems, resulting in superior overall performance.
-Furthermore, replacing the decoder with POMO or ReLD confirms the strong cross-problem robustness of URS, while the integration of the decoder in ICAM (Zhou et al., 2024a) leads to a further improvement.
+Furthermore, replacing the decoder with POMO or ReLD confirms the strong cross-problem robustness of URS, while integrating the decoder introduced in Zhou et al.
+(2026) further improves cross-problem performance.
 More detailed results and analyses can be found in Appendix K.
+Figure 4: Effects of different components of URS.
+Table 7: Results comparing performance with and without MBM, $\mathrm{WEIGHT}(\bm{\lambda})$, and $\mathrm{BIAS}(\bm{\lambda})$ on seen and unseen variants.
+MBM $\mathrm{WEIGHT}(\bm{\lambda})$ $\mathrm{BIAS}(\bm{\lambda})$ Avg.gap Best
+Seen (11) Unseen (11) Solution
+$\times$ $\checkmark$ $\checkmark$ 6.33% 10.41% 3/22
+$\checkmark$ $\times$ $\times$ 4.20% 8.12% 0/22
+$\checkmark$ $\times$ $\checkmark$ 3.84% 8.02% 1/22
+$\checkmark$ $\checkmark$ $\times$ 3.63% 7.81% 1/22
+$\checkmark$ $\checkmark$ $\checkmark$ 2.74% 7.05% 17/22
+MBM & Parameter Generator
+To further evaluate the effects of MBM, $\mathrm{WEIGHT}(\bm{\lambda})$, and $\mathrm{BIAS}(\bm{\lambda})$ for cross-problem zero-shot generalization, we conduct tests on 11 seen problems and 11 unseen CVRP variants widely studied in (Liu et al., 2024).
+As shown in Table 7, the full URS model achieves the lowest average gap on both seen and unseen problems, and provides the best solution on the majority of tasks (17/22).
+This shows that their synergy is essential for robust cross-problem zero-shot generalization.
+We provide detailed results in Appendix K.5.
+Table 8: Performance comparison of ablating each UDR component (i.e., $\bm{\rho}$, $\bm{\omega}$, and $\bm{\xi}$) on seen and unseen variants.
+Since removing positional identifiers $\bm{\rho}$ prevents the use of instance augmentation, this augmentation technique is disabled across all results to maintain a consistent inference strategy.
+$\bm{\rho}$ $\bm{\omega}$ $\bm{\xi}$ Avg.gap Best
+Seen (11) Unseen (11) Solution
+$\times$ $\checkmark$ $\checkmark$ 4.65% 18.12% 3/22
+$\checkmark$ $\times$ $\checkmark$ 9.81% 18.60% 0/22
+$\checkmark$ $\checkmark$ $\times$ 5.43% 18.60% 4/22
+$\checkmark$ $\checkmark$ $\checkmark$ 4.17% 8.51% 15/22
+Effects of Each UDR Component
+To clarify the independent contribution of each UDR component, we conduct ablations by removing each component in turn.
+As shown in Table 8, the absence of any single component leads to notable performance degradation.
+A complete UDR achieves the lowest average gap, affirming the necessity of our UDR design for robust zero-shot generalization.
+The detailed results for each VRP variant can be found in Appendix K.6.
+Hypernetwork vs. Adapter
+We train URS-Adapter, a URS variant that adopts a distinct decoder for each of the 11 training problems, and compare it with the original URS with a hypernetwork on the same 11 seen problems and large-scale CVRP variants (see Appendix K.7 for detailed results).
+URS consistently outperforms URS-Adapter across the 11 problems seen, achieving higher parameter efficiency.
+This performance advantage holds consistently across large-scale CVRP variants, further solidifying our conclusion.
 6 Conclusion, Limitation, and Future Work
-In this work, we propose a novel RL-based Unified Routing Solver (URS) for cross-problem zero-shot generalization.
-The core of URS is a Unified Data Representation (UDR), which broadens problem coverage by substituting problem enumeration with data unification.
-Its generalization capability is further enhanced by a Mixed Bias Module (MBM) and an adaptive parameter generator.
-Moreover, an LLM-based constraint satisfaction mechanism guarantees solution feasibility and reduces dependency on domain expertise.
-Extensive experiments demonstrate that URS can produce high-quality solutions for more than 100 distinct VRP variants without any fine-tuning, more than 90 of which are unseen during training.
+In this work, we propose a novel RL-based URS for cross-problem zero-shot generalization.
+The core of URS is the proposed UDR, which broadens problem coverage by substituting problem enumeration with data unification.
+Its generalization capability is further enhanced by an MBM and a problem-conditioned parameter generator.
+Extensive experiments show that URS can produce high-quality solutions for 110 VRP variants without any fine-tuning, including 99 unseen variants, significantly broadening the problem coverage and reducing reliance on domain expertise.
 Limitation and Future Work
-While URS achieves impressive results on a wide range of VRP variants, its current training scheme does not accommodate their inherent differences in complexity.
-In the future, we aim to explore more effective training schemes to further improve training efficiency.
+While URS achieves impressive results across a wide range of VRP variants, its current training scheme does not fully account for the inherent disparities in problem complexity.
+In the future, we aim to explore more effective strategies, such as hierarchical training and difficulty-adaptive sampling, to improve training efficiency.
+Additionally, extending our method to accommodate more complex variants(e.g., dynamic VRP) is a promising direction for future research.
+Acknowledgements
+This work was supported by the National Natural Science Foundation of China (Grant No.
+62476118), the Guangdong Provincial Key Laboratory of Fully Actuated System Control Theory and Technology (Grant No.
+2024B1212010002), the Natural Science Foundation of Guangdong Province (Grant No.
+2024A1515011759), and the Center for Computational Science and Engineering at Southern University of Science and Technology.
+We would like to thank the anonymous reviewers and (S)ACs of ICML 2026 for their constructive comments and dedicated service to the community.
+Additionally, we would like to thank Mr.
+Rongsheng Chen for his valuable advice and assistance during this research.
+Impact Statement
+This paper presents work whose goal is to advance the field of Machine Learning.
+There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
 Appendix A Related Work
@@
 Subsequently, the Transformer-style heavy-encoder and light-decoder auto-regressive architecture has become the dominant paradigm (Kool et al., 2019), while POMO (Kwon et al., 2020) exploits permutation symmetry to enhance exploration diversity.
-Many advancements within this paradigm have pushed promising results on small instances (e.g., 100-node TSPs) (Xin et al., 2020; Kim et al., 2022; Xiao et al., 2024a; b; Sun et al., 2024).
+Many advancements within this paradigm have pushed promising results on small instances (e.g., 100-node TSPs) (Xin et al., 2020; Kim et al., 2022; Xiao et al., 2024a, b; Sun et al., 2024).
 However, models trained on a fixed instance scale often degrade sharply when evaluated out of scale.
-To mitigate the poor generalization performance, recent studies introduce auxiliary local policies (Gao et al., 2024), adaptive perception across instance scales (Zhou et al., 2024a), or different search space reduction methods (Fang et al., 2024; Zhou et al., 2025; Chen et al., 2025).
+To mitigate the poor generalization performance, recent studies introduce auxiliary local policies (Gao et al., 2024), adaptive perception across instance scales (Zhou et al., 2026), or different search space reduction methods (Fang et al., 2024; Zhou et al., 2025; Chen et al., 2025).
 In parallel, ”heavy decoder” designs that dynamically recompute embeddings at each construction step (Drakulic et al., 2023; Luo et al., 2023) demonstrate strong generalization on large-scale instances.
-To improve the performance on large-scale routing instances, complementary directions are reducing solving difficulty via problem decomposition (Zheng et al., 2024b; Ye et al., 2024b; Li et al., 2025b) or adopting non-autoregressive generation augmented with additional searches (e.g., Monte Carlo tree search (MCTS) and 2-opt) (Fu et al., 2021; Sun & Yang, 2023; Qiu et al., 2022; Li et al., 2023; Ye et al., 2023).
-Moreover, heterogeneous vehicle routing (Li et al., 2022b), asymmetric distance settings (Kwon et al., 2021), multiple VRPs (Zheng et al., 2024a), and pickup–delivery problems (Li et al., 2021; Kong et al., 2024) have each motivated tailored neural designs, which have yielded competitive results.
-Despite these advancements, they still train a specialized model per problem type, limiting cross-problem transfer.
+To improve the performance on large-scale routing instances, complementary directions are reducing solving difficulty via problem decomposition (Zheng et al., 2024b; Ye et al., 2024; Li et al., 2025b) or adopting non-autoregressive generation augmented with additional searches (e.g., Monte Carlo tree search (MCTS) and 2-opt) (Fu et al., 2021; Sun and Yang, 2023; Qiu et al., 2022; Li et al., 2023; Ye et al., 2023).
+Moreover, heterogeneous vehicle routing (Li et al., 2022), asymmetric distance settings (Kwon et al., 2021), multiple VRPs (Zheng et al., 2024a), and pickup-delivery problems (Li et al., 2021; Kong et al., 2024) have each motivated tailored neural designs, which have yielded competitive results.
+Despite these advancements, they still train a specialist model per problem type, limiting cross-problem transfer.
 We instead focus on a unified neural routing solver that aims to generalize across multiple routing problems without retraining from scratch for each specification.
@@
 Existing efforts can broadly fall into two categories: (1) constraint combination-based multi-task learning and (2) adapter-based fine-tuning frameworks.
-The former line (e.g., MTPOMO (Liu et al., 2024a)) treats VRP variants as combinations of a predefined set of attributes and trains a shared backbone, achieving promising results on up to $16$ VRP variants.
-Follow-up work extends this paradigm by introducing mixture-of-experts (MoE) modules (Zhou et al., 2024b), heavy decoder designs plus knowledge distillation (Zheng et al., 2025a), and constraint-aware dual-attention (Li et al., 2025a) or updated Transformer structure (Berto et al., 2024) to improve model performance.
+The former line (e.g., MTPOMO (Liu et al., 2024)) treats VRP variants as combinations of a predefined set of attributes and trains a shared backbone, achieving promising results on up to $16$ VRP variants.
+Follow-up work extends this paradigm by introducing mixture-of-experts (MoE) modules (Zhou et al., 2024), heavy decoder designs plus knowledge distillation (Zheng et al., 2025), and constraint-aware dual-attention (Li et al., 2025a) or updated Transformer structure (Berto et al., 2025) to improve model performance.
 While this enables knowledge sharing across problems, its coverage is inherently bounded by the hand-specified attribute set.
@@
 In contrast, we focus on a constructive neural routing solver supported with zero-shot capability: once trained, it can directly generate high-quality solutions for previously unseen problems without any fine-tuning.
-A.3 LLM-based Algorithm Generation in VRPs
-Employing LLMs to solve VRPs is a new research direction.
-As an early attempt, Yang et al.
-(2023) utilizes the LLM as a direct optimizer, prompting an LLM to directly generate solutions for TSP instances.
-A more predominant stream leverages LLMs as automated algorithm designers to generate algorithms for solving VRPs.
-These methods often integrate LLMs with a search procedure.
-EoH (Liu et al., 2024b) combined an LLM with Evolutionary Computation to design heuristic components for a Guided Local Search (GLS) framework.
-ReEvo (Ye et al., 2024a) incorporated a reflection mechanism to refine heuristics for iterative solution construction.
-HSEvo (Dat et al., 2025) additionally introduced a diversity mechanism for designing heuristics solving TSP and OP.
-And AHD-MCTS (Zheng et al., 2025b) employed Monte Carlo Tree Search (MCTS) to guide the design of heuristics for solving TSP and CVRP.
-Despite their promise, these methods share common limitations rooted in their reliance on the LLM’s implicit, pre-trained knowledge as the core design logic.
-The process of discovering effective algorithms demands extensive search procedures and is computationally expensive.
-Furthermore, the heuristics generated by the LLM are susceptible to overfitting to the specific instances used during the design phase, limiting their generalization capabilities.
-In contrast, our URS overcomes these challenges by establishing a clear division of labor: the LLM is tasked with satisfying complex constraints, while our URS model is responsible for ensuring solution quality.
-This approach significantly reduces the reasoning burden on the LLM, mitigating the risk of overfitting and enabling URS to effectively solve a more extensive range of VRPs.
 Appendix B Setups of VRP Variants
-In this paper, we consider over 100 VRP variants which may simultaneously have one or more constraints from the following categories: (1) Capacity (C); (2) Open Route (O); (3) Backhaul (B); (4) Backhaul and Priority (BP); (5) Duration Limit (L); (6) Time Windows (TW); (7) Multi-Depot (MD); (8) Prize Collecting (PC); (9) Asymmetric (A); (10) Pickup and Delivery (PD).
+In this paper, we consider 110 VRP variants which may simultaneously have one or more constraints from the following categories: (1) Capacity (C); (2) Open Route (O); (3) Backhaul (B); (4) Backhaul and Priority (BP); (5) Duration Limit (L); (6) Time Windows (TW); (7) Multi-Depot (MD); (8) Prize Collecting (PC); (9) Asymmetric (A); (10) Pickup and Delivery (PD).
 The node coordinates are randomly sampled from the unit square when available (in symmetric geometric cases, i.e., $d_{ij}=d_{ji}$), following AM (Kool et al., 2019).
@@
 The introduction of backhauls extends this formulation to the CVRP with Backhauls (CVRPB), where vehicles must also load goods from backhaul customers and return them to the depot, in addition to serving the standard linehaul (delivery) customers.
-Following the experimental setup of MTPOMO (Liu et al., 2024a), we generate customer demands by sampling from a discrete uniform distribution over the integers $\{1,\ldots,9\}$.
+Following the experimental setup of MTPOMO (Liu et al., 2024), we generate customer demands by sampling from a discrete uniform distribution over the integers $\{1,\ldots,9\}$.
 Subsequently, 20% of these nodes are randomly designated as backhaul customers, and their corresponding demand values are negative, while the remainder are classified as linehaul customers.
-Notably, no precedence constraints are imposed between the servicing of linehaul and backhaul customers.To ensure the feasibility of the solutions constructed by our neural solver, we adhere to the MVMoE (Zhou et al., 2024b) framework’s policy that the first customer visited must be a linehaul node.
+Notably, no precedence constraints are imposed between the servicing of linehaul and backhaul customers.
+To ensure the feasibility of the solutions constructed by our neural solver, we adhere to the MVMoE (Zhou et al., 2024) framework’s policy that the first customer visited must be a linehaul node.
 Backhaul and Priority (BP)
@@
 Duration Limit (L)
-Following MTPOMO (Liu et al., 2024a), we impose a duration limit (i.e., the maximum length of each vehicle route) of 3 in symmetric cases.
+Following MTPOMO (Liu et al., 2024), we impose a duration limit (i.e., the maximum length of each vehicle route) of 3 in symmetric cases.
 Since the node coordinates are randomly sampled from the unit square, the maximum possible distance between any two nodes is $\sqrt{2}$.
@@
 Time Windows (TW)
-Consistent with the MVMoE (Zhou et al., 2024b), we set time windows of the depot to $e_{0},l_{0}]$=[0,3] and its service time to 0.
+Consistent with the MVMoE (Zhou et al., 2024), we set time windows of the depot to $e_{0},l_{0}]$=[0,3] and its service time to 0.
 The service time $s_{i}$ for each customer node is set to 0.2.
-For the time windows for each customer node, we generate them following the methodology outlined in MVMoE (Zhou et al., 2024b).
+For the time windows for each customer node, we generate them following the methodology outlined in MVMoE (Zhou et al., 2024).
 Multi-Depot (MD)
-The generation of coordinates for the multiple depots aligns with that of other nodes, with the number of depots set to 3 (Berto et al., 2024).
+The generation of coordinates for the multiple depots aligns with that of other nodes, with the number of depots set to 3 (Berto et al., 2025).
 A vehicle is required to return to the same depot it departed from.
@@
 A key constraint is that a minimum amount of total prize must be collected before the vehicle is permitted to return to the depot and terminate its route.
+In this context, the minimum prize requirement for the four examined problems (PCTSP, SPCTSP, APCTSP, and ASPCTSP) is standardized to 1.0.
 Consistent with the AM (Kool et al., 2019) framework, the prizes and penalties are generated as follows: $prize\sim\text{Uniform}\left(0,\frac{4}{n}\right)$, and $penalty\sim\text{Uniform}\left(0,3\cdot\frac{k_{n}}{n}\right)$, where $n=100$ and $k_{n}=4$ in our experiments.
+For the OP, the objective is to maximize collected prizes without exceeding a predefined distance limit (set to 4.0 for OP and 1.0 for AOP), where unvisited nodes carry no penalties.
+The prize for each node is assigned proportional to its Euclidean distance from the depot.
+By normalizing distances to the instance’s maximum distance and discretizing them into $[0.01,1.00]$, nodes further from the depot systematically yield higher rewards, with the depot’s prize set to zero.
 Asymmetric (A)
@@
 In this section, we provide a detailed description of our UDR $\bm{U}=\{\mathbf{u}_{i}\}_{i=0}^{n}$ (see Table 9).
-Each node $\mathbf{u}_{i}$ consists of three components, i.e., $\mathbf{u}_{i}=\{\bm{\rho}_{i},\ \bm{\omega}_{i},\ \bm{\xi}_{i}\}$.
-where $\bm{\rho}_{i},\ \bm{\omega}_{i},\ \bm{\xi}_{i}$ are a position identifier, unified attribute set, and node-type indicator, respectively.
-To distinguish linehaul/delivery from backhaul/pickup nodes, we assign negative demand values to backhaul/pickup nodes (positive for linehaul/delivery), making demand the only signed component, following  (Liu et al., 2024a; Zhou et al., 2024b).
+Each node $\mathbf{u}_{i}$ consists of three components, i.e., $\mathbf{u}_{i}=\{\bm{\rho}_{i},\ \bm{\omega}_{i},\ \bm{\xi}_{i}\}$, where $\bm{\rho}_{i},\ \bm{\omega}_{i},\ \bm{\xi}_{i}$ are a positional identifier, unified attribute set, and node-type indicator, respectively.
+To distinguish linehaul/delivery from backhaul/pickup nodes, we assign negative demand values to backhaul/pickup nodes (positive for linehaul/delivery), making demand the only signed component, following  (Liu et al., 2024; Zhou et al., 2024).
 Table 9: Detailed description of each component in UDR.
@@
 As shown in Table 10, we train URS on eleven classic and varied routing problems to ensure that each feature in the UDR has been seen at least once, i.e., $\mathcal{P}=\{P_{i}\}_{i=1}^{11}$.
-Then, we evaluate generalization on unseen problems on different combinations of ten constraint attributes (see Appendix B), and the detailed representations of each unseen problem can be found in Table 7 (documented VRP variants) and Table 8 (undocumented VRP variants).
+We adhere to the principle of minimal redundancy and strictly align with existing work on CVRP variant selection (e.g., MVMoE (Zhou et al., 2024) and ReLD (Huang et al., 2025)): most features in UDR appear in at least two training problems, helping prevent URS from overfitting to a single problem.
+The only exception is the Penalty attribute, as there is structural similarity between the seen PCTSP and unseen SPCTSP.
+For more detailed analysis and experimental results, please see Appendix J.1.
+Then, we evaluate generalization on unseen problems on different combinations of ten constraint attributes (see Appendix B), and the detailed representations of each unseen problem can be found in Table 11 (unseen well-studied VRP variants) and Table 12 (unseen less-explored VRP variants).
 Table 10: The representations for seen routing problems in our experiments.
@@
 PDTSP LKH-3 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-ACVRP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+ACVRP PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
 CVRP PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
 OCVRPTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-Table 7: The problem representations of $44$ VRP variants that have been extensively explored in current literature (Liu et al., 2024a; Berto et al., 2024).
+Frequency 2/11 9/11 6/11 2/11 1/11 2/11 2/11 2/11 9/11 2/11 7/11 6/11 2/11
+Table 11: The problem representations of unseen $44$ well-studied VRP variants.
 Here, RI represents the random identifier, and EAT, LAT, and ST indicate the earliest arrival time, the latest arrival time, and the service time, respectively.
@@
 OCVRPBPTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-CVRPBLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+CVRPBPLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
 CVRPBPTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
 OCVRPBPL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-OCVRPBLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+OCVRPBPLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
 MDCVRP PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
 MDOCVRPBPTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-MDCVRPBLTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-MDCVRPBTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+MDCVRPBPLTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+MDCVRPBPTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
 MDOCVRPBPL PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-MDOCVRPBLTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+MDOCVRPBPLTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
 SPCTSP ILS $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-Table 8: The problem representations of $52$ VRP variants that have not yet been explored in current literature.
+Table 12: The problem representations of unseen $55$ less-explored VRP variants.
 Here, RI represents the random identifier, and EAT, LAT, and ST indicate the earliest arrival time, the latest arrival time, and the service time, respectively.
@@
 AOCVRPBPTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-ACVRPBLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-ACVRPBTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+ACVRPBPLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+ACVRPBPTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
 AOCVRPBPL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-AOCVRPBLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+AOCVRPBPLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
 AMDCVRP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
 AMDOCVRPBPTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-AMDCVRPBLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-AMDCVRPBTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+AMDCVRPBPLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+AMDCVRPBPTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
 AMDOCVRPBPL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-AMDOCVRPBLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+AMDOCVRPBPLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
 APDTSP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+APDCVRP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+AOPDCVRP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
 PDCVRP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
 OPDCVRP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-APDCVRP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-AOPDCVRP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-Appendix E Model Architecture
+AOP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$
+APCTSP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+ASPCTSP N/A $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+Appendix E Model Architecture and Training
 E.1 Encoder
 Similar to previous work (Kool et al., 2019; Kwon et al., 2020), we also use the encoder composed of $L$ stacked attention layers, where each attention layer consists of two sub-layers: an attention sub-layer and a Feed-Forward (FF) sub-layer, both of which use Instance Normalization (Ulyanov et al., 2016) and skip-connection (He et al., 2016).
-Let ${H}^{(\ell-1)}=\{\mathbf{h}_{i}^{(\ell-1)}\}_{i=0}^{n}$ denote the input to the $\ell$-th attention layer for $\ell=1,\ldots,L$. The outputs for the $i$-th node are computed as follows:
+The encoder $\bm{\theta}_{enc}$ with $L$ stacked attention layers transforms $H^{(0)}$ into advanced node embeddings $H^{(L)}=\{\mathbf{h}_{i}^{(L)}\}_{i=0}^{n}$.
+Let ${H}^{(\ell-1)}=\{\mathbf{h}_{i}^{(\ell-1)}\}_{i=0}^{n}$ denote the input to the $\ell$-th attention layer for $\ell=1,\ldots,L$.
+The outputs for the $i$-th node are computed as follows:
 $\widetilde{\mathbf{h}}_{i}^{(\ell)}=\mathrm{IN}^{(\ell)}\left(\mathbf{h}_{i}^{(\ell-1)}+\mathrm{MBM}^{(\ell)}\left(\mathbf{h}_{i}^{(\ell-1)},H^{(\ell-1)},\bm{D},\bm{D}^{\mathrm{T}},\bm{R}\right)\right),$ (11)
@@
 It can be computed via three consecutive linear projections:
-$\bm{H}_{\text{dec}}(\bm{\lambda})=((\bm{\lambda}W_{1}+\mathbf{b}_{1})W_{2}+\mathbf{b}_{2})W_{3}+b_{3},$ (12)
-where $W_{1}\in\mathbb{R}^{|\lambda|\times d_{h}},\ W_{2}\in\mathbb{R}^{d_{h}\times d_{h}},\ W_{3}\in\mathbb{R}^{d_{h}\times(5\times|\bm{\lambda}|)},\ \mathbf{b}_{1}\in\mathbb{R}^{d_{h}},\ \mathbf{b}_{2}\in\mathbb{R}^{d_{h}},\mathbf{b}_{2}\in\mathbb{R}^{(5\times|\bm{\lambda}|)}$ are learnable parameters, $d_{h}$ is $256$ in this work, following Lin et al.
+$\bm{H}_{\text{dec}}(\bm{\lambda})=((\bm{\lambda}W_{1}+\mathbf{b}_{1})W_{2}+\mathbf{b}_{2})W_{3}+\mathbf{b}_{3},$ (13)
+where $W_{1}\in\mathbb{R}^{|\bm{\lambda}|\times d_{h}},\ W_{2}\in\mathbb{R}^{d_{h}\times d_{h}},\ W_{3}\in\mathbb{R}^{d_{h}\times(5\times|\bm{\lambda}|)},\ \mathbf{b}_{1}\in\mathbb{R}^{d_{h}},\ \mathbf{b}_{2}\in\mathbb{R}^{d_{h}},\mathbf{b}_{3}\in\mathbb{R}^{(5\times|\bm{\lambda}|)}$ are learnable parameters, $d_{h}$ is $256$ in this work, following Lin et al.
 (2022).
@@
 $W_{K}(\bm{\lambda})=\bm{h}_{\text{dec}}^{4}(\bm{\lambda})W^{4}_{\textbf{hyper}},\ W_{V}(\bm{\lambda})=\bm{h}_{\text{dec}}^{5}(\bm{\lambda})W^{5}_{\textbf{hyper}},$ (15)
-where $W^{1}_{\textbf{hyper}}\in\mathbb{R}^{|\lambda|\times(d\times d)}$, $W^{2}_{\textbf{hyper}}\in\mathbb{R}^{|\lambda|\times(d\times d)}$, $W^{3}_{\textbf{hyper}}\in\mathbb{R}^{|\lambda|\times(1\times d)}$, $W^{4}_{\textbf{hyper}}\in\mathbb{R}^{|\lambda|\times(d\times d)}$, $W^{5}_{\textbf{hyper}}\in\mathbb{R}^{|\lambda|\times(d\times d)}$ are learnable parameters.
+where $W^{1}_{\textbf{hyper}}\in\mathbb{R}^{|\bm{\lambda}|\times(d\times d)}$, $W^{2}_{\textbf{hyper}}\in\mathbb{R}^{|\bm{\lambda}|\times(d\times d)}$, $W^{3}_{\textbf{hyper}}\in\mathbb{R}^{|\bm{\lambda}|\times(1\times d)}$, $W^{4}_{\textbf{hyper}}\in\mathbb{R}^{|\bm{\lambda}|\times(d\times d)}$, $W^{5}_{\textbf{hyper}}\in\mathbb{R}^{|\bm{\lambda}|\times(d\times d)}$ are learnable parameters.
 In this way, we use a simple MLP hypernetwork $\mathrm{WEIGHT}(\bm{\lambda})$ to generate used linear projection matrices $\{W_{Q}(\bm{\lambda}),\ W_{K}(\bm{\lambda}),\ W_{V}(\bm{\lambda})\}$ conditioned on the multi-hot problem representation $\bm{\lambda}$, enabling the adaptive construction of high-quality solutions for different problems.
@@
 For all problems with the capacity constraint, we explicitly impose only remaining load, while diverse additional constraints (e.g., time windows, duration limit, and backhaul) are enforced implicitly as $\mathcal{M}_{t}$ masks infeasible candidate nodes to guarantee valid solutions, instead of being fully enumerated in the input of the decoder.
-For OP, we keep track of the remaining maximum length at time $t$, and $C_{t}$ represents the remaining routing length that can be moved, and we normalize it to $[0,1]$ by dividing by the maximum length (i.e., 4.0 in OP100).
+For OP and AOP, we keep track of the remaining maximum length at time $t$, and $C_{t}$ represents the remaining routing length that can be moved, and we normalize it to $[0,1]$ by dividing by the maximum length (i.e., 4.0 in OP100 and 1.0 in AOP100).
 The setting is the same as AM (Kool et al., 2019).
-For PCTSP and SPCTSP, $C_{t}$ is the remaining prize to collect, and we do not provide any information about the penalties, as this is irrelevant for the remaining decisions, following Kool et al.
+For (A)PCTSP and (A)SPCTSP, $C_{t}$ is the remaining prize to collect, and we do not provide any information about the penalties, as this is irrelevant for the remaining decisions, following Kool et al.
 (2019).
@@
 $\nabla_{\theta}\mathcal{L}(\bm{\theta})=\mathbb{E}_{p(\bm{\pi}|\mathcal{G},{\bm{\theta}})}[(f(\bm{\pi}|\mathcal{G})-\frac{1}{n}\sum_{i=1}^{n}f(\bm{\pi}^{i}|\mathcal{G}))\nabla_{\bm{\theta}}\log p_{\bm{\theta}}\left(\bm{\pi}|\mathcal{G}\right)],$ (16)
-$p_{\bm{\theta}}(\bm{\pi}\mid\mathcal{G})=\prod_{t=2}^{n}p_{\bm{\bm{\theta}}}(\pi_{t}\mid\mathcal{G},\pi_{1:t-1}),$ (16)
+$p_{\bm{\theta}}(\bm{\pi}\mid\mathcal{G})=\prod_{t=2}^{n}p_{\bm{\theta}}(\pi_{t}\mid\mathcal{G},\pi_{1:t-1}),$ (17)
 where the objective function $f(\bm{\pi}|\mathcal{G})$ represents the total reward (e.g., the negative value of tour length) of instance $\mathcal{G}$ given a specific solution $\bm{\pi}$.
@@
 Following Zhou et al.
-(2024a), we implement $\mathrm{Attention}(\cdot)$ using an adaptation attention free module (AAFM) to enhance geometric pattern recognition for routing problems.
-Given the input $X$, AAFM first transforms it into $Q$, $K$, and $V$ through corresponding linear projection operations:
+(2026), we implement $\mathrm{Attention}(\cdot)$ using an adaptation attention free module (AAFM) to enhance geometric pattern recognition for routing problems.
+Given the input $X$, the AAFM computation is then expressed as:
 $Q=XW^{Q},\quad K=XW^{K},\quad V=XW^{V},$ (18)
-where $W^{Q}$, $W^{K}$, and $W^{V}$ are learnable matrices. The AAFM computation is then expressed as:
 $\mathrm{Attention}(Q,K,V,A)=\sigma(Q)\odot\frac{\exp(A)(\exp(K)\odot V)}{\exp(A)\exp(K)},$ (19)
-where $\sigma$ denotes the sigmoid function, $\odot$ represents the element-wise product, and $A=\{a_{ij}\}$ denotes the pair-wise adaptation bias.
+where $W^{Q}$, $W^{K}$, and $W^{V}$ are learnable matrices.
+$\sigma$ denotes the sigmoid function, $\odot$ represents the element-wise product, and $A=\{a_{ij}\}$ denotes the pair-wise adaptation bias.
 For distance matrices, the corresponding adaptation bias $f\left(\alpha,N,d_{ij}\right)=-\alpha\cdot\log_{2}{N}\cdot d_{i,j}$.
@@
 Compared to multi-head attention (MHA) (Vaswani et al., 2017), AAFM enables the model to explicitly capture instance-specific knowledge by updating pair-wise adaptation biases while exhibiting lower computational overhead.
-Further details are provided in the related work section mentioned above.
-Appendix G LLM-driven Constraint Satisfaction Mechanism
-To operationalize the UDR framework for a vast spectrum of routing problems, we introduce an automated constraint satisfaction mechanism based on LLMs.
-This mechanism obviates the need for manual, problem-specific masks by automatically designing executable mask generators.
-This approach follows the generate-and-validate pipeline, transforming constraint specifications into reliable code.
-G.1 Mask Generator Design
-This process begins by prompting an LLM with a prompt that comprises two key elements: 1) a natural language task description outlining the constraint logic, and 2) a standardized Python code template that specifies the function signature, inputs, and outputs.
-Here is an example for generating a mask for the TSP that enforces the fundamental visit-once constraint:
-Task Description: Given the current node, visited nodes, and the number of nodes, design a function to mask visited nodes.
-Code Template:
-1import torch
-2
-3def mask_visited_nodes(current_node: torch.Tensor, visited_nodes: torch.Tensor, num_nodes: int) -> torch.Tensor:
-4 """
-5 Design a novel algorithm to mask the visited nodes in each step.
-6 The operation processes batches in parallel, so no loops over the batch dimension are allowed-inputs must use batch-first shapes.
-7
-8 Args:
-9 current_node: ID of the current node. Shape: [batch_size, 1]
-10 visited_nodes: Array of IDs of visited nodes. Shape: [batch_size, num_visited_nodes]
-11 num_nodes: The number of nodes.
-12
-13 Return:
-14 mask: Boolean mask indicating valid nodes (True = selectable, False = masked). Shape: [batch_size, num_nodes]
-15 """
-16 return mask
-The LLM completes the function body with the prompt guidance. The candidate mask generator is an executable Python snippet implementing the described logic:
-1import torch
-2
-3def mask_visited_nodes(current_node: torch.Tensor, visited_nodes: torch.Tensor, num_nodes: int) -> torch.Tensor:
-4 """
-5 Design a novel algorithm to mask the visited nodes in each step.
-6 The operation processes batches in parallel, so no loops over the batch dimension are allowed-inputs must use batch-first shapes.
-7
-8 Args:
-9 current_node: ID of the current node. Shape: [batch_size, 1]
-10 visited_nodes: Array of IDs of visited nodes. Shape: [batch_size, num_visited_nodes]
-11 num_nodes: The number of nodes.
-12
-13 Return:
-14 mask: Boolean mask indicating valid nodes (True = selectable, False = masked). Shape: [batch_size, num_nodes]
-15 """
-16 mask = torch.ones(current_node.size(0), num_nodes, dtype=torch.bool)
-17 mask.scatter_(1, visited_nodes, 0)
-18 mask.scatter_(1, current_node, 0)
-19 return mask
-G.2 Automated Validation
-Each candidate generator is then validated.
-We execute it within the URS solver or a simple simulated solver on a set of problem instances and use a problem-specific checker function to verify if the resulting solutions are constraint-compliant.
-For the TSP example, the checker ensures that every node is visited exactly once:
-1import torch
-2
-3def checker(solutions: torch.Tensor, problem_size: int, num_instances: int) -> float:
-4 """
-5 Check the validity of TSP solutions that use the generated masks.
-6
-7 Args:
-8 solutions: A batch of generated solutions to be validated. Each row represents a single solution sequence. Shape: [batch_size, problem_size].
-9 problem_size: The problem size.
-10
-11 Returns:
-12 valid_rate: The validity rate, representing the fraction of valid solutions in the batch.
-13 """
-14 num_instances = solutions.shape[0]
-15 sorted, _ = torch.sort(solutions, dim=1)
-16 valid = (sorted == torch.arange(problem_size)[None, :].expand(num_instances, problem_size)).all(1)
-17 valid_rate = torch.mean(valid.float()).item()
-18 return valid_rate
-G.3 AHD Framework Integration
-The mask generator design process can be formulated as an optimization problem aimed at maximizing the solution validity rate across test instances.
-This allows us to leverage sophisticated LLM-based automated algorithm design frameworks, such as EoH (Liu et al., 2024b), to improve the efficiency, especially for the complex VRP variants.
-The search terminates as soon as a fully reliable generator is found, making the process significantly more efficient than general optimization tasks.
-This LLM-driven synthesis, combined with UDR, empowers URS with true zero-shot generalization capabilities while rigorously upholding solution feasibility across an unprecedented range of routing problems.
+Further details are provided in the related work mentioned above.
+Appendix G Comparison between Different Attention Mechanisms
+Table 13: Comparison between different attention mechanisms.
+Here, $N,d,m$ denote the sequence length, feature dimension, and extra hidden layer dimension of mixed-score attention, respectively.
+Values in parentheses represent the percentage increase or decrease relative to URS and Avg.
+Peak VRAM represents the average peak VRAM per instance.
+Attention Mechanism Time Complexity Space Complexity #Params Avg.Peak VRAM Inference Time
+Multi-Head Attention $\mathcal{O}(N^{2}d)$ $\mathcal{O}(N^{2}+Nd)$ 3.323.32M(×0.7)($\times$ 0.7) 1.871.87MB(×0.99)($\times$ 0.99) $0.92$s(×0.9)($\times$ 0.9)
+Mixed-Score Attention (Parallel) $\mathcal{O}(N^{2}(d+m))$ $\mathcal{O}(N^{2}m+Nd)$ 16.1716.17M(×3.3)($\times$ 3.3) 25.6425.64MB(×13.6)($\times$ 13.6) $3.64$s(×3.7)($\times$ 3.7)
+Mixed Bias Module (Ours) $\mathcal{O}(N^{2}d)$ $\mathcal{O}(N^{2}+Nd)$ 4.964.96M 1.891.89MB $0.97s$
+To validate the efficiency of our proposed MBM, we have conducted a new comparative experiment between different attentions and tested them on TSP instances to ensure consistent decoding steps.
+Specifically, we replace our MBM with both a vanilla MHA (Vaswani et al., 2017) and the parallel mixed-score attention (MSA) used in MatNet (Kwon et al., 2021), and compare their complexities, memory overheads, and inference time.
+The results in Table 13 show that the increased VRAM and time overhead of MBM compared to standard MHA are small, and both maintain the same time complexity and space complexity.
+Compared to MatNet’s parallel MSA, our MBM drastically reduces both VRAM consumption and inference latency.
+These results show that MBM imposes only modest computational penalties while successfully capturing crucial problem-specific biases, which enhances cross-problem generalization while reducing architectural redundancy.
 Appendix H Training and Model Settings
-The detailed information about the hyperparameter settings can be found in Table 9.
-The training of URS is conducted on a single NVIDIA GeForce RTX 4090 GPU (24GB memory).
-The training procedure requires only about $95$ hours ($\approx 4$ days) in total (500 epochs at approximately $12$ minutes each) to obtain a unified parameter set capable of addressing a broad class of routing problems.
+H.1 Multi-Task Version
 Table 14: The hyperparameter settings of URS.
@@
 Training Epochs $500$
-Appendix I Detailed Experimental Results
-To comprehensively evaluate zero-shot generalization of URS, we test its performance on a diverse set of 96 unseen problems.
-This set is designed to cover both established and novel challenges, including 44 well-studied VRP variants from the NCO literature and 52 newly introduced variants that have not been explored before.
+The detailed information about the hyperparameter settings can be found in Table 14.
+The URS training is conducted on a single NVIDIA GeForce RTX 4090 GPU (24GB of memory).
+The training procedure requires only about $95$ hours ($\approx 4$ days) in total (500 epochs at approximately $12$ minutes each) to obtain a unified parameter set capable of addressing a broad class of routing problems.
+H.2 Single-Task Version
+URS-STL is trained on data from only a single problem type (e.g., ”URS-STL (CVRP)” is trained only on CVRP).
+The URS-STL variant uses exactly the same model architecture, hyperparameters (e.g., embedding dimension, number of layers, optimizer), and training setup (e.g., initial learning rate, batch size) as the multi-task URS described in Section 4.1 and Appendix H.1.
+The only modification we made is to the training duration.
+We observe that the single-task models converge significantly faster than the multi-task model.
+Therefore, we train URS-STL models for 300 epochs, with a learning rate decay applied at the 251st epoch.
+This contrasts with the multi-task URS, which is trained for 500 epochs with decay applied at the 451st epoch.
+This adjustment is made solely to reflect the faster convergence of the single-task training.
+All other hyperparameters and model architecture remain identical.
+Appendix I Generalization to Unseen VRPs
+To comprehensively evaluate URS’s zero-shot generalization, we test its performance on a diverse set of 99 unseen problems.
+This set is designed to cover both established and novel challenges, including 44 well-studied VRP variants and 55 less-explored VRP variants.
 The detailed results are presented in Table 15 and Table 16.
 On widely-studied constraints (i.e., C, O, L, TW, MD, B, BP), URS continues to deliver high-quality solutions.
-Notably, URS merely selects the next node from a given feasible candidates without any additional domain knowledge, making its performance especially meaningful.
-Benefiting from our UDR, URS can directly address previously unexplored variants, which compound asymmetry and PD relations.
-Without any fine-tuning, URS still constructs high-quality solutions for them, strengthening the broad problem generality and deployment efficiency.
-Table 10: Zero-shot generalization performance across $44$ VRP variants that have been explored in current literature (Liu et al., 2024a; Berto et al., 2024) ($N=100$).
+Notably, URS merely selects the next node from a given set of feasible candidates without any additional domain knowledge, making its performance especially meaningful.
+Benefiting from our UDR, URS can directly address 55 less-explored VRP variants that compound asymmetry and PD relations.
+Without any fine-tuning, URS still produces high-quality solutions for them, thereby strengthening its broad problem generality and deployment efficiency.
+Table 15: Zero-shot generalization performance across $44$ unseen well-studied VRP variants ($N=100$).
 Method CVRPBP OCVRPBP CVRPBPL OCVRPBPTW CVRPBPLTW
 Gap Time Gap Time Gap Time Gap Time Gap Time
-Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
-MVMoE 13.95% 14s 17.36% 14s 13.38% 14s 2.29% 15s 10.46% 16s
-MTPOMO 13.71% 7s 16.94% 8s 13.44% 8s 3.26% 9s 11.35% 10s
-ReLD-MTL 13.57% 9s 14.83% 9s 12.96% 10s 2.27% 11s 10.80% 12s
-[…]
+Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
+MVMoE 13.95% 14s 18.27% 14s 13.38% 14s 9.42% 15s 16.66% 16s
+MTPOMO 13.71% 7s 17.90% 8s 13.44% 8s 9.94% 9s 17.15% 10s
+ReLD-MTL 13.57% 9s 15.81% 9s 12.96% 10s 8.43% 11s 15.95% 12s
+URS 12.95% 7s 16.96% 7s 12.65% 8s 10.05% 8s 18.33% 10s
 Method CVRPBPTW OCVRPBPL OCVRPBPLTW CVRPL CVRPBL
 Gap Time Gap Time Gap Time Gap Time Gap Time
-Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 16m 0.00% 3.5h
-MVMoE 10.85% 15s 16.88% 14s 1.84% 16s 0.26% 12s 1.35% 12s
-MTPOMO 11.53% 9s 16.39% 8s 2.77% 10s 0.48% 6s 1.79% 6s
-[…]
-[…]
-Method OCVRPB CVRPLTW OCVRPBTW CVRPBLTW OCVRPL
-Gap Time Gap Time Gap Time Gap Time Gap Time
-Oracle Solver 0.00% 3.5h 0.00% 3.5h 0.00% 3.5h 0.00% 3.5h 0.00% 3.5h
-MVMoE 7.09% 12s 1.47% 16s 9.95% 15s 7.33% 16s 3.15% 13s
-MTPOMO 7.34% 5s 1.92% 9s 10.45% 6s 7.75% 8s 3.44% 6s
-[…]
-URS 9.35% 6s 2.67% 9s 13.77% 7s 9.18% 8s 3.22% 7s
+Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 16m 0.00% 3.5h
+MVMoE 16.86% 15s 17.75% 14s 8.87% 16s 0.26% 12s 1.35% 12s
+MTPOMO 17.22% 9s 17.31% 8s 9.47% 10s 0.48% 6s 1.79% 6s
+ReLD-MTL 16.09% 11s 15.21% 10s 8.06% 11s 0.02% 8s 1.01% 7s
+URS 18.15% 9s 16.37% 8s 9.51% 9s 0.43% 7s 1.65% 6s
 Method CVRPBTW OCVRPBL OCVRPLTW OCVRPBLTW MDCVRPBP
 Gap Time Gap Time Gap Time Gap Time Gap Time
-Oracle Solver 0.00% 3.5h 0.00% 3.5h 0.00% 3.5h 0.00% 3.5h 0.00% 6.6m
-MVMoE 7.08% 15s 7.12% 12s 3.90% 15s 10.01% 15s 64.91% 33s
-MTPOMO 7.41% 7s 7.34% 7s 4.37% 8s 10.50% 7s 45.66% 24s
-[…]
-[…]
+Oracle 0.00% 3.5h 0.00% 3.5h 0.00% 3.5h 0.00% 3.5h 0.00% 6.6m
+MVMoE 7.08% 15s 7.12% 12s 3.90% 15s 10.01% 15s 65.98% 33s
+MTPOMO 7.41% 7s 7.34% 7s 4.37% 8s 10.50% 7s 45.77% 24s
+ReLD-MTL 6.74% 8s 5.41% 7s 3.16% 9s 9.22% 8s 40.70% 27s
+URS 8.94% 8s 9.47% 7s 5.12% 9s 13.72% 8s 35.44% 24s
 Method MDOCVRPBP MDCVRPBPL MDOCVRPBPTW MDCVRPBPLTW MDCVRPBPTW
 Gap Time Gap Time Gap Time Gap Time Gap Time
-Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
-MVMoE 52.16% 33s 64.79% 36s 46.18% 38s 54.23% 42s 54.76% 39s
-MTPOMO 40.15% 24s 45.84% 26s 34.19% 29s 40.26% 32s 40.32% 30s
-ReLD-MTL 34.53% 27s 39.06% 29s 34.08% 34s 36.59% 37s 37.16% 35s
-[…]
+Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
+MVMoE 56.60% 33s 65.81% 36s 63.77% 38s 68.49% 42s 69.00% 39s
+MTPOMO 42.94% 24s 45.98% 26s 47.09% 29s 50.45% 32s 50.74% 30s
+ReLD-MTL 37.74% 27s 39.57% 29s 45.16% 34s 45.23% 37s 45.62% 35s
+URS 24.44% 22s 35.07% 26s 26.31% 26s 36.49% 32s 36.71% 30s
 Method MDOCVRPBPL MDOCVRPBPLTW MDCVRP MDCVRPTW MDOCVRP
 Gap Time Gap Time Gap Time Gap Time Gap Time
-Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
-MVMoE 51.72% 34s 45.23% 40s 31.90% 31s 53.64% 41s 29.08% 31s
-MTPOMO 39.34% 27s 33.37% 32s 23.75% 23s 36.10% 31s 27.46% 23s
-[…]
-[…]
+Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
+MVMoE 56.26% 34s 62.47% 40s 33.06% 31s 53.64% 41s 29.08% 31s
+MTPOMO 41.98% 27s 46.26% 32s 23.38% 23s 36.10% 31s 27.46% 23s
+ReLD-MTL 37.45% 29s 44.74% 36s 18.39% 25s 30.90% 34s 22.32% 25s
+URS 24.22% 24s 26.20% 28s 15.05% 23s 37.26% 32s 22.01% 23s
 Method MDCVRPL MDCVRPB MDOCVRPTW MDOCVRPB MDCVRPBL
 Gap Time Gap Time Gap Time Gap Time Gap Time
-Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
+Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
 MVMoE 33.42% 34s 21.70% 26s 51.57% 40s 28.21% 26s 20.35% 29s
 MTPOMO 24.32% 25s 12.89% 19s 35.71% 29s 25.06% 20s 11.10% 20s
-[…]
-[…]
+ReLD-MTL 18.48% 28s 9.38% 21s 33.55% 33s 18.50% 21s 8.26% 23s
+URS 15.32% 26s 13.14% 19s 26.05% 25s 15.17% 18s 11.19% 21s
 Method MDCVRPLTW MDOCVRPBTW MDCVRPBLTW MDOCVRPL MDCVRPBTW
 Gap Time Gap Time Gap Time Gap Time Gap Time
-Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
+Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
 MVMoE 55.55% 44s 62.47% 35s 63.18% 37s 29.33% 33s 65.90% 35s
 MTPOMO 37.34% 32s 45.55% 25s 45.07% 27s 27.21% 25s 45.94% 25s
-[…]
-[…]
+ReLD-MTL 31.56% 36s 43.36% 28s 40.57% 30s 22.05% 27s 41.28% 28s
+URS 38.04% 35s 30.32% 22s 37.70% 27s 22.15% 25s 37.78% 25s
 Method MDOCVRPBL MDOCVRPLTW MDOCVRPBLTW SPCTSP -
 Gap Time Gap Time Gap Time Gap Time
-Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 1.6h -
+Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 1.6h -
 MVMoE 28.70% 28s 51.09% 43s 62.10% 37s - -
@@
 ReLD-MTL 18.50% 22s 33.26% 35s 43.27% 30s - -
-AM - - - 2.69% 5s -
-[…]
-Table 11: Zero-shot generalization performance across $52$ VRP variants that have not yet been explored in the current literature ($N=100$).
+URS 14.90% 20s 25.59% 27s 30.41% 24s -2.37% 5s -
+Table 16: Zero-shot generalization performance across $55$ unseen less-explored VRP variants ($N=100$).
 Method ACVRPL ACVRPBL ACVRPLTW ACVRPBLTW AOCVRPL
 Gap Time Gap Time Gap Time Gap Time Gap Time
-Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
-[…]
+Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
+URS -2.24% 1.9m 6.18% 2.4m 12.22% 2.7m 10.18% 2.4m 37.36% 3.1m
 Method AOCVRPBL AOCVRPLTW AOCVRPBLTW ACVRPBPL ACVRPBPLTW
 Gap Time Gap Time Gap Time Gap Time Gap Time
-Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
-[…]
+Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
+URS 17.23% 2.7m 16.35% 2.8m 14.32% 2.5m 24.03% 2.6m 13.82% 2.8m
 Method AOCVRPBPL AOCVRPBPLTW ACVRPB ACVRPTW AOCVRP
 Gap Time Gap Time Gap Time Gap Time Gap Time
-Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
-[…]
+Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
+URS 32.50% 2.7m 16.43% 2.7m 7.10% 2.1m 11.28% 2.5m 37.63% 2.9m
 Method AOCVRPTW AOCVRPB AOCVRPBTW ACVRPBTW ACVRPBP
 Gap Time Gap Time Gap Time Gap Time Gap Time
-Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
-[…]
+Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
+URS 17.47% 2.6m 18.05% 2.5m 14.45% 2.3m 9.95% 2.2m 24.20% 2.4m
 Method AOCVRPBP AOCVRPBPTW ACVRPBPTW APDTSP AMDCVRPL
 Gap Time Gap Time Gap Time Gap Time Gap Time
-Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
-[…]
+Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
+URS 32.58% 2.5m 17.45% 2.5m 14.52% 2.6m 6.21% 1m 7.15% 6.1m
 Method AMDCVRPBL AMDCVRPLTW AMDCVRPBLTW AMDOCVRPL AMDOCVRPBL
 Gap Time Gap Time Gap Time Gap Time Gap Time
-Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
-[…]
+Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
+URS 17.51% 6.5m 19.98% 8.8m 16.73% 7.5m 47.80% 8.6m 27.23% 7.1m
 Method AMDOCVRPLTW AMDOCVRPBLTW AMDCVRPBPL AMDCVRPBPLTW AMDOCVRPBPL
 Gap Time Gap Time Gap Time Gap Time Gap Time
-Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
-[…]
+Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
+URS 24.90% 7.9m 22.84% 6.7m 35.00% 7.2m 20.50% 8.3m 38.78% 7.3m
 Method AMDOCVRPBPLTW AMDCVRP AMDCVRPTW AMDOCVRP AMDCVRPB
 Gap Time Gap Time Gap Time Gap Time Gap Time
-Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
-[…]
+Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
+URS 24.05% 7.5m 8.11% 5.3m 20.96% 8.1m 48.02% 7.8m 17.35% 5.8m
 Method AMDOCVRPTW AMDOCVRPB AMDOCVRPBTW AMDCVRPBTW AMDCVRPBP
 Gap Time Gap Time Gap Time Gap Time Gap Time
-Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
-[…]
+Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
+URS 26.02% 7.3m 27.78% 6.5m 23.67% 6.2m 17.52% 6.8m 36.18% 6.5m
 Method AMDOCVRPBP AMDOCVRPBPTW AMDCVRPBPTW PDCVRP OPDCVRP
 Gap Time Gap Time Gap Time Gap Time Gap Time
-Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
-[…]
-Method APDCVRP AOPDCVRP - - -
-Gap Time Gap Time
-Oracle Solver 0.00% 6.6m 0.00% 6.6m - - -
-[…]
-Appendix J Results on Benchmark Dataset
-Our test results on Benchmark datasets are presented in Table 12.
+Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
+URS 39.31% 6.7m 24.99% 6.9m 20.98% 7.6m -1.47% 4.3s 4.93% 4.3s
+Method APDCVRP AOPDCVRP AOP APCTSP ASPCTSP
+Gap Time Gap Time Gap Time Gap Time Value Time
+Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m -
+URS 7.03% 1.2m 11.13% 1.2m -5.47% 1.2m 43.79% 1.2m 2.22 1.2m
+Appendix J More Experiments
+J.1 Training Problem Selection
+To empirically validate our selection strategy of training problems, we train two additional variants and compare them with our original URS: URS-Diff (Different Selection) and URS-Expand (Broader Problems).
+The first variant, URS-Diff, uses a different problem set where OCVRPTW is replaced by attribute-independent CVRPL.
+This modification means that the Open Route and Time Window attributes each appear in only one training problem.
+The second variant, URS-Expand, is trained on a broader set of 22 problems, incorporating all variants evaluated in MTPOMO (Liu et al., 2024) and MVMoE (Zhou et al., 2024).
+As shown in Table 17, URS-Diff performs significantly worse on unseen problems involving Open Route and Time Window constraints.
+Meanwhile, URS-Expand achieves a slightly better performance than URS due to seeing more problem variants during training.
+However, the original URS remains highly competitive.
+Remarkably, despite being trained on only half the number of problem types (11 vs.
+22), URS achieves the best solution on 12 out of the 22 evaluated problems.
+These results demonstrate that the choice of our selected 11 problems offers a balanced trade-off, providing sufficient attribute coverage for strong generalization while maintaining training efficiency.
+Expanding the training set yields only marginal gains, while a poorly selected set leads to poor cross-problem generalization.
+In future work, we plan to explore more effective training problem selection strategies to further enhance model performance and training efficiency.
+Table 17: Comparison under different training problem settings.
+The symbol * indicates the original problem selection, and $\Delta$ represents a different problem selection of equal number.
+All test problems are encompassed within URS-Expand.
+Method ATSP (* $\Delta$) TSP(* $\Delta$) OP(* $\Delta$) PCTSP(* $\Delta$) PDTSP(* $\Delta$) ACVRP($\Delta$)
+URS-Diff 4.52% 1.05% 0.82% 1.31% 6.09% 5.10%
+URS-Expand 4.77% 0.93% 0.99% 1.41% 5.85% 5.48%
+URS 2.26% 0.57% 0.45% 1.06% 4.98% 3.06%
+CVRP(* $\Delta$) CVRPTW(* $\Delta$) CVRPB(* $\Delta$) OCVRP(* $\Delta$) OCVRPTW(*) CVRPL($\Delta$)
+URS-Diff 1.97% 6.26% 1.80% 3.75% 15.45% 0.58%
+URS-Expand 2.04% 5.62% 1.77% 3.45% 4.34% 0.65%
+URS 1.81% 6.13% 1.46% 3.24% 5.07% 0.43%
+OCVRPB CVRPBL CVRPLTW OCVRPBTW CVRPBLTW OCVRPL
+URS-Diff 7.55% 2.02% 2.76% 27.48% 10.25% 3.73%
+URS-Expand 5.39% 1.88% 2.03% 9.13% 6.29% 3.41%
+URS 9.35% 1.65% 2.67% 13.77% 9.18% 3.22%
+CVRPBTW OCVRPBL OCVRPLTW OCVRPBLTW Avg.gap Best Sol.
+URS-Diff 10.00% 7.56% 15.61% 27.80% 7.43% 0/22
+URS-Expand 6.07% 5.47% 4.38% 9.23% 4.12% 10/22
+URS 8.94% 9.47% 5.12% 13.72% 4.89% 12/22
+J.2 Comparison with RouteFinder and CaDA
+We strictly adhered to the problem settings established by MVMoE (Zhou et al., 2024) and ReLD (Huang et al., 2025) to ensure a fair and consistent evaluation.
+Notably, as shown in Table 6, URS already outperforms RF (Berto et al., 2025) and CaDA (Li et al., 2025a) on the large-scale CVRPLib benchmark dataset.
+In our cross-problem test, RF and CaDA are excluded due to discrepancies in the problems, including the number of trained CVRP variants (16 in their work vs.
+5 in ours) and constraint settings (i.e., Backhauls and Time Windows).
+To ensure a fair comparison with RF and CaDA, we conduct two additional experiments:
+  • We evaluate URS against CaDA and the best-performing RouteFinder variant (RF-TE) on a shared subset of both seen and unseen problems, strictly restricting the evaluation to variants with perfectly aligned constraint definitions across all methods.
+As shown in Table 18, when evaluated under strictly matched conditions, URS still performs best overall among comparable methods.
+Note that URS can provide the strongest zero-shot generalization across all shared unseen problems.
+  • We retrain both RouteFinder-TE (denoted as RF-Retrain) and CaDA (denoted as CaDA-Retrain) from scratch using the exact same training settings as our URS and evaluate the three models on 48 CVRP variants.
+Since RF fails to train beyond 150 epochs due to the occurrence of NaN probabilities in our training setting, we evaluate all three methods at exactly 150 epochs to ensure a fair comparison.
+The experimental results in Table 19 show that under consistent training settings, URS still achieves the best overall performance among the three methods and provides the best solutions for the majority of problems (32/48), highlighting our strong zero-shot generalization.
+Table 18: Comparison of shared seen and unseen VRP variants between URS, RF-TE, and CaDA.
+Method Seen Unseen Avg.gap Best Sol.
+CVRP OCVRP MDCVRP MDOCVRP MDCVRPL MDOCVRPL
+RF-TE 2.06% 2.86% 63.60% 33.52% 66.75% 34.40% 33.86% 0/6
+CaDA 2.17% 2.80% 40.96% 53.21% 42.40% 53.34% 32.48% 1/6
+URS 1.81% 3.24% 15.04% 22.01% 15.32% 22.15% 13.26% 5/6
+Table 19: Comparison between URS and the retrained RouteFinder and CaDA on all 48 CVRP variants. Note that all models are trained for 150 epochs to ensure a fair comparison.
+Method CVRPTW OCVRP CVRP CVRPB OCVRPTW
+Gap Time Gap Time Gap Time Gap Time Gap Time
+RF-Retrain 8.38% 8s 6.03% 6s 3.58% 6s 4.47% 5s 8.18% 7s
+CaDA-Retrain 5.28% 9s 3.22% 7s 1.92% 7s 1.61% 6s 4.42% 8s
+URS 6.09% 8s 3.71% 6s 2.07% 6s 1.90% 5s 4.89% 8s
+Method OCVRPB CVRPBL CVRPLTW OCVRPBTW CVRPBLTW
+Gap Time Gap Time Gap Time Gap Time Gap Time
+RF-Retrain 11.61% 5s 4.73% 6s 4.77% 9s 14.17% 7s 10.11% 8s
+CaDA-Retrain 8.46% 6s 4.42% 7s 4.63% 11s 10.39% 8s 10.87% 10s
+URS 8.01% 6s 2.12% 6s 2.53% 9s 12.37% 7s 8.74% 8s
+Method OCVRPL CVRPBTW OCVRPBL OCVRPLTW OCVRPBLTW
+Gap Time Gap Time Gap Time Gap Time Gap Time
+RF-Retrain 5.98% 7s 9.74% 7s 11.58% 6s 8.34% 8s 14.46% 7s
+CaDA-Retrain 5.81% 8s 7.17% 8s 12.48% 7s 8.55% 11s 15.94% 10s
+URS 3.72% 7s 8.42% 7s 8.05% 6s 5.06% 8s 12.37% 8s
+Method CVRPBP OCVRPBP CVRPBPL OCVRPBPTW CVRPBPLTW
+Gap Time Gap Time Gap Time Gap Time Gap Time
+RF-Retrain 14.40% 7s 18.31% 7s 14.01% 7s 11.43% 9s 18.70% 10s
+CaDA-Retrain 13.37% 8s 18.25% 8s 16.71% 9s 9.64% 10s 20.17% 12s
+URS 13.37% 7s 16.08% 7s 12.94% 8s 10.07% 9s 18.47% 10s
+Method CVRPBPTW OCVRPBPL OCVRPBPLTW CVRPL MDCVRP
+Gap Time Gap Time Gap Time Gap Time Gap Time
+RF-Retrain 18.84% 9s 17.76% 8s 10.88% 9s 2.21% 7s 27.10% 18s
+CaDA-Retrain 17.11% 10s 21.44% 9s 13.54% 12s 2.29% 8s 27.09% 23s
+URS 18.63% 9s 15.42% 8s 9.57% 10s 0.66% 7s 38.17% 23s
+Method MDCVRPTW MDOCVRP MDCVRPL MDCVRPB MDOCVRPTW
+Gap Time Gap Time Gap Time Gap Time Gap Time
+RF-Retrain 76.46% 25s 19.09% 18s 27.41% 34s 21.40% 15s 35.23% 22s
+CaDA-Retrain 51.08% 30s 30.01% 19s 27.59% 21s 16.47% 19s 54.46% 29s
+URS 38.25% 31s 21.34% 21s 39.22% 26s 20.83% 18s 28.40% 22s
+Method MDOCVRPB MDCVRPBL MDCVRPLTW MDOCVRPBTW MDCVRPBLTW
+Gap Time Gap Time Gap Time Gap Time Gap Time
+RF-Retrain 17.84% 15s 20.02% 29s 79.33% 40s 44.11% 19s 82.19% 23s
+CaDA-Retrain 26.72% 16s 15.02% 18s 43.43% 29s 64.29% 25s 50.11% 24s
+URS 15.20% 17s 17.36% 21s 39.98% 33s 25.68% 19s 33.24% 27s
+Method MDOCVRPL MDCVRPBTW MDOCVRPBL MDOCVRPLTW MDOCVRPBLTW
+Gap Time Gap Time Gap Time Gap Time Gap Time
+RF-Retrain 19.02% 19s 84.94% 21s 17.72% 16s 35.06% 24s 44.12% 21s
+CaDA-Retrain 30.61% 20s 61.39% 25s 26.94% 17s 39.49% 29s 49.12% 25s
+URS 21.56% 24s 33.58% 25s 14.84% 19s 28.60% 24s 25.84% 21s
+Method MDCVRPBP MDOCVRPBP MDCVRPBPL MDOCVRPBPTW MDCVRPBPLTW
+Gap Time Gap Time Gap Time Gap Time Gap Time
+RF-Retrain 54.74% 21s 36.08% 21s 54.19% 23s 47.24% 26s 84.51% 30s
+CaDA-Retrain 54.45% 25s 52.38% 22s 50.69% 24s 69.07% 32s 55.63% 32s
+URS 41.49% 24s 24.18% 23s 41.08% 27s 24.19% 25s 33.99% 35s
+Method MDCVRPBPTW MDOCVRPBPL MDOCVRPBPLTW Avg.gap Best Solutions
+Gap Time Gap Time Gap Time
+RF-Retrain 85.12% 28s 36.06% 23s 47.63% 28s 29.24% 3/48
+CaDA-Retrain 66.78% 33s 53.48% 23s 53.06% 31s 28.45% 13/48
+URS 34.55% 32s 24.10% 25s 24.32% 27s 19.01% 32/48
+J.3 Detailed Results on CVRPLib Set-X Dataset
+We further evaluate the generalization ability of URS on benchmark instances from CVRPLib Set-X (Uchoa et al., 2017) with scale $\in[500,1000]$.
+Our test results on CVRPLib Set-X datasets are presented in Table 20.
 All models are trained on instances of size $N=100$.
-Some results for comparative models are sourced from MVMoE (Zhou et al., 2024b) and RouteFinder (Berto et al., 2024).
-URS maintains its position as the best overall performance across instances of varying scales.
-Notably, RL-based URS surpasses the SL-based representative specialized model LEHD (8.678% vs.
-12.836%), underscoring its practical applicability in real-world scenarios.
+In the Set-X dataset, some results for comparative models are sourced from MVMoE (Zhou et al., 2024) and RouteFinder (Berto et al., 2025).
+As shown in Table 20, URS maintains its position as the best overall performance across instances of varying scales.
+Notably, RL-based URS surpasses the SL-based representative specialist model LEHD (8.678% vs.
+12.836%) and many representative multi-task neural solvers (e.g., MVMoE and RouteFinder), underscoring its practical applicability in real-world scenarios.
 Table 20: Results on large-scale CVRPLib instances (Set-X) (Uchoa et al., 2017) ($500\leq N\leq 1000$). All models are trained on instances of size $N=100$.
@@
 Instance Opt. Obj. Gap Obj. Gap Obj. Gap Obj. Gap Obj. Gap Obj. Gap Obj. Gap Obj. Gap
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
+X-n502-k39 69226 75617 9.232% 71438 3.195% 77284 11.640% 73533 6.222% 74429 7.516% 76338 10.274% 71791 3.705% 71281 2.969%
+X-n513-k21 24201 30518 26.102% 25624 5.880% 28510 17.805% 32102 32.647% 31231 29.048% 32639 34.866% 28465 17.619% 26166 8.119%
+X-n524-k153 154593 201877 30.586% 280556 81.480% 192249 24.358% 186540 20.665% 182392 17.982% 170999 10.612% 174381 12.800% 175250 13.362%
+X-n536-k96 94846 106073 11.837% 103785 9.425% 106514 12.302% 109581 15.536% 108543 14.441% 105847 11.599% 103272 8.884% 102969 8.564%
+X-n548-k50 86700 103093 18.908% 90644 4.549% 94562 9.068% 95894 10.604% 95917 10.631% 104289 20.287% 100956 16.443% 89768 3.539%
+X-n561-k42 42717 49370 15.575% 44728 4.708% 47846 12.007% 56008 31.114% 51810 21.287% 53383 24.969% 49454 15.771% 45964 7.601%
+X-n573-k30 50673 83545 64.871% 53482 5.543% 60913 20.208% 59473 17.366% 57042 12.569% 61524 21.414% 55952 10.418% 54361 7.278%
+X-n586-k159 190316 229887 20.792% 232867 22.358% 208893 9.761% 215668 13.321% 214577 12.748% 212151 11.473% 205575 8.018% 202645 6.478%
+X-n599-k92 108451 150572 38.839% 115377 6.386% 120333 10.956% 128949 18.901% 125279 15.517% 126578 16.714% 116560 7.477% 114423 5.507%
+X-n613-k62 59535 68451 14.976% 62484 4.953% 67984 14.192% 82586 38.718% 74945 25.884% 73456 23.383% 67267 12.987% 65901 10.693%
+X-n627-k43 62164 84434 35.825% 67568 8.693% 73060 17.528% 70987 14.193% 70905 14.061% 70414 13.271% 67572 8.700% 66499 6.973%
+X-n641-k35 63682 75573 18.672% 68249 7.172% 72643 14.071% 75329 18.289% 72655 14.090% 71975 13.023% 70831 11.226% 67005 5.218%
+X-n655-k131 106780 127211 19.134% 117532 10.069% 116988 9.560% 117678 10.206% 118475 10.952% 119057 11.497% 112202 5.078% 110237 3.237%
+X-n670-k130 146332 208079 42.197% 220927 50.977% 190118 29.922% 197695 35.100% 183447 25.364% 168226 14.962% 168999 15.490% 184010 25.748%
+X-n685-k75 68205 79482 16.534% 72946 6.951% 80892 18.601% 97388 42.787% 89441 31.136% 82269 20.620% 77847 14.137% 75942 11.344%
+X-n701-k44 81923 97843 19.433% 86327 5.376% 92075 12.392% 98469 20.197% 94924 15.870% 90189 10.090% 89932 9.776% 86038 5.023%
+X-n716-k35 43373 51381 18.463% 46502 7.214% 52709 21.525% 56773 30.895% 52305 20.593% 52250 20.467% 49669 14.516% 46496 7.200%
+X-n733-k159 136187 159098 16.823% 149115 9.493% 161961 18.925% 178322 30.939% 167477 22.976% 156387 14.833% 148463 9.014% 147743 8.485%
+X-n749-k98 77269 87786 13.611% 83439 7.985% 90582 17.229% 100438 29.985% 94497 22.296% 92147 19.255% 85171 10.227% 83759 8.399%
+X-n766-k71 114417 135464 18.395% 131487 14.919% 144041 25.891% 152352 33.155% 136255 19.086% 130505 14.061% 129935 13.563% 139371 21.810%
+X-n783-k48 72386 90289 24.733% 76766 6.051% 83169 14.897% 100383 38.677% 92960 28.423% 96336 33.087% 83185 14.919% 77437 6.978%
+X-n801-k40 73305 124278 69.536% 77546 5.785% 85077 16.059% 91560 24.903% 87662 19.585% 87118 18.843% 86164 17.542% 77369 5.544%
+X-n819-k171 158121 193451 22.344% 178558 12.925% 177157 12.039% 183599 16.113% 185832 17.525% 179596 13.581% 174441 10.321% 171024 8.160%
+X-n837-k142 193737 237884 22.787% 207709 7.212% 214207 10.566% 229526 18.473% 221286 14.220% 230362 18.904% 208528 7.635% 203457 5.017%
+X-n856-k95 88965 152528 71.447% 92936 4.464% 101774 14.398% 99129 11.425% 106816 20.065% 105801 18.924% 98291 10.483% 94547 6.274%
+X-n876-k59 99299 119764 20.609% 104183 4.918% 116617 17.440% 119619 20.463% 114333 15.140% 114016 14.821% 107416 8.174% 105417 6.161%
+X-n895-k37 53860 70245 30.421% 58028 7.739% 65587 21.773% 79018 46.710% 64310 19.402% 69099 28.294% 64871 20.444% 58137 7.941%
+X-n916-k207 329179 399372 21.324% 385208 17.021% 361719 9.885% 383681 16.557% 374016 13.621% 373600 13.494% 352998 7.236% 346556 5.279%
+X-n936-k151 132715 237625 79.049% 196547 48.097% 186262 40.347% 220926 66.466% 190407 43.471% 161343 21.571% 163162 22.942% 172675 30.110%
+X-n957-k87 85465 130850 53.104% 90295 5.651% 98198 14.898% 113882 33.250% 105629 23.593% 123633 44.659% 102689 20.153% 90485 5.874%
+X-n979-k58 118976 147687 24.132% 127972 7.561% 138092 16.067% 146347 23.005% 139682 17.404% 131754 10.740% 129952 9.225% 125353 5.360%
+X-n1001-k43 72355 100399 38.759% 76689 5.990% 87660 21.153% 114448 58.176% 94734 30.929% 88969 22.962% 85929 18.760% 77739 7.441%
+Avg. Gap 29.658% 12.836% 16.796% 26.408% 19.607% 18.795% 12.303% 8.678%
 Appendix K Ablation Study
@@
 Method ATSP TSP OP PCTSP PDTSP ACVRP CVRP CVRPTW CVRPB OCVRP OCVRPTW Avg.gap
-[…]
-[…]
+w/o $\bm{\xi}_{i}$ 2.14% 0.48% 0.45% 1.06% 9.28% 2.84% 3.66% 6.61% 1.57% 7.87% 5.66% 3.78%
+w/ $\bm{\xi}_{i}$ 2.26% 0.57% 0.45% 1.06% 4.98% 3.06% 1.81% 6.13% 1.46% 3.24% 5.07% 2.74%
 To evaluate the effectiveness of node-type indicator $\bm{\xi}$, we conduct an ablation study about including and excluding $\bm{\xi}$.
-Table 13 shows that adding the node-type indicator reduces the average optimality gap from 3.78% to 2.74%, with large gains on the routing problem with relatively rich instance node roles, such as PDTSP, where the coexistence of a depot and paired pickup–delivery nodes (with inherent precedence coupling) allows the explicit node-role encoding to further boost performance.
+Table 21 shows that adding the node-type indicator reduces the average optimality gap from 3.78% to 2.74%, with large gains on the routing problem with relatively rich instance node roles, such as PDTSP, where the coexistence of a depot and paired pickup-delivery nodes (with inherent precedence coupling) allows the explicit node-role encoding to further boost performance.
 Meanwhile, URS maintains a competitive performance in addressing the problem of single-node roles.
@@
 Method ATSP TSP OP PCTSP PDTSP ACVRP CVRP CVRPTW CVRPB OCVRP OCVRPTW Avg.gap
-[…]
-[…]
+w/o $C_{t}$ 2.09% 0.45% 0.60% 1.20% 4.59% 4.03% 2.70% 5.71% 2.64% 4.02% 4.62% 2.97%
+w/ $C_{t}$ 2.26% 0.57% 0.45% 1.06% 4.98% 3.06% 1.81% 6.13% 1.46% 3.24% 5.07% 2.74%
 As described in Appendix E.3, the majority of VRP variants are already solved without supplying an additional problem state $C_{t}$.
@@
 The results in Table 22 show that URS maintains its competitive performance even in the absence of $C_{t}$.
-This indicates that explicit constraint-state features are not essential for the effectiveness of URS.
+This indicates that explicit constraint-state features are not essential to URS effectiveness.
 Interestingly, on (O)CVRPTW instances, the model without any explicit constraint features outperforms the version that supplies only capacity information.
@@
 K.3 Effects of Mixed Bias Module
-Table 15: Ablation of prior components in the Mixed Bias Module (MBM).
+Table 23: Ablation of prior components in the MBM.
 $\bm{D}$ $\bm{D^{\mathrm{T}}}$ $\bm{R}$ ATSP TSP OP PCTSP PDTSP ACVRP CVRP CVRPTW CVRPB OCVRP OCVRPTW Avg.gap
-[…]
-[…]
-[…]
-[…]
-To further assess the effectiveness of MBM, we conduct an ablation study that incorporates different combinations of problem-specific priors into the MBM.
-Owing to its flexibility, the MBM permits imposing multiple priors on the same shared attention layer.
-The results in Appendix K.3 underscore the complementary roles of the three prior components: (1) The transposed distance matrix $\bm{D^{\mathrm{T}}}$ captures directional asymmetry absent from the raw distance matrix, and its inclusion markedly stabilizes optimization on asymmetric variants; (2) Removing the pickup–delivery relation matrix $\bm{R}$ degrades performance on PDTSPs, confirming that explicit relational coupling is indispensable for these instances; (3) When all priors are fused, MBM can simultaneously learn the geometric and relational biases inherent in various problems, yielding the best overall performance.
+$\checkmark$ $\times$ $\times$ 16.64% 0.44% 0.25% 0.67% 7.33% 16.62% 2.42% 6.44% 3.96% 9.44% 5.47% 6.33%
+$\checkmark$ $\checkmark$ $\times$ 2.09% 0.45% 0.48% 0.82% 7.66% 9.11% 2.52% 6.89% 2.68% 5.70% 6.36% 4.07%
+$\checkmark$ $\times$ $\checkmark$ 16.60% 0.46% 0.49% 0.94% 4.49% 26.22% 4.16% 6.57% 3.07% 12.23% 5.53% 7.34%
+$\checkmark$ $\checkmark$ $\checkmark$ 2.26% 0.57% 0.45% 1.06% 4.98% 3.06% 1.81% 6.13% 1.46% 3.24% 5.07% 2.74%
+To further assess the effectiveness of MBM, we conduct an ablation study that examines different combinations of problem-specific priors within MBM.
+Owing to its flexibility, the MBM allows multiple priors to be imposed on the same shared attention layer.
+The results in Table 23 underscore the complementary roles of the three prior components: (1) The transposed distance matrix $\bm{D^{\mathrm{T}}}$ captures directional asymmetry absent from the raw distance matrix, and its inclusion markedly stabilizes optimization on asymmetric variants; (2) Removing the pickup-delivery relation matrix $\bm{R}$ degrades performance on PDTSPs, confirming that explicit relational coupling is indispensable for these instances; (3) When all priors are fused, MBM can simultaneously learn the geometric and relational biases inherent in various problems, yielding the best overall performance.
 K.4 Effects of Different Decoder Architectures
-Table 16: Effect of different decoder architectures on URS cross‑problem performance
+Table 24: Effect of different decoder architectures on URS cross-problem performance.
 Decoder ATSP TSP OP PCTSP PDTSP ACVRP CVRP CVRPTW CVRPB OCVRP OCVRPTW Avg.gap
@@
 URS-ReLD 6.48% 0.90% 0.86% 1.51% 5.81% 6.51% 2.33% 6.56% 2.33% 3.62% 5.64% 3.87%
-[…]
+URS 2.26% 0.57% 0.45% 1.06% 4.98% 3.06% 1.81% 6.13% 1.46% 3.24% 5.07% 2.74%
 We adopt an AM (Kool et al., 2019) as our basic encoder-decoder model.
-Multiple decoder refinements have been proposed around AM (Kwon et al., 2020; Huang et al., 2025; Zhou et al., 2024a).
+Multiple decoder refinements have been proposed around AM (Kwon et al., 2020; Huang et al., 2025; Zhou et al., 2026).
 To assess how decoder structure impacts multi-task performance, we additionally train two URS variants whose decoders follow POMO and ReLD, denoted URS-POMO and URS-ReLD.
 For stronger and fairer baselines, we apply Distance-aware Attention Reshaping (DAR) (Wang et al., 2025b) to both, and all decoder parameters are generated by a hypernetwork conditioned on the problem representation $\bm{\lambda}$ (see Appendix E.2).
-As shown in Table 16, URS-POMO and URS-ReLD still exhibit strong cross-problem robustness, while replacing the decoder with ICAM (Zhou et al., 2024a) yields a further overall improvement.
+As shown in Table 24, URS-POMO and URS-ReLD still exhibit strong cross-problem robustness, while replacing the decoder with ICAM (Zhou et al., 2026) yields a further overall improvement.
+Furthermore, the results in Table 24 reveal a notable optimization challenge.
+By incorporating an additional Feed-Forward Network layer compared to POMO, the ReLD decoder forces the hypernetwork to map to a larger parameter space, thereby increasing its learning difficulty and degrading performance.
+This optimization difficulty prevents URS-ReLD from matching the performance of the original ReLD-MTL, where weights are learned directly rather than being dynamically generated.
+Given the outstanding performance of ReLD and other heavy-decoder paradigms in single-task optimization (Luo et al., 2023; Drakulic et al., 2023), exploring how to adapt problem-conditioned generation mechanisms for heavier decoders is a highly promising direction.
+K.5 Effects of MBM & Parameter Generator on Zero-shot Generalization
+Table 25: Ablation of MBM, $\mathrm{WEIGHT}(\bm{\lambda})$, and $\mathrm{BIAS}(\bm{\lambda})$ on cross-problem zero-shot generalization. The symbol (*) indicates the seen problems in training.
+MBM $\mathrm{WEIGHT}(\bm{\lambda})$ $\mathrm{BIAS}(\bm{\lambda})$ ATSP (*) TSP (*) OP (*) PCTSP (*) PDTSP (*) ACVRP (*)
+$\times$ $\checkmark$ $\checkmark$ 16.64% 0.44% 0.25% 0.67% 7.33% 16.62%
+$\checkmark$ $\times$ $\times$ 4.24% 1.10% 1.04% 1.42% 6.38% 4.07%
+$\checkmark$ $\times$ $\checkmark$ 3.83% 1.16% 1.09% 1.52% 6.65% 4.27%
+$\checkmark$ $\checkmark$ $\times$ 2.79% 1.58% 1.39% 1.92% 7.05% 3.39%
+$\checkmark$ $\checkmark$ $\checkmark$ 2.26% 0.57% 0.45% 1.06% 4.98% 3.06%
+MBM $\mathrm{WEIGHT}(\bm{\lambda})$ $\mathrm{BIAS}(\bm{\lambda})$ CVRP (*) CVRPTW (*) CVRPB (*) OCVRP (*) OCVRPTW (*) CVRPL
+$\times$ $\checkmark$ $\checkmark$ 2.42% 6.44% 3.96% 9.44% 5.47% 0.98%
+$\checkmark$ $\times$ $\times$ 6.23% 7.22% 5.00% 3.99% 5.41% 2.92%
+$\checkmark$ $\times$ $\checkmark$ 4.94% 6.08% 3.00% 4.61% 5.21% 4.79%
+$\checkmark$ $\checkmark$ $\times$ 2.90% 6.60% 3.12% 4.47% 4.76% 1.46%
+$\checkmark$ $\checkmark$ $\checkmark$ 1.81% 6.13% 1.46% 3.24% 5.07% 0.43%
+MBM $\mathrm{WEIGHT}(\bm{\lambda})$ $\mathrm{BIAS}(\bm{\lambda})$ OCVRPB CVRPBL CVRPLTW OCVRPBTW CVRPBLTW OCVRPL
+$\times$ $\checkmark$ $\checkmark$ 15.95% 4.18% 2.89% 18.78% 10.93% 9.51%
+$\checkmark$ $\times$ $\times$ 9.54% 5.30% 2.73% 14.37% 9.80% 4.64%
+$\checkmark$ $\times$ $\checkmark$ 9.58% 3.25% 3.07% 14.16% 9.91% 3.88%
+$\checkmark$ $\checkmark$ $\times$ 9.92% 3.34% 3.08% 14.02% 9.92% 4.52%
+$\checkmark$ $\checkmark$ $\checkmark$ 9.35% 1.65% 2.67% 13.77% 9.18% 3.22%
+MBM $\mathrm{WEIGHT}(\bm{\lambda})$ $\mathrm{BIAS}(\bm{\lambda})$ CVRPBTW OCVRPBL OCVRPLTW OCVRPBLTW Avg.gap Best Sol.
+$\times$ $\checkmark$ $\checkmark$ 10.74% 16.14% 5.57% 18.80% 8.37% 3/22
+$\checkmark$ $\times$ $\times$ 9.67% 10.67% 5.23% 14.49% 6.16% 0/22
+$\checkmark$ $\times$ $\checkmark$ 9.63% 10.42% 5.51% 13.99% 5.93% 1/22
+$\checkmark$ $\checkmark$ $\times$ 9.79% 10.02% 5.81% 13.98% 5.72% 1/22
+$\checkmark$ $\checkmark$ $\checkmark$ 8.94% 9.47% 5.12% 13.72% 4.89% 17/22
+To empirically evaluate the necessity of these components, we have conducted tests on the 11 seen problems as well as 11 additional unseen CVRP variants widely investigated in current multi-task research (e.g., MVMoE (Zhou et al., 2024)).
+The results are presented in Table 25.
+We can observe that removing MBM leads to a significant performance drop on problems with complex geometric properties, such as ATSP, ACVRP, and PDTSP, which confirms MBM is critical for perceiving specific geometric biases.
+In addition, removing $\mathrm{WEIGHT}(\bm{\lambda})$ or $\mathrm{BIAS}(\bm{\lambda})$ also results in performance degradation across most tasks.
+In terms of overall performance, URS significantly outperforms any model where a single module is removed.
+The full URS model achieves the lowest average gap and provides the best solution on the majority of tasks (17/22).
+This demonstrates that the synergy of these components is essential for robust cross-problem generalization.
+K.6 Effects of Each UDR Component
+To clarify the independent contribution of each UDR component, we conduct ablations by removing each component in turn.
+As shown in Table 26, the absence of any single component leads to notable performance degradation.
+A complete UDR achieves the lowest average gap and provides the best solution on the majority of tasks (15/22), affirming the necessity of our UDR design for robust zero-shot generalization.
+Table 26: The ablation study of each UDR component.
+The symbol (*) indicates the seen problems in training.
+$\bm{\rho}$, $\bm{\omega}$, and $\bm{\xi}$ represent positional identifier, unified attribute set, and node-type indicator, respectively.
+Since removing positional identifiers prevents the use of instance augmentation, we have excluded it from all results to ensure consistency in the inference strategy.
+$\bm{\rho}$ $\bm{\omega}$ $\bm{\xi}$ ATSP (*) TSP (*) OP (*) PCTSP (*) PDTSP (*) ACVRP (*)
+$\times$ $\checkmark$ $\checkmark$ 7.64% 1.76% 1.17% 1.96% 8.54% 7.16%
+$\checkmark$ $\times$ $\checkmark$ 10.07% 1.98% 2.70% 7.05% 13.50% 13.36%
+$\checkmark$ $\checkmark$ $\times$ 5.43% 1.02% 1.08% 1.86% 12.72% 5.81%
+$\checkmark$ $\checkmark$ $\checkmark$ 5.51% 1.17% 0.96% 1.79% 7.71% 6.11%
+$\bm{\rho}$ $\bm{\omega}$ $\bm{\xi}$ CVRP (*) CVRPTW (*) CVRPB (*) OCVRP (*) OCVRPTW (*) CVRPL
+$\times$ $\checkmark$ $\checkmark$ 2.68% 6.88% 2.84% 4.64% 5.91% 1.29%
+$\checkmark$ $\times$ $\checkmark$ 4.15% 23.71% 5.24% 7.10% 19.10% 2.70%
+$\checkmark$ $\checkmark$ $\times$ 4.72% 7.67% 2.83% 9.64% 6.96% 3.33%
+$\checkmark$ $\checkmark$ $\checkmark$ 2.62% 6.80% 2.69% 4.54% 6.00% 1.23%
+$\bm{\rho}$ $\bm{\omega}$ $\bm{\xi}$ OCVRPB CVRPBL CVRPLTW OCVRPBTW CVRPBLTW OCVRPL
+$\times$ $\checkmark$ $\checkmark$ 11.87% 3.12% 3.34% 34.34% 44.03% 4.63%
+$\checkmark$ $\times$ $\checkmark$ 15.02% 5.47% 19.53% 30.88% 29.64% 6.98%
+$\checkmark$ $\checkmark$ $\times$ 26.06% 2.69% 4.13% 34.62% 26.55% 9.68%
+$\checkmark$ $\checkmark$ $\checkmark$ 11.54% 3.03% 3.38% 15.47% 10.70% 4.55%
+$\bm{\rho}$ $\bm{\omega}$ $\bm{\xi}$ CVRPBTW OCVRPBL OCVRPLTW OCVRPBLTW Avg.gap Best Sol.
+$\times$ $\checkmark$ $\checkmark$ 43.99% 11.99% 6.02% 34.66% 11.38% 3/22
+$\checkmark$ $\times$ $\checkmark$ 29.34% 14.84% 19.22% 31.02% 14.21% 0/22
+$\checkmark$ $\checkmark$ $\times$ 28.60% 25.96% 7.00% 35.93% 12.01% 4/22
+$\checkmark$ $\checkmark$ $\checkmark$ 10.51% 11.68% 6.08% 15.38% 6.34% 15/22
+K.7 Hypernetwork vs. Adapter
+We train URS-Adapter, a URS variant that adopts a distinct decoder for each of the 11 variants, and compare it with the original URS using a hypernetwork on 11 seen problems (see Table 27) and large-scale CVRP variants (see Table 28).
+URS consistently outperforms URS-Adapter across 11 seen problems, achieving higher parameter efficiency.
+This performance advantage holds consistently on large-scale CVRP variants, which further solidifies our conclusion.
+Table 27: Performance comparison between URS-Hypernet and URS-Adapter on 11 seen VRP variants.
+Method OP PCTSP PDTSP ACVRP
+Gap Time Gap Time Gap Time Gap Time
+URS-Adapter 0.88% 4s 1.47% 4s 5.69% 4s 3.83% 1.5m
+URS 0.45% 4s 1.06% 4s 4.98% 4s 3.06% 1.5m
+Method CVRPTW CVRPB OCVRP OCVRPTW
+Gap Time Gap Time Gap Time Gap Time
+URS-Adapter 6.34% 8s 3.21% 5s 4.06% 6s 4.94% 7s
+URS 6.13% 8s 1.46% 5s 3.24% 6s 5.07% 7s
+Method ATSP TSP CVRP Average Best
+Gap Time Gap Time Gap Time Gap Solution
+URS-Adapter 2.97% 1.2m 0.82% 4s 2.85% 6s 3.37% 1/11
+URS 2.26% 1.2m 0.57% 4s 1.81% 6s 2.74% 10/11
+Table 28: Generalization comparison between URS and URS-Adapter across diverse large-scale CVRP variants.
+Problem Method N = 1000 N = 2000 N = 3000 N = 4000 N = 5000
+Obj.(Gap) Time Obj.(Gap) Time Obj.(Gap) Time Obj.(Gap) Time Obj.(Gap) Time
+CVRPTW PyVRP 159.35(0.00%) 20m 337.15(0.00%) 40m 516.97(0.00%) 1h 618.58(0.00%) 2.7h 787.25(0.00%) 3.3h
+URS-Adapter 173.68(8.99%) 1.3m 367.22(8.92%) 10.8m 564.55(9.21%) 30.8m 677.91(9.59%) 1.2h 860.63(9.32%) 2.3h
+URS 172.58(8.30%) 1.3m 365.27(8.34%) 11.1m 559.03(8.14%) 30.1m 673.63(8.90%) 1.2h 853.79(8.45%) 2.3h
+CVRPB PyVRP 36.52(0.00%) 20m 51.80(0.00%) 40m 75.12(0.00%) 1h 93.52(0.00%) 2.7h 112.21(0.00%) 3.3h
+URS-Adapter 42.71(16.96%) 43s 62.73(21.11%) 5.5m 85.72(14.11%) 17.5m 107.17(14.59%) 41.9m 127.93(14.01%) 1.3h
+URS 34.86(-4.54%) 40s 50.39(-2.72%) 5.1m 72.01(-4.15%) 17.3m 91.74(-1.91%) 41.2m 111.93(-0.24%) 1.3h
+OCVRPTW PyVRP 90.91(0.00%) 20m 164.00(0.00%) 40m 224.16(0.00%) 1h 299.45(0.00%) 2.7h 367.86(0.00%) 3.3h
+URS-Adapter 105.96(16.56%) 1.1m 203.52(24.10%) 9.9m 287.70(28.35%) 28.6m 391.06(30.59%) 1h 483.19(31.35%) 1.9h
+URS 98.76(8.63%) 1m 178.07(8.58%) 7.6m 243.88(8.80%) 25.4m 325.64(8.74%) 1h 398.72(8.39%) 1.9h
 Appendix L Licenses for Used Resources
@@
 Resource Type Link License
-HGS-PyVRP (Wouda et al., 2024) Code https://github.com/PyVRP/PyVRP MIT License
+PyVRP (Wouda et al., 2024) Code https://github.com/PyVRP/PyVRP MIT License
 LKH3 (Helsgaun, 2017) Code http://webhotel4.ruc.dk/~keld/research/LKH-3/ Available for academic research use
-OR-Tools (Perron & Furnon, 2023) Code https://github.com/google/or-tools Apache-2.0 License
+OR-Tools (Perron and Furnon, 2023) Code https://github.com/google/or-tools Apache-2.0 License
 POMO (Kwon et al., 2020) Code https://github.com/yd-kwon/POMO/tree/master/NEW_py_ver MIT License
@@
 BQ (Drakulic et al., 2023) Code https://github.com/naver/bq-nco CC BY-NC-SA 4.0 license
-ICAM (Zhou et al., 2024a) Code https://github.com/CIAM-Group/ICAM MIT License
+ICAM (Zhou et al., 2026) Code https://github.com/CIAM-Group/ICAM MIT License
+ELG (Gao et al., 2024) Code https://github.com/gaocrr/ELG MIT License
 MatNet (Kwon et al., 2021) Code https://github.com/yd-kwon/MatNet/tree/main MIT License
 Heter-AM (Li et al., 2021) Code https://github.com/jingwenli0312/Heterogeneous-Attentions-PDP-DRL MIT License
-MTPOMO (Liu et al., 2024a) Code https://github.com/FeiLiu36/MTNCO MIT License
-MVMoE (Zhou et al., 2024b) Code https://github.com/RoyalSkye/Routing-MVMoE MIT License
-RouteFinder (Berto et al., 2024) Code https://github.com/ai4co/routefinder MIT License
+MTPOMO (Liu et al., 2024) Code https://github.com/FeiLiu36/MTNCO MIT License
+MVMoE (Zhou et al., 2024) Code https://github.com/RoyalSkye/Routing-MVMoE MIT License
+RouteFinder (Berto et al., 2025) Code https://github.com/ai4co/routefinder MIT License
 CaDA (Li et al., 2025a) Code https://github.com/CIAM-Group/CaDA MIT License
@@
 CVRPLib Set-X (Uchoa et al., 2017) Dataset http://vrp.galgos.inf.puc-rio.br/index.php/en/ Available for academic research use
-We list the used existing codes and datasets in Table 17, and all of them are open-sourced resources for academic usage.
-Appendix M LLM usage
-During the preparation of this manuscript, we utilized a Large Language Model (LLM) as a writing assistance tool.
-The primary purpose of its use was to improve the linguistic quality of the paper.
-Specifically, the LLM is used to polish the writing for enhanced clarity and fluency and to identify and correct grammatical, spelling, and punctuation errors.
-We confirm that all core research ideas, study design, data analysis, and conclusions are independently developed by the authors.
-The LLM does not contribute to the research ideation, methodology, or substantive academic content of the paper.
-Its function was strictly confined to language editing and proofreading to ensure the manuscript’s professionalism and readability.
+CVRPLib Set-AGS (Arnold et al., 2019) Dataset http://vrp.galgos.inf.puc-rio.br/index.php/en/ Available for academic research use
+We list the existing codes and datasets in Table 29, all of which are open-source resources for academic use.